### v2

# Smooth Tests for Normality in ANOVA

Abstract
The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models, yet it is rarely formally tested in practice. In this paper, we propose Neyman’s smooth tests for assessing the normality assumption across various types of ANOVA models. The proposed test statistics are constructed based on the Gaussian probability integral transformation of ANOVA residuals. Under the null hypothesis of normality, the test statistics are asymptotically Chi-square distributed, with degrees of freedom determined by the dimension of the smooth model (the number of orthonormal functions). A data-driven selection of the model dimension using a modified Schwarz’s criterion is also discussed. Simulation studies demonstrate the effectiveness of our proposed method.
Keywords: ANOVA; Estimation effect; Normality; Schwarz’s selection rule; Smooth test.
1 Introduction
Analysis of Variance (ANOVA) is an important and widely used method in exploratory and confirmatory data analysis (Gelman,, 2005), particularly for comparing group means and assessing the significance of factors in experimental designs. The theory of ANOVA models has been well established and developed in the literature, see, for example, Scheffé, (1959), Miller Jr, (1997), Wickens and Keppel, (2004), Dean et al., (2017), Hirotsu, (2017), Montgomery, (2017). A standard assumption in ANOVA models is that the random errors are normally distributed. Nonnormality may make the normal-based inference invalid. In terms of estimation, departures from normality generally undermine the validity of the normal-theory variance formulae for point estimates of variance components, because these formulae rely on theoretical results that represent the mean squares of random effects as constant multiples of Chi-square variables (Scheffé,, 1959). From the view of testing, the $F$-test employed in ANOVA is sensitive to nonnormality if the group sizes are unequal (Ali and Sharma,, 1996). The effects of nonnormality on size and power of the $F$-test are also extensively investigated in details (Pearson,, 1931; Gayen,, 1950; David and Johnson, 1951a, ; David and Johnson, 1951b, ; Srivastava,, 1959; Atiqullah,, 1962; Tiku,, 1964, 1971; Donaldson,, 1968; Janic-Wróblewska and Ledwina,, 2000). Therefore, it is crucial to evaluate the normality of ANOVA models.
While a broad range of normality tests have been proposed for observed data and linear models, comparatively little effort has been devoted to systematically checking the normality assumption in ANOVA. A direct approach to normality diagnosis is to plot a normal probability plot of ANOVA residuals (Dean et al.,, 2017; Hirotsu,, 2017), which is intuitive but lacks a theoretical guarantee. To our knowledge, formal hypothesis testing procedures for assessing normality in ANOVA appear to be relatively scarce in the existing literature. Existing methods are merely a simple application of classical normality tests to ANOVA residuals, see, for instance, Bonett and Woodward, (1990), Bonett and Seier, (2002) and Hwang and Wei, (2006). Although applicable in practice, these methods do not precisely provide a rigorous assessment of normality within the ANOVA framework, nor do they account for the specific structure and constraints inherent in ANOVA models.
In this article, following the spirit of Neyman, (1937), we propose a unified smooth test framework for assessing the normality assumption in various types of ANOVA models. Neyman’s smooth test has gained popularity across various scientific disciplines due to its well-known advantages in both theoretical and practical aspects. It has proven to be powerful in detecting against a broad class of alternatives; see, e.g., Bera and Ghosh, (2002), Thas, (2010), and Lehmann and Romano, (2022). Specifically, we first re-express the problem of testing for normality as testing for uniformity via the Gaussian probability integral transformation (PIT), and then construct the test statistics through the Gaussian PIT of ANOVA residuals. Under mild conditions for each model, the proposed test statistics are asymptotically Chi-square distributed under the null. Furthermore, we analyze the power properties of the tests under fixed and local alternatives. A modified Schwarz’s selection rule is proposed to determine the order of tests. Monte Carlo experiments demonstrate the good performance of our methodology in finite samples.
The advantages and contributions of our proposed tests are as follows. First, the proposed asymptotic Chi-square tests are simple to implement, in the sense that the critical values can be obtained directly from the limiting Chi-square distributions without bootstrapping. Second, we systematically investigate the properties of test statistics for three types of one-way fixed effects models with potentially diverging numbers of groups in a unified asymptotic framework. The parameter estimation effects in each model are well interpreted through our analysis procedure. The heterogeneity in group means and group variances does contribute to the specific form of test statistics and to the corresponding regularity conditions, but does not alter the convergence rate, which is only related to the total sample sizes under both the null and the alternatives. Lastly, we clarify the motivation and theoretical properties of data-driven smooth tests under ANOVA settings, and provide a revised limiting null distribution of the data-driven test statistics in finite samples.
The rest of this paper is organized as follows. Section 2 presents the general testing framework. Section 3 establishes the theoretical results for smooth tests in three types of ANOVA models. In Section 4, a data-driven approach is proposed to determine the order of tests. Simulation results in Section 5 demonstrate good finite-sample performance of our proposed methodology. Some extensions are discussed in Section 6. Section 7 concludes. The proof of theoretical results and some practical issues are gathered in the supplementary material.
2 The testing framework
Our problem of interest is whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$ (denoted as $\mathcal{N}(0,\sigma^{2})$) for some $\sigma>0$. We aim to test for the null hypothesis $H_{0}:e\sim\mathcal{N}(0,1)$ against the alternatives of nonnormality. To motivate our methodology, consider the transformed random variable $Z=\Phi(e)$, with $\Phi(\cdot)$ denoting the CDF of the standard normal distribution. Under the null hypothesis of normality, the CDF of $Z$ is given by
$G(z)=\mathrm{P}(Z\leq z)=\mathrm{P}\left(e\leq\Phi^{-1}(z)\right)=\Phi(\Phi^{-1}(z))=z.$
Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1$ for $z\in[0,1]$. That is, $Z\sim\mathcal{U}[0,1]$ under $H_{0}$, where $\mathcal{U}[0,1]$ marks the uniform distribution on the interval $[0,1]$. Under the alternatives, the PDF $g(z)$ differs from 1. The distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ provides the basis for Neyman’s smooth test. Neyman, (1937) considered the following smooth alternative to the uniform density:
$h\left(z\right)=c\left(\bm{\theta}_{K}\right)\exp\left[\sum_{k=1}^{K}\theta_{k}\pi_{k}\left(z\right)\right],\quad 0\leq z\leq 1,$ (2.1)
where $c(\bm{\theta}_{K})$ is a constant of integration depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ is an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
[…]
The null hypothesis that $Z\sim\mathcal{U}[0,1]$ can then be tested by testing $\theta_{1}=\theta_{2}=\ldots=\theta_{K}=0$ in (2.1).
If we had independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ and $\sigma$ were known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
$\Psi_{K}^{2}=\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}(Z_{i})\right)^{2},$ (2.3)
where $K$ is some fixed positive integer and $Z_{i}=\Phi(e_{i})=\Phi(\varepsilon_{i}/\sigma)$. Under $H_{0}$, $n\Psi_{K}^{2}$ converges asymptotically to a Chi-square distribution with $K$ degrees of freedom (denoted as $\chi_{K}^{2}$), and the test has the local optimal properties of Rao’s score test.
However, it is always the case that $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown. We may use some approximation $\{\hat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and some estimator $\hat{\sigma}$ for $\sigma$ (e.g., the residuals and the sample standard deviation, respectively). Then, consider the following approximation for $Z_{i}$:
$\hat{Z}_{i}=\Phi(\hat{e}_{i})=\Phi\left(\frac{\hat{\varepsilon}_{i}}{\hat{\sigma}}\right),\quad i=1,\ldots,n.$
With $\{\hat{Z}_{i}\}_{i=1}^{n}$ replacing $\{Z_{i}\}_{i=1}^{n}$, we can construct a feasible smooth test statistic of the following form:
$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\hat{Z}_{i}\right)\right)^{2}.$
Due to the replacement of $Z_{i}$ by $\hat{Z}_{i}$, the sequence $\{\hat{Z}_{i}\}_{i=1}^{n}$ may not be i.i.d. any longer. Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\hat{Z}_{i})$. Compared with $\Psi_{K}^{2}$ in (2.3), $\hat{Z}_{i}$ has non-negligible estimation effects, as demonstrated in the theoretical results below. Consequently, the presence of estimation effects invalidates the above form of the test statistic and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution. This motivates us to study the estimation effects carefully and then derive the correct quadratic-form test statistics.
It is also important to address the issue of sample sizes in ANOVA models. Throughout this section, to highlight the essence of smooth tests, we do not directly consider the ANOVA models but first examine the testing for a single random sequence of sample size $n$. This simplified setting captures the core ideas of the methodology and facilitates the derivation of key asymptotic properties. In the complete ANOVA framework, additional complications arise due to multiple sample sizes and the indexing structure, which require a more elaborate treatment. Besides, from a practical perspective, our method is expected to accommodate scenarios in which the number of groups grows with the total sample size. In subsequent analysis, we will systematically tackle these issues for specific ANOVA models.
3 Testing for normality in one-way fixed effects models
In this section, we establish smooth tests for normality in three types of one-way fixed effects models where either the group means or the group variances are equal. We examine the testing procedure and the theoretical properties separately for each case. Let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observations of interest coming from $J$ groups, with $N_{j}$ being the sample sizes in group $j$ for $1\leq j\leq J$. Define $N=\sum_{j=1}^{J}N_{j}$ as the total sample size. In our theoretical framework, the number of groups $J$ is allowed to diverge as $N$ increases.
3.1 Same mean and same variance
We first investigate the ANOVA models with equal group means and equal group variances. Consider the following model:
$Y_{ij}=\mu+\varepsilon_{ij}=\mu+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.1)
where $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. random errors with mean zero and variance one, $\mu$ is the common group mean, and $\sigma$ is the standard deviation of random errors. We also define $Z_{ij}=\Phi(e_{ij})$.
We estimate $\mu$ and $\sigma^{2}$ in (3.1) by the sample mean $\hat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}\right)^{2}$, respectively. Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu})/\hat{\sigma}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To derive the test statistic and its asymptotic distribution, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij}),k=1,\ldots,K$. The following two assumptions are imposed for our asymptotic analysis.
Assumption 3.1.
$\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. with continuous CDF $F(x)$, PDF $f(x)$, mean zero, variance one, and finite fourth moments.
Assumption 3.2.
For $k=1,\ldots,K$, $\pi_{k}(\cdot)$ are twice differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
The two mild conditions are in line with similar assumptions adopted in the literature of Neyman’s smooth tests, including Bera et al., (2013) and Song and Xiao, (2022). Introduce two constants:
$c_{1k}=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z,\quad c_{2k}=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z,\quad k=1,\ldots,K.$
Theorem 1.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $N\to\infty$.
Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ is asymptotically equivalent to the sum of three terms: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$. The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$.
With the assistance of Theorem 1, by the central limit theorem (CLT), we can obtain the asymptotic normality of the $K$-dimensional vector
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$ (3.2)
$\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Sigma}_{K}\right),$
where $\bm{\Sigma}_{K}=\left(\sigma_{kl}\right)_{K\times K}$ is given by
$\sigma_{kl}$ $=\mathrm{E}\left[\left(\pi_{k}(Z)-c_{1k}e-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right)\left(\pi_{l}(Z)-c_{1l}e-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right)\right]$
$=\delta_{kl}-c_{1k}c_{1l}-\frac{1}{2}c_{2k}c_{2l}.$
Based on this fact, our proposed feasible smooth test statistic is defined as
$\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\bm{\Sigma}_{K}^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.3)
The following corollary establishes the limiting null distribution of $\hat{\Psi}_{K}^{2}$ defined in (3.3).
Corollary 1.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$,
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2},$
as $N\to\infty$.
Corollary 1 demonstrates an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\hat{\Psi}_{K}^{2}$. Given the asymptotic significance level of $\alpha$, we reject the null hypothesis if
$N\hat{\Psi}_{K}^{2}>\chi_{K,1-\alpha}^{2},$
where $\chi_{K,1-\alpha}^{2}$ denotes the $(1-\alpha)$-th quantile of the $\chi_{K}^{2}$ distribution.
Now we turn to the asymptotic behavior of $\hat{\Psi}_{K}^{2}$ under the fixed alternatives and a Pitman-type sequence of local alternatives. For the fixed alternatives, we consider the following form:
[…]
Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed, which leads to different asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ from those under $H_{0}$. To formally present the corresponding theoretical results under $H_{1}$, we need some notations. Let
$d_{1k}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)f(x)\mathrm{d}x,$
$d_{2k}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)xf(x)\mathrm{d}x,$
$d_{3k}=\mathrm{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)xf(x)\mathrm{d}x,$
and
$d_{4k}=\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)x^{2}f(x)\mathrm{d}x,$
for $k=1,\ldots,K$. Note that under $H_{0}$, $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ due to $f(x)=\phi(x)$. The following theorem demonstrates the asymptotic properties under $H_{1}$.
Theorem 2.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.5)
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $N\to\infty$. Furthermore,
$\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K},$
and
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right),$ (3.6)
where $\bm{a}_{K}\equiv(a_{1},\ldots,a_{K})^{\top}=(\mathrm{E}[\pi_{1}(Z)],\ldots,\mathrm{E}[\pi_{K}(Z)])^{\top}$ and $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ is given by
$\xi_{kl}=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
From Theorem 2, one can see that under $H_{1}$ in (3.4), if there exists at least one $1\leq k\leq K$ such that
$a_{k}\equiv\mathrm{E}[\pi_{k}(Z)]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))f(x)\mathrm{d}x=\int_{0}^{1}\pi_{k}(z)\frac{f(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z\neq 0,$
the smooth test statistic $N\hat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, and the asymptotic power of the test is 1. Specifically, the power function is given by
$1-\Phi\left(\left(4N\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)^{-1/2}\chi_{K,1-\alpha}^{2}-\left(N\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}/4\right)^{1/2}\right).$
For further power analysis, we consider a Pitman-type sequence of local alternatives converging to the null hypothesis at a suitable rate, which is specified as follows:
$H_{1L}:F(x)=(1-\delta_{N})\Phi(x)+\delta_{N}Q(x),$ (3.7)
where $Q\left(x\right)$ (which admits PDF $q\left(x\right)$) represents some distribution function that is different from $\Phi(x)$, and $\delta_{N}\to 0$ as $N\to\infty$. Denote
$\Delta_{k}=\int_{0}^{1}\pi_{k}(z)\frac{q(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z,\quad k=1,\ldots,K.$
The following theorem presents the theoretical properties of $N\hat{\Psi}_{K}^{2}$ under the local alternatives. Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the local power of the test.
Theorem 3.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.8)
$+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $N\to\infty$. Furthermore,
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
where $\bm{\Delta}_{K}=(\Delta_{1},\ldots,\Delta_{K})^{\top}$, and $\chi^{2}_{K}\left(\tau\right)$ denotes the noncentral $\chi^{2}$ distribution with $K$ degrees of freedom and nonnegative noncentrality parameter $\tau$.
It is worthwhile to mention that $N^{-1/2}$ is the critical order of magnitude for our test to have a nontrivial power against the Pitman sequence of local alternatives. It would be impossible to distinguish the alternatives from the null if $\delta_{N}=o(N^{-1/2})$. Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\hat{\Psi}_{K}^{2}$ has nontrivial asymptotic power against the local alternatives.
Remark 1.
We should emphasize that the proposed smooth test is not a “consistent” test in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives. Only under alternatives where $\mathrm{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power go to 1. Otherwise, if $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathrm{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\hat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$. For example, if $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ such that $\mathrm{E}[e]=0$ and $\operatorname{Var}[e]=1$, and we take the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$, then for $Z=\Phi(e)$, we have
$\mathrm{E}[Z]$ $=\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}\Phi(z)\mathrm{d}z$
$=\frac{1}{2\sqrt{3}}z\Phi(z)\Big{|}_{-\sqrt{3}}^{\sqrt{3}}-\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}z\mathrm{d}\Phi(z)$
$=\frac{1}{2}\left(\Phi\left(\sqrt{3}\right)+\Phi\left(-\sqrt{3}\right)\right)$
$=\frac{1}{2},$
which leads to $\mathrm{E}[\pi_{1}(Z)]=0$. In fact, smooth tests are neither directional nor omnibus; that is, they maintain decent power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see Bera et al., (2013) for details.
3.2 Different means and same variance
Next we allow the group means $\mu$ in model (3.1) to vary across the group index $j$, in the sense that
$Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.9)
where $\{\mu_{j}\}_{j=1}^{J}$ represents the potentially different means of each group. Let $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ be estimators of $\mu_{j},j=1,\ldots,J$ and $\sigma^{2}$, respectively. Correspondingly we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. The following theorem presents the null limiting properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ under the settings for model (3.9).
Theorem 4.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\hat{\Psi}_{K}^{2}$ defined in (3.3), we still have
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}.$
Theorem 4 is parallel with Theorem 1 and Corollary 1 except for extra conditions regarding the group sample sizes and the number of groups. These additional conditions are imposed for technical considerations and to account for the heterogeneity across group means from model (3.9). We illustrate the intuition behind the conditions as follows. First, we require that the sample size of each group, $N_{j}$, tends to infinity. For model (3.9) where $\mu_{j}$ is allowed to be distinct, only the observations in group $j$ can be used to estimate $\mu_{j}$, thus $N_{j}\to\infty$ is necessary for the consistency of $\hat{\mu}_{j}$. Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must be satisfied jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$, which rules out the case that the number of groups grows faster than the group sample sizes. In contrast, for model (3.1) in the previous section, only the total sample size $N\to\infty$ suffices to derive the asymptotic properties regardless of the magnitude of $N_{j}$s and $J$. This is because in model (3.1), all groups share the same mean $\mu$ and variance $\sigma^{2}$. Thus, the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are all i.i.d., and we can just pool them together for efficient estimation and inference.
The following two theorems establish the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ (defined by (3.3)) under the alternatives. The conditions are the same as those of Theorem 4, and the results are parallel with those in Theorems 2 and 3.
Theorem 5.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore,
$\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K},$
and
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right).$
Theorem 6.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore,
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right).$
Theorems 4, 5, and 6 provide an asymptotic $\chi^{2}$ test for the normality of model (3.9) and demonstrate its power properties. Such a test is quite similar to that in Section 3.1. When the group means in (3.9) are all equal in the sense that $\mu_{1}=\ldots=\mu_{J}=\mu$, the methodology in this section can be reduced to that in Section 3.1 for model (3.1), which validates the unified inferential framework and reflects the effects of group heterogeneity in means.
3.3 Same mean and different variances
Now we allow the group variances $\sigma$ in model (3.1) to be different among the groups, in the sense that
$Y_{ij}=\mu+\varepsilon_{ij}=\mu+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.10)
where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially different standard deviations of each group. This time, we adopt the estimators $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\hat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2},j=1,\ldots,J$, respectively. Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of variation across groups. Let $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu})/\hat{\sigma}_{j}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. Define $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ regarding the relative proportion among the group sample sizes. Here we further impose the following conditions for model (3.10): (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$. They both imply that the sample sizes and the standard deviations of each group are of the same magnitude.
Like the procedure in Sections 3.1 and 3.2 above, we start from the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ here, see the following theorem.
Theorem 7.
Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.11)
$+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$.
Note that the decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4. The difference lies in the parameter estimation effect of $\hat{\mu}$ characterized by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right hand side of (3.11). Unlike the adoption of the sample mean for estimating $\mu$ in model (3.1), here we consider a new form of $\hat{\mu}$ for model (3.10), which leads to a different estimation effect term. Specially, when all groups share the same variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term can be simplified as $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect reduces to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$. In contrast, the estimation effect of the variances is still aggregated as the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$ can be derived from (3.11). Under the conditions of Theorem 7, by CLT,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
$\to_{d}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{Z}_{j},$
where $\bm{Z}_{1},\ldots,\bm{Z}_{J}$ are independently distributed as
$\bm{Z}_{j}\sim\mathcal{N}_{K}\left(\bm{0},\bm{\Omega}_{K}^{(j)}\right),$
with $\bm{\Omega}_{K}^{(j)}=\left(\omega_{kl}^{(j)}\right)_{K\times K}$ given by
$\omega_{kl}^{(j)}$ $=\mathrm{E}\left\{\pi_{k}(Z)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right\}$
$\qquad\left\{\pi_{l}(Z)-c_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right\}$
$=\delta_{kl}-\frac{2c_{1k}c_{1l}\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{c_{1k}c_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}-\frac{c_{2k}c_{2l}}{2}.$
Therefore, under $H_{0}$, we have
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right).$
By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ by $\hat{\sigma}_{j}$, $\hat{p}_{j}=N_{j}/N$ and $\hat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted as $\widehat{\bm{\Omega}}_{K}^{(j)}$. Correspondingly, the test statistic takes the form
$\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}\hat{p}_{j}\widehat{\bm{\Omega}}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.12)
We also define the infeasible version,
$\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.13)
Corollary 2.
Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under $H_{0}:e\sim\Phi(x)$,
[…]
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$.
Corollary 2 demonstrates that under the null hypothesis, both the feasible test statistic $\hat{\Psi}_{K}^{2}$ (in (3.12)) and its infeasible version $\tilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed after multiplied by $N$, the total sample size. Although the test statistics considered here take a slightly different form from those in the earlier setting, the limiting Chi-square distribution is retained and the convergence rate remains $N$, thereby reinforcing the validity of our unified theoretical framework. The following two theorems establish the asymptotic behavior of $\tilde{\Psi}_{K}^{2}$ under the fixed and local alternatives, which is also analogous to that of $\hat{\Psi}_{K}^{2}$ in Sections 3.1 and 3.2.
Theorem 8.
Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right])-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$. Furthermore,
$\tilde{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K},$
and
$\sqrt{N}\left(\tilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Upsilon}_{K}\bm{a}_{K}\right),$ (3.15)
where
$\bm{\Upsilon}_{K}=\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\sum_{j=1}^{J}p_{j}\bm{\Lambda}_{K}^{(j)}\right)\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1},$
and $\bm{\Lambda}_{K}^{(j)}=\left(\lambda_{kl}^{(j)}\right)_{K\times K}$ is given by
$\lambda_{kl}^{(j)}=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathrm{E}\left[e^{3}\right]$
$+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
Theorem 9.
Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right])-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$. Furthermore,
$N\tilde{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{\Delta}_{K}\right).$
Remark 2.
To conclude this section, we summarize the regularity conditions of the three models above in Table 1. On the one hand, one can find that more complex model structures lead to more restrictions on parameters and sample sizes. The heterogeneity across groups indeed poses additional challenges for theoretical analysis. On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can always be established for each of the three cases. This demonstrates that our proposed method possesses a unified theoretical framework while retaining the ability to capture the distinctive features of different ANOVA models.
Table 1: The regularity conditions for the three models
Model Conditions
Same mean & same variance (3.1) $N\to\infty$
Different means & same variance (3.9)
$\min\{N_{1},\ldots,N_{J}\}\to\infty$,
$J=o(N^{1/2})$, $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$
Same mean & different variances (3.10)
$\min\{N_{1},\ldots,N_{J}\}\to\infty$,
$J=o(N^{1/2})$ , $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$,
$\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$,
$\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$
4 Data-driven choice of $K$
The methodology in Section 3 is based on the fact that the order of smooth alternatives (2.1), $K$, is a predetermined fixed integer before the tests are performed. In fact, the order $K$ has a key impact on the performance of smooth tests (Ledwina,, 1994; Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). If $K$ is large, the alternatives we test against contain redundant components that do not contribute to the test statistic markedly but increase the number of degrees of freedom. This will make smooth tests more omnibus and lead to power dilution. If $K$ is small, as described in Section 3, it is more likely to happen that $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, which makes the tests exhibit no power. Therefore, it is crucial to select an appropriate $K$ before applying the test. In this section, we focus on the data-driven choice of $K$ for the three cases in Section 3.
In the literature of Neyman’s smooth tests, the spirit of data-driven methods was first introduced by Ledwina, (1994) for the problem of testing for uniformity. The data-driven procedure consists of two steps. First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to find a suitable dimension of the smooth model for the data. Second, Neyman’s smooth test is performed within the selected model space, yielding a data-driven test statistic. Therefore, the data-driven smooth test serves as a kind of first selection followed by a more precise instrument, being Neyman’s test in the “right” direction. Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were shown in Kallenberg and Ledwina, 1995a and Kallenberg and Ledwina, 1995b , Inglot and Ledwina, (1996). Later, some modified versions of Schwarz’s selection rule were employed to address more complicated issues, such as testing for composite hypotheses (Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Inglot et al.,, 1997), testing for independence (Kallenberg and Ledwina,, 1999), two-sample testing problems (Janic-Wróblewska and Ledwina,, 2000), and so on. The modified selection rules utilize the smooth test statistics and circumvent the need to compute the maximized log-likelihood function of smooth models; thus, they are easier to implement.
Motivated by the strategies mentioned above, we adopt a modified Schwarz’s rule to determine the order $K$ for our problem. The selection rule is specified as
$\hat{K}=\min\left\{\mathop{\arg\max}\limits_{1\leq k\leq D}\left(N\hat{\Psi}_{k}^{2}-k\log N\right)\right\},$ (4.1)
where $D$ is some fixed positive integer, and $\hat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) up to specific scenarios. The corresponding data-driven test statistic is $N\Psi_{\hat{K}}^{2}$, where $\hat{K}$ is given by (4.1). The form of (4.1) is mainly inspired by Ducharme and Fontez, (2004), Ducharme and Lafaye de Micheaux, (2004), Kraus, (2007) and Duchesne et al., (2016) where the upper bound of selection $D$ is fixed. An alternative is to set $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000). Although the diverging upper bound helps to improve the consistency of smooth tests theoretically, the rate of divergence is required to be quite slow, and the empirical power levels off rapidly as $D$ increases in their simulation studies. Therefore, our proposed selection rule (4.1) with fixed $D$ is reasonable and simple to implement in practice.
To establish the properties of $\hat{K}$ and $N\Psi_{\hat{K}}^{2}$, we need a revised version of the alternative hypotheses $H_{1}$ (3.4) and Assumption 3.2. Consider the following alternatives:
$H_{1}^{\prime}:\mathrm{E}[\pi_{1}(Z)]=\ldots=\mathrm{E}[\pi_{K_{0}-1}(Z)]=0,\mathrm{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D.$ (4.2)
In other words, $H_{1}^{\prime}$ encompasses a broader range of alternatives than $H_{1}$ due to the replacement of the fixed order $K$ with the larger $D$. Analogously, the condition on the orthonormal system is now broadened to include all functions up to order $D$; see the revised assumption below.
Assumption 4.1.
For $k=1,\ldots,D$, $\pi_{k}(\cdot)$ are two times differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
The following theorem demonstrates the properties of data-driven tests under the null and the revised alternatives.
Theorem 10.
Suppose Assumptions 3.1 and 4.1 hold.
  • (1).
    For model (3.1) in Section 3.1, as $N\to\infty$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$ and $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\hat{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
  • (2).
    For model (3.9) in Section 3.2, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$ and $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\hat{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
  • (3).
    For model (3.10) in Section 3.3, with additional conditions such that (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$, $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$ and $N\tilde{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\tilde{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
Theorem 10 presents the unified results of data-driven smooth tests for the three cases in Section 3. On the one hand, under the null hypothesis, the probability of $\{\hat{K}=1\}$ tends to $1$ asymptotically, which implies that the first-order smooth model fits $\{\hat{Z}_{ij}\}_{i=1,j=1}^{N_{j}.J}$ best among the $D$ candidate models. From the perspective of hypothesis testing, it is $\hat{\Psi}_{1}^{2}$ that is informative and adequate for diagnosing the uniformity/normality. On the other hand, the data-driven selection procedure helps identify the mechanism behind nonuniformity/nonnormality and leads to a more powerful test against a broader range of alternatives, $H_{1}^{\prime}$. These findings are parallel with existing results about data-driven smooth tests.
However, in practice, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics is often inaccurate in finite samples (Kallenberg and Ledwina, 1995b, ; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). To overcome this issue in our problem, provided the nested orthonormal system (i.e. $\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the instruction of Kallenberg and Ledwina, 1995b , Inglot et al., (1997), Kallenberg and Ledwina, (1997), Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007) and adopt a finite-sample approximation for the null distribution $N\hat{\Psi}_{\hat{K}}^{2}$:
[…]
Such an approximation primarily accounts for the selection uncertainty of $\hat{K}$. In finite samples, under the null, the empirical probability of observing $\{\hat{K}=1\}$ may not be exactly one, as the event $\{\hat{K}=2\}$ can still occur with a small but negligible probability. Accordingly, (4.3) is obtained from $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x)\approx\mathrm{P}(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1)+\mathrm{P}(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2)$. The reliability and effectiveness of $H(x)$ are demonstrated through simulation studies in the following section.
5 Simulation
In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the proposed tests developed in Sections 3 and 4. Following Neyman, (1937) and Song and Xiao, (2022), we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$. Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of Duchesne et al., (2016), and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material. From a computational perspective, the test statistic $N\hat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in Duchesne et al., (2016), which leverages the closed-form of $\bm{\Sigma}_{K}^{-1}$ (see equations (12) and (13) therein). Since Section 3.1 is a special case of Section 3.2 or 3.3, we focus our evaluation on the settings described in Sections 3.2 and 3.3 through two numerical experiments.
In the first experiment, we generate observations with different means and the same variances as follows:
$Y_{ij}^{(0,1)}\sim\mathcal{N}\left(5j,4\right),\quad Y_{ij}^{(1,1)}\sim\chi_{2}^{2}+(5j-2),\quad i=1,\ldots,jm,\quad j=1,\ldots,5,$
where $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$, $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$ and $H_{1}$ respectively. Note that
$\mathrm{E}[Y_{ij}^{(0,1)}]=\mathrm{E}[Y_{ij}^{(1,1)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,1)}]=\operatorname{Var}[Y_{ij}^{(1,1)}]=4$. The sample size parameter $m$ varies from 10 to 150 in increments of 10. We set the significance level $\alpha=5\%$ and perform $500$ replications. We evaluate the performance of the test statistic $N\hat{\Psi}_{K}^{2}$ introduced in Section 3.2, with $1\leq K\leq 5$, as well as the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ from Section 4, where $\hat{K}$ is determined by (4.1) with $D=5$. Table 2 reports the empirical rejection rates under $H_{0}$. For fixed values of $K=1,\ldots,5$, $N\hat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes, such as $m=10$ (corresponding to a total sample size of $N=150$). For the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$, the results indicate that the limiting $\chi^{2}_{1}$ distribution does not adequately approximate the null distribution, whereas the revised distribution $H(x)$ achieves better size control. The empirical frequency of $\hat{K}$ under $H_{0}$ and $H_{1}$ is presented in Tables 3 and 4, respectively. Under $H_{0}$, the frequency of $\{\hat{K}=1\}$ approaches 1 as $m$ increases, while under $H_{1}$, $\hat{K}$ is consistently greater than 1, which aligns with our theoretical expectations. Additionally, Figure 1 displays the sample means and corresponding error bars (sample mean $\pm$ sample standard deviation) of $\hat{K}$ across different values of $m$. Here, under our data-generating processes, under $H_{0}$, the sample mean of $\hat{K}$ remains close to 1 regardless of the sample size, and the variability diminishes as $m$ increases, indicating consistent model selection under the null. Under $H_{1}$, $\hat{K}$ increases steadily and converges to the maximum selection range $D=5$ as $m$ grows, with reduced variability. These findings align with our theoretical results and demonstrate the adaptiveness of the data-driven procedure.
Table 2: Empirical rejection rates under $H_{0}$ in the first experiment, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
$10$ 0.050 0.046 0.048 0.048 0.052 0.076 0.050
2020 0.072 0.060 0.050 0.046 0.050 0.054 0.048
$30$ 0.054 0.046 0.054 0.038 0.056 0.074 0.062
$40$ 0.046 0.048 0.060 0.054 0.048 0.054 0.050
$50$ 0.044 0.060 0.052 0.052 0.048 0.040 0.038
$60$ 0.066 0.044 0.048 0.052 0.058 0.048 0.042
$70$ 0.044 0.048 0.038 0.046 0.048 0.072 0.064
$80$ 0.050 0.044 0.048 0.038 0.042 0.074 0.068
$90$ 0.040 0.060 0.046 0.054 0.046 0.048 0.042
$100$ 0.062 0.056 0.042 0.050 0.048 0.058 0.052
$110$ 0.060 0.046 0.066 0.052 0.050 0.068 0.054
$120$ 0.034 0.044 0.038 0.038 0.036 0.058 0.056
$130$ 0.038 0.050 0.042 0.052 0.050 0.058 0.048
$140$ 0.038 0.050 0.048 0.042 0.044 0.044 0.040
$150$ 0.064 0.056 0.048 0.058 0.058 0.046 0.042
Table 3: Empirical frequency of $\hat{K}$ under $H_{0}$ in the first experiment
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.976 0.024 0 0 0
2020 0.974 0.018 0.006 0 0.002
$30$ 0.986 0.014 0 0 0
$40$ 0.990 0.010 0 0 0
$50$ 0.998 0.002 0 0 0
$60$ 0.988 0.012 0 0 0
$70$ 0.992 0.008 0 0 0
$80$ 0.990 0.008 0.002 0 0
$90$ 0.988 0.012 0 0 0
$100$ 0.996 0.004 0 0 0
$110$ 0.994 0.004 0.002 0 0
$120$ 0.992 0.006 0.002 0 0
$130$ 0.998 0.002 0 0 0
$140$ 0.992 0.008 0 0 0
$150$ 0.996 0.004 0 0 0
Table 4: Empirical frequency of $\hat{K}$ under $H_{1}$ in the first experiment
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.032 0.004 0 0.418 0.546
2020 0 0 0 0.314 0.686
$30$ 0 0 0 0.174 0.826
$40$ 0 0 0 0.098 0.902
$50$ 0 0 0 0.064 0.936
$60$ 0 0 0 0.036 0.964
$70$ 0 0 0 0.022 0.978
$80$ 0 0 0 0.018 0.982
$90$ 0 0 0 0.008 0.992
$100$ 0 0 0 0.008 0.992
$110$ 0 0 0 0.002 0.998
$120$ 0 0 0 0 1
$130$ 0 0 0 0 1
$140$ 0 0 0 0.002 0.998
$150$ 0 0 0 0 1
Figure 1: The sample means and error bars of $\hat{K}$ in the first experiment
Figure 2: The sample means and error bars of $\hat{K}$ in the second experiment
In the second experiment, we generate observations with the same means and different variances as follows:
$Y_{ij}^{(0,2)}\sim\mathcal{N}\left(8,j^{2}\right),\ Y_{ij}^{(1,2)}\sim\mathcal{U}\left[8-\sqrt{3}j,8+\sqrt{3}j\right],\ i=1,\ldots,jm,\ j=1,\ldots,5,$
where $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$, $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$ and $H_{1}$ respectively. Note that
$\mathrm{E}[Y_{ij}^{(0,2)}]=\mathrm{E}[Y_{ij}^{(1,2)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,2)}]=\operatorname{Var}[Y_{ij}^{(1,2)}]=j^{2}$. We adopt the same settings as in the first experiment. The results are analogous to those observed previously; see Tables 5, 6, and 7, as well as Figure 2. However, from Table 5, we observe that under $H_{1}$, the test statistic $N\hat{\Psi}_{1}^{2}$ fails to demonstrate power. This is because $\mathrm{E}[\pi_{1}(Z)]=\mathrm{E}[\pi_{1}(\Phi((Y_{ij}^{(1,2)}-8)/j))]=0$, indicating that $N\hat{\Psi}_{K}^{2}$ may lack consistency against certain directions of alternatives, particularly when $K$ is small. Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, the event $\{\hat{K}=1\}$ does not occur, which is consistent with Theorem 10 (3), where it is established that $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ under $H^{\prime}_{1}$ with $K_{0}=2$.
To summarize the simulation results, we conclude that the proposed test statistic $N\hat{\Psi}_{K}^{2}$ maintains the nominal significance level well under $H_{0}$, and demonstrates high power under the alternatives unless $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$. The data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ using the approximated null distribution $H(x)$ performs reliably in controlling the Type  error and achieves excellent power performance.
Table 5: Empirical rejection rates in the second experiment, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
$H_{0}$ $H_{1}$
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
$10$ 0.034 0.044 0.040 0.030 0.028 0.062 0.042 0.040 0.998 0.998 0.996 0.994 0.998 0.998
2020 0.062 0.064 0.060 0.056 0.058 0.060 0.050 0.034 1 1 1 1 1 1
$30$ 0.040 0.040 0.038 0.042 0.048 0.060 0.044 0.034 1 1 1 1 1 1
$40$ 0.038 0.042 0.044 0.028 0.032 0.052 0.040 0.024 1 1 1 1 1 1
$50$ 0.044 0.042 0.042 0.042 0.044 0.060 0.048 0.036 1 1 1 1 1 1
$60$ 0.062 0.080 0.052 0.064 0.074 0.082 0.066 0.016 1 1 1 1 1 1
$70$ 0.062 0.066 0.060 0.050 0.062 0.052 0.042 0.050 1 1 1 1 1 1
$80$ 0.046 0.048 0.042 0.058 0.036 0.040 0.038 0.042 1 1 1 1 1 1
$90$ 0.060 0.058 0.050 0.052 0.046 0.052 0.042 0.032 1 1 1 1 1 1
$100$ 0.052 0.038 0.046 0.058 0.042 0.054 0.048 0.018 1 1 1 1 1 1
$110$ 0.040 0.034 0.038 0.032 0.040 0.052 0.046 0.032 1 1 1 1 1 1
$120$ 0.056 0.044 0.066 0.046 0.046 0.066 0.058 0.040 1 1 1 1 1 1
$130$ 0.052 0.058 0.054 0.056 0.050 0.058 0.052 0.028 1 1 1 1 1 1
$140$ 0.050 0.060 0.054 0.042 0.040 0.074 0.068 0.028 1 1 1 1 1 1
$150$ 0.056 0.064 0.054 0.058 0.060 0.054 0.052 0.028 1 1 1 1 1 1
Table 6: Empirical frequency of $\hat{K}$ under $H_{0}$ in the second experiment
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.972 0.024 0.004 0 0
2020 0.986 0.012 0 0.002 0
$30$ 0.980 0.018 0.002 0 0
$40$ 0.992 0.008 0 0 0
$50$ 0.990 0.010 0 0 0
$60$ 0.996 0.004 0 0 0
$70$ 0.994 0.006 0 0 0
$80$ 0.996 0.004 0 0 0
$90$ 0.996 0.004 0 0 0
$100$ 0.990 0.010 0 0 0
$110$ 0.994 0.006 0 0 0
$120$ 0.992 0.006 0.002 0 0
$130$ 0.992 0.008 0 0 0
$140$ 0.986 0.014 0 0 0
$150$ 0.998 0.002 0 0 0
Table 7: Empirical frequency of $\hat{K}$ under $H_{1}$ in the second experiment
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0 0.990 0.008 0.002 0
2020 0 0.982 0.018 0 0
$30$ 0 0.990 0.008 0 0.002
$40$ 0 0.994 0.006 0 0
$50$ 0 0.996 0.004 0 0
$60$ 0 0.996 0.004 0 0
$70$ 0 0.994 0.006 0 0
$80$ 0 0.996 0.004 0 0
$90$ 0 1 0 0 0
$100$ 0 0.998 0.002 0 0
$110$ 0 0.996 0.004 0 0
$120$ 0 1 0 0 0
$130$ 0 1 0 0 0
$140$ 0 0.988 0.012 0 0
$150$ 0 1 0 0 0
6 Discussion of other ANOVA models
In this section, we briefly discuss smooth tests for normality of other ANOVA models beyond those in Section 3. The testing approach above is also applicable to these models.
6.1 One-way fixed effects model with different group means and variances
We first consider the one-way ANOVA model with different group means and different group variances, which is given by
$Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J.$ (6.1)
Here $\{\mu_{j}\}_{j=1}^{J},\{\sigma_{j}\}_{j=1}^{J}$ represent the potentially different group means and group standard deviations, respectively. Due to the heterogeneity across groups in model (6.1), the estimation of each group mean and group variance only depends on observations in the corresponding group, in the sense that $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij},\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ for $j=1,\ldots,J$. Accordingly, we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}_{j}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To implement the smooth test, one can follow the steps in Section 3 and derive a formula like regarding the estimation effects in $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij}),k=1,\ldots,K$. Then the test statistic is constructed as a quadratic form of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$, normalized by the inverse of its asymptotic covariance matrix. It is expected that an asymptotic $\chi^{2}$ test can be obtained under mild conditions for model (6.1), which leads to a unified testing framework for one-way fixed effects models.
6.2 One-way random effects model
Now we turn to the one-way random effects model. The observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are generated from following equation:
$Y_{ij}=\mu+A_{j}+\varepsilon_{ij}=\mu+A_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (6.2)
where the $\{A_{j}\}_{j=1}^{J}$ are i.i.d. random effects independent of the random errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$, with $\mathrm{E}[A_{j}]=0$ and $\operatorname{Var}[A_{j}]=\sigma_{a}^{2}$. In this setting, we estimate $\sigma^{2}$ by
$\hat{\sigma}^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\bar{Y}_{\cdot j}\right)^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\hat{\varepsilon}_{ij}^{2},$
where $\bar{Y}_{\cdot j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\varepsilon}_{ij}=Y_{ij}-\bar{Y}_{\cdot j}$. Note that $\bar{Y}_{\cdot j}$ coincides with $\hat{\mu}_{j}$ in Section 3.2, leading to an identical expression for $\hat{\sigma}^{2}$ here as that in Section 3.2, too. Therefore, the methodology in Section 3.2 can be directly applied to test the normality of $e_{ij}$ in (6.2), since the random effects $A_{j}$ are effectively eliminated by within-group centering via $Y_{ij}-\bar{Y}_{\cdot j}$.
6.3 Two-way models
Next, we extend the discussion to two-way ANOVA models. Suppose the observations
$\{Y_{ij_{1}j_{2}}\}_{i=1,j_{1}=1,j_{2}=1}^{N_{j_{1}j_{2}},J_{1},J_{2}}$ are generated from the two-way fixed effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\left(\mu+\alpha_{j_{1}}+\beta_{j_{2}}+\gamma_{j_{1}j_{2}}\right)+\sigma e_{ij_{1}j_{2}},$ (6.3)
subject to the identifiability constraints
$\sum_{j_{1}=1}^{J_{1}}\alpha_{j_{1}}=\sum_{j_{2}=1}^{J_{2}}\beta_{j_{2}}=\sum_{j_{1}=1}^{J_{1}}\gamma_{j_{1}j_{2}}=\sum_{j_{2}=1}^{J_{2}}\gamma_{j_{1}j_{2}}=0.$
Inspired by the methodology developed in Section 3, we can construct analogous estimators for $\mu$, $\{\alpha_{j_{1}}\}_{j_{1}=1}^{J_{1}}$, $\{\beta_{j_{2}}\}_{j_{2}=2}^{J_{2}}$, $\{\gamma_{j_{1}j_{2}}\}_{j_{1}=1,j_{2}=1}^{J_{1},J_{2}}$, and $\sigma^{2}$, and derive the corresponding test statistic. The theoretical framework in Section 3 remains valid in this setting.
Similarly, for the two-way random effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\left(\mu+A_{j_{1}}+B_{j_{2}}+D_{j_{1}j_{2}}\right)+\sigma e_{ij_{1}j_{2}},$ (6.4)
where the i.i.d. zero-mean random effects $\{A_{j_{1}}\}_{j_{1}=1}^{J_{1}}$, $\{B_{j_{2}}\}_{j_{2}=1}^{J_{2}}$, $\{D_{j_{1}j_{2}}\}_{j_{1}=1,j_{2}=1}^{J_{1},J_{2}}$, and the random errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$, are mutually independent. To test the normality of $e_{ij}$ in (6.4), one can mimic the strategy used in Section 6.2 to eliminate the random effects first through within-cell centering, and then construct the test statistic based on the resulting residuals.
To conclude, the methodology proposed in Section 3 can be seamlessly extended to more general ANOVA settings. This unified framework offers a flexible and effective approach to testing for normality in a wide range of linear models.
7 Concluding remarks
In this paper, we establish Neyman’s smooth tests for assessing the normality assumption in ANOVA models. For the proposed tests in three types of one-way fixed effects models, we demonstrate the asymptotic properties and validate the finite-sample performance through extensive simulations. Our approach provides a reliable tool for validating the normality assumption in diverse ANOVA settings.
Several directions remain open for future research. First, although we have developed tests within the ANOVA framework, the true data-generating mechanism may not always be known a priori. A promising approach is to conduct preliminary structural tests to identify the most suitable model before applying the proposed procedures. Second, the independence assumption on random errors can be relaxed to enable further theoretical analysis. For example, repeated clinical trials frequently exhibit autocorrelated errors (Ma et al.,, 2012; Schober and Vetter,, 2018; Langenberg et al.,, 2022). In such cases, our smooth test framework could potentially be adapted by incorporating estimated correlation structures. Finally, for the data-driven selection rule, it is intriguing to explore the case where the upper bound $D=D(N)\to\infty$ as $N\to\infty$ from the theoretical perspective.
Appendix A Appendix
We provide the proof of our main theoretical results and some practical issues in this Appendix. The symbol $\lesssim$” means that the left side is bounded by a positive constant times the right side. The symbol $\asymp$” means that both sides are asymptotically equivalent.
Proof of Theorem 1.
We first study the estimation effects of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}-\sigma^{2}$. Note that
$\hat{\mu}-\mu=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\mu\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{\sigma}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.1)
Under $H_{0}$, $e_{ij}\sim\mathcal{N}(0,1)i.i.d$., then $\sqrt{N}(\hat{\mu}-\mu)\sim\mathcal{N}(0,\sigma^{2})$, and $\hat{\mu}-\mu=O_{p}(N^{-1/2})$ as $N\to\infty$. Besides,
$\hat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}\right)^{2}-\sigma^{2}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\left(\hat{\mu}-\mu\right)^{2}$
$=$ $\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+O_{p}\left(\frac{1}{N}\right).$ (A.2)
Under $H_{0}$, by CLT, $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\mathcal{N}(0,2\sigma^{4})$ as $N\to\infty$, then $\hat{\sigma}^{2}-\sigma^{2}=O_{p}(N^{-1/2})$.
Now we deal with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$. By second order Taylor expansion of $\pi_{k}(\hat{Z}_{ij})$ with respect to $\hat{Z}_{ij}$ at $Z_{ij}$, we obtain
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left(\hat{Z}_{ij}-Z_{ij}\right)$
$+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\tilde{Z}_{ij}\right)\left(\hat{Z}_{ij}-Z_{ij}\right)^{2},$ (A.3)
[…]
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2},$ (A.4)
where $\tilde{\varepsilon}_{ij}/\tilde{\sigma}$ an intermediate point between $^{\hat{\varepsilon}_{ij}}/{\hat{\sigma}}$ and $\varepsilon_{ij}/\sigma$, whose exact value may vary across different scenarios.
We focus on the first term in (A.4). Straightforward calculation leads to
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}$
$=$ $-\frac{\hat{\mu}-\mu}{\hat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\frac{\hat{\sigma}^{2}-\sigma^{2}}{\hat{\sigma}(\hat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
$=$ $-\frac{\hat{\mu}-\mu}{\sigma}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]-\frac{\hat{\sigma}^{2}-\sigma^{2}}{2\sigma^{2}}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]$
$+o_{p}\left(\hat{\mu}-\mu\right)+o_{p}\left(\hat{\sigma}^{2}-\sigma^{2}\right),$ (A.5)
where the last step follows due to the law of large numbers (LLN) for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)]+o_{p}(1)$ and $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)e]+o_{p}(1)$ as $N\to\infty$, as well as the consistency of $\hat{\sigma}$ to $\sigma$ from the arguments above. Under $H_{0}$, through integration by parts, we have
$\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z=c_{1k},$
and
$\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{0}^{1}\pi_{k}(z)\left[\left(\Phi^{-1}(z)\right)^{2}-1\right]\mathrm{d}z=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z=c_{2k},$
where we use the fact that $\mathrm{d}\Phi^{-1}(z)=\mathrm{d}z/\phi(\Phi^{-1}(z))$ and $\int_{0}^{1}\pi_{k}(z)\mathrm{d}z=0$. Thus, along with (A.1) and (A.2), (A.5) can be re-expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})-c_{1k}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{c_{2k}}{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.6)
It suffices to show that
$R_{N1}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\tilde{Z}_{ij}\right)\left(\hat{Z}_{ij}-Z_{ij}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.7)
and
$R_{N2}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.8)
For $R_{N1}$, by the boundedness of $\ddot{\pi}_{k}$ (in Assumption 3.2) and $\phi$,
$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left[\Phi\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}\right)-\Phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\right]^{2}$
$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\phi^{2}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}.$
By the inequality $(a-b)^{2}\leq 2(a^{2}+b^{2})$ for any $a,b\in\mathbb{R}$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\asymp$ $\left(\hat{\mu}-\mu\right)^{2}+(\hat{\sigma}^{2}-\sigma^{2})^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
$=$ $O_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{N}\right)\left(\sigma^{2}+o_{p}(1)\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
where the last step is the result of (A.1), (A.2) and LLN for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}=\sigma^{2}+o_{p}(1)$. For $R_{N2}$, we have
$|R_{N2}|$ $=\left|\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left[-\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\phi\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\right]\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}\right|$
$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left|\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right|\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$\leq\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}\right)^{1/2}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}\right)^{1/2},$
where the first “==” is due to $\dot{\phi}(x)=-x\phi(x)$, and the last $\leq$” is due to Cauchy–Schwarz inequality. Similar arguments as above yield that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}$
$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}\right]$
$\asymp$ $\left(\hat{\mu}-\mu\right)^{4}+\left(\hat{\sigma}^{2}-\sigma^{2}\right)^{4}\left(\mathrm{E}\left[e^{4}\right]+o_{p}(1)\right)=O_{p}\left(\frac{1}{N^{2}}\right),$
and
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}$ $\leq\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{{\varepsilon}_{ij}}{{\sigma}}\right)^{2}$
$=1+\left(1+o_{p}(1)\right)=O_{p}(1).$ (A.9)
Thus, $|R_{N2}|=O_{p}(N^{-1})=o_{p}(N^{-1/2})$. Hence (A.7) and (A.8) are verified. Combining (A.3), (A.4), (A.5), (A.6), (A.7) and (A.8), we complete the proof of Theorem 1. $\square$
Proof of Theorem 2.
The proof of Theorem 2 is similar to that of Theorem 1. Under $H_{1}$, (A.1) still holds, and $\sqrt{N}(\hat{\mu}-\mu)\to_{d}\mathcal{N}(0,\sigma^{2})$ since $e_{ij}\sim F\neq\Phi$. Analogously, (A.2) still holds, and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\sigma^{2}\mathcal{N}(0,\mathrm{E}[e^{4}]-1)$.
We start from (A.3). Under $H_{1}$, the arguments and results in (A.3), (A.4) and (A.5) still hold. Unlike (A.6), here in the context of $H_{1}$ the last line of (A.5) should be expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})-d_{1k}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{d_{2k}}{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.10)
where we replace $c_{1k}$ with $d_{1k}$ and $c_{2k}$ with $d_{2k}$. Besides, the remainder terms $R_{N1}$ in (A.7) and $R_{N2}$ in (A.8) are still $O_{p}(N^{-1/2})$ through similar derivation above. Combining (A.3), (A.4), (A.5), (A.7), (A.8) and (A.10), we obtain (3.5) in Theorem 2. From (3.5), we know that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}(1),$
as $N\to\infty$, which implies $\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$. Furthermore,
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)$
$=$ $\sqrt{N}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\bm{\Sigma}_{K}^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]$
$-\sqrt{N}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$
$=$ $2a^{\top}\bm{\Sigma}_{K}^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}\bm{\Sigma}_{K}^{-1}$
$\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.11)
The following part demonstrates the asymptotic normality of
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\hat{Z}_{ij})-\bm{a}_{K}\right).$
Indeed, the $k$-th component in it is
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\pi_{k}(\hat{Z}_{ij})-a_{k}\right)=\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\zeta_{k}(e_{ij})-\mathrm{E}\left[\zeta_{k}(e_{ij})\right]\right)+o_{p}(1),$
where
$\zeta_{k}(e)\equiv\pi_{k}(\Phi(e))-d_{1k}e-\frac{d_{2k}}{2}(e^{2}-1).$
Then, by CLT,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\hat{Z}_{ij})-\bm{a}_{K}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Xi}_{K}\right),$
with $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ given by
$\xi_{kl}=$ $\mathrm{E}\left[\zeta_{k}(e)\zeta_{l}(e)\right]$
$=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
Along with (A.11), (3.6) in Theorem 2 is validated. This completes the proof of Theorem 2. $\square$
Proof of Theorem 3.
The proof of Theorem 3 is similar to that of Theorems 1 and 2. We just follow the arguments in the previous proof and carefully specify the behavior of the estimation effects. Under $H_{1L}$ in (3.7), the PDF of $e$ is given by
$f(x)=(1-\delta_{N})\phi(x)+\delta_{N}q(x).$
We first examine the convergence of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}-\sigma^{2}$ as before. For $\hat{\mu}-\mu$, (A.1) still holds under $H_{1L}$, and $\sqrt{N}(\hat{\mu}-\mu)\to_{d}\mathcal{N}(0,\sigma^{2})$ as $N\to\infty$. For $\hat{\sigma}^{2}-\sigma^{2}$, under $H_{1L}$, (A.2) still holds, and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\mathcal{N}(0,2\sigma^{4})$  [^1] Here under $H_{1L}$, the asymptotic normality of $\sqrt{N}(\hat{\mu}-\mu)$ and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})$ are both derived from CLT for triangular arrays since the distribution of $e_{ij}$ changes over $N$. It can be verified that the conditions of Lindeberg–Feller CLT hold for both random sequences, see, for example, Proposition 2.27 of Van der Vaart, (2000) . since
$\mathrm{E}\left[e^{4}\right]=$ $\int_{-\infty}^{\infty}x^{4}f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}x^{4}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}x^{4}q(x)\mathrm{d}x$
$=$ $3+o(1).$
Next, we clarify the relationship of the denoted constants. Recall the definitions of the constants $d_{1k}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)f(x)\mathrm{d}x$ and $d_{2k}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)e]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)xf(x)\mathrm{d}x$. Then under $H_{1L}$, as $N\to\infty$, we have
$d_{1k}=$ $\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi^{2}\left(x\right)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)q(x)\mathrm{d}x$
$=$ $c_{1k}+o(1),$ (A.12)
$d_{2k}=$ $\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)xf(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))x\phi^{2}\left(x\right)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))x\phi\left(x\right)q(x)\mathrm{d}x$
$=$ $c_{2k}+o(1),$ (A.13)
and
$\mathrm{E}[\pi_{k}(Z)]=$ $\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)f(x)\mathrm{d}x$
$=$ $\left(1-\delta_{N}\right)\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)q(x)\mathrm{d}x$
$=$ $\delta_{N}\Delta_{k}.$ (A.14)
Now we start from (A.3) again. Under $H_{1L}$, we still have (A.3), (A.4) and (A.5). (A.6) also holds as a result of (A.12) and (A.13). The remainder terms $R_{N1}$ in (A.7) and $R_{N2}$ in (A.8) are still $O_{p}(N^{-1/2})$. Combining (A.3), (A.4), (A.5), (A.6), (A.7), (A.8) and (A.14), we obtain (3.8) in Theorem 3.
It suffices to derive the asymptotic normality of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$. Under $H_{1L}$, straightforward calculation yields that
$\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]=$ $\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{l}(\Phi(x))f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{k}(\Phi(x))\phi(x)\mathrm{d}x$
$+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{k}(\Phi(x))q(x)\mathrm{d}x$
$=$ $\delta_{kl}+o(1),$
$d_{3k}=$ $\mathrm{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xf(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xq(x)\mathrm{d}x$
$=$ $c_{1k}+o(1),$
$d_{4k}=$ $\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}q(x)\mathrm{d}x$
$=$ $c_{2k}+o(1),$
and
$\mathrm{E}\left[e^{3}\right]=$ $\int_{-\infty}^{\infty}x^{3}f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}x^{3}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}x^{3}q(x)\mathrm{d}x$
$=$ $o(1).$
Thus, by (3.8) and CLT,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
$\to_{d}\mathcal{N}_{K}\left(\bm{\Delta}_{K},\bm{\Sigma}_{K}\right),$
since the asymptotic covariance matrix is still $\bm{\Sigma}_{K}$ from the calculation above. Besides,
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
by definition of the non-central $\chi^{2}$ distribution. This completes the proof of Theorem 3. $\square$
Proof of Theorem 4, 5 and 6.
The proof is similar to that of Theorems 1, 2, and 3. Note that in the case of different means and same variance we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ without distinction of the null and alternatives. Our goal is to prove that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]e_{ij}-\frac{\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.15)
We first focus on the estimation effects of $\hat{\mu}_{j}-\mu_{j}$ and $\hat{\sigma}^{2}-\sigma^{2}$. For $\hat{\mu}_{j}-\mu_{j}$, by CLT, we have $\sqrt{N_{j}}(\hat{\mu}_{j}-\mu_{j})\to_{d}\mathcal{N}(0,\sigma^{2})$ for each $1\leq j\leq J$ as $\min\{N_{1},\ldots,N_{J}\}\to\infty$. Besides, since
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\mathrm{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\right]=\frac{J}{N}\sigma^{2},$ (A.16)
we have $N^{-1}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}=O_{p}(J/N)=o_{p}(N^{-1/2})$ from $J=o(N^{1/2})$. Analogously, by
$\mathrm{E}\left[\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{1}{N_{j}^{4}}\left[3N_{j}(N_{j}-1)+N_{j}\mathrm{E}\left[e^{4}\right]\right]=\frac{3}{N_{j}^{2}}+\frac{\mathrm{E}\left[e^{4}\right]-3}{N_{j}^{3}},$
we obtain
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}\right]$
$=$ $\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\mathrm{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\left(\frac{3}{N_{j}}+\frac{\mathrm{E}\left[e^{4}\right]-3}{N_{j}^{2}}\right),$ (A.17)
which implies $N^{-1}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}=N^{-1}O_{p}(\sum_{j=1}^{J}N_{j}^{-1})=o_{p}(N^{-1})$ since $\sum_{j=1}^{J}N_{j}^{-1}=o_{p}(1)$. For $\hat{\sigma}^{2}-\sigma^{2}$, we have
$\hat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}-\sigma^{2}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}$
$=$ $\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
with the last step following (A.16). Note that the main term in the last line is the same as (A.2).
Next, we will show that
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$=$ $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.18)
Indeed, since
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)e_{i_{2}j},$
and
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}=\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}e_{i_{2}j},$
then we have
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{ij}$
$+\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
$\equiv$ $A_{N1}+A_{N2}.$
For $A_{N1}$, we have
$\mathrm{E}[A_{N1}]=\frac{J}{N}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right],$
and
$\operatorname{Var}[A_{N1}]=\operatorname{Var}\left[\left(\dot{\pi}_{k}(Z)\phi\left(e\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e\right]\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}},$
then $A_{N1}=o_{p}(N^{-1/2})$ since $J=o(N^{1/2})$ as $N\to\infty$. For $A_{N2}$, we have $\mathrm{E}[A_{N2}]=0$, and
$\operatorname{Var}[A_{N2}]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\operatorname{Var}\left[\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\mathrm{E}\left[\left(\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right)^{2}\right]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{N_{j}(N_{j}-1)}{N_{j}^{2}}\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathrm{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)$
$=$ $\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathrm{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)\frac{1}{N^{2}}\left(J-\sum_{j=1}^{J}\frac{1}{N_{j}}\right),$
then $A_{N2}=o_{p}(N^{-1/2})$. Hence, (A.18) is verified.
Now we start from (A.3). In the case of different means and same variance, (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-e_{ij}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}$
$=$ $-\frac{1}{\hat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$-\frac{\hat{\sigma}^{2}-\sigma^{2}}{\hat{\sigma}(\hat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
$=$ $-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{1}{2\sigma^{2}}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\left(\hat{\sigma}^{2}-\sigma^{2}\right)$
$+o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}\right)+o_{p}\left(\hat{\sigma}^{2}-\sigma^{2}\right)$
$=$ $-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
$-\frac{1}{2N}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.19)
where we apply (A.2), (A.18), LLN for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$, and CLT for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$ as $N\to\infty$.
It suffices to show $R_{N1}$ (in (A.7)) and $R_{N2}$ (in (A.8)) are $o_{p}(N^{-1/2})$ in order to obtain (A.15). The technique is similar to that in the proof of Theorem 1. For $R_{N1}$, we have
$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\lesssim\frac{1}{N\sigma^{2}}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}+\left(\hat{\sigma}^{2}-\sigma^{2}\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
$\asymp O_{p}\left(\frac{J}{N}\right)+O_{p}\left(\frac{1}{N}\right)\left(\sigma^{2}+o_{p}(1)\right)=O_{p}\left(\frac{J}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
by (A.16) and the assumption $J=o(N^{1/2})$. For $R_{N2}$, we first have
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}$
$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}\right]$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}+(\hat{\sigma}^{2}-\sigma^{2})^{4}\left(\mathrm{E}\left[e^{4}\right]+o_{p}(1)\right)$
$\asymp$ $o_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{N^{2}}\right)=o_{p}\left(\frac{1}{N}\right),$
by (A.17). Along with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\tilde{\varepsilon}_{ij}^{2}/\tilde{\sigma}^{2}=O_{p}(1)$ from (A.9), we obtain
$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
Thus, (A.15) is verified. The form of (A.15) can be further refined under $H_{0}$, $H_{1}$ and $H_{1L}$ respectively. For example, under $H_{0}$, we can write $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ and $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ in (A.15), and then obtain the results of Theorem 4. $\square$
Proof of Theorem 7.
The technique is still the same as the previous ones. Recall that in the context of same mean and different variances, we denote $p_{j}=N_{j}/N$, $q_{j}=JN_{j}/N$ for $1\leq j\leq J$, and we further assume that as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$.
We first tackle the estimation effects of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}$. For $\hat{\mu}-\mu$, note that $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ here. Then,
$\hat{\mu}-\mu=\frac{1}{J}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{1}{J}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.20)
From (A.20) we obtain $\mathrm{E}[\hat{\mu}-\mu]=0$, and $\operatorname{Var}[\hat{\mu}-\mu]=J^{-2}\sum_{j=1}^{J}N_{j}^{-1}\sigma_{j}^{2}\asymp J^{-2}\sum_{j=1}^{J}N_{j}^{-1}=N^{-1}J^{-1}\sum_{j=1}^{J}q_{j}^{-1}=O_{p}(N^{-1})$. Then $\hat{\mu}-\mu=O_{p}(N^{-1/2})$. For $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}$, we have
$\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}$ $=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}\right)^{2}-\sigma_{j}^{2}\right)=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma_{j}^{2}\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}$ (A.21)
$=\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}$
$=O_{p}\left(\frac{1}{\sqrt{N_{j}}}\right)+O_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{\sqrt{N_{j}N}}\right)=O_{p}\left(\frac{1}{\sqrt{N_{j}}}\right),$
by CLT for $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}e_{ij}\to_{d}\mathcal{N}(0,1)$ and $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)\to_{d}\mathcal{N}(0,\mathrm{E}[e^{4}]-1)$ as $N_{j}\to\infty$ and (A.20).
We next prove some preliminary results such that
$\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}+O_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.22)
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
$=$ $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.23)
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.24)
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}=O_{p}(1),$ (A.25)
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$. For (A.22), straightforward calculation leads to
$\mathrm{E}\left[\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right]=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}},$
and
$\operatorname{Var}\left[\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right]=\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p^{2}_{j}}{N_{j}\sigma^{2}_{j}}$
$=$ $\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{N_{j}}{\sigma^{2}_{j}}=O\left(\frac{1}{N}\right).$
Thus, (A.22) is verified. The derivation of (A.23) is analogous to that of (A.18) if we first rewrite $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ on the right side as
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}(e_{i_{2}j}^{2}-1),$
and so is (A.24). (A.25) is obtained by
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\}\right]=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)e]\frac{1}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}=O(1),$
and
$\operatorname{Var}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\right]=O\left(\frac{1}{N}\right).$
Now we start from (A.3). In the case of the same mean and different variances, (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(\frac{\varepsilon_{ij}}{\sigma_{j}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}$
$-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}.$
$\equiv$ $D_{N1}-D_{N2}.$
For $D_{N1}$, under $H_{0}$, we have
$D_{N1}$ $=-\frac{\hat{\mu}-\mu}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\left(1+o_{p}(1)\right)$
$=-\frac{1}{N}\left(\sum_{j=1}^{J}\frac{N}{JN_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}\right)\left(\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right)\left(1+o_{p}(1)\right)$
$=-\frac{1}{N}\left(\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\sigma_{j}e_{ij}}{q_{j}}\right)\left(c_{1k}\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}+O_{p}\left(\frac{1}{\sqrt{N}}\right)\right)\left(1+o_{p}(1)\right)$
$=-\frac{c_{1k}}{N}\left[\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\sigma_{j}e_{ij}}{q_{j}}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.26)
where the second last line is followed by (A.22), and $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ under $H_{0}$. The last line is followed by $N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=o_{p}(1)$ using LLN. For $D_{N2}$, under $H_{0}$, we have
$D_{N2}=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left\{\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}\right$.
$\left.-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{2}}{N_{j}\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
$+\left(\hat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
$-2\left(\hat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
$=$ $\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{N_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}\left(1+o_{p}(1)\right)$
$+\left(\hat{\mu}-\mu\right)^{2}\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left(1+o_{p}(1)\right)$
$-\left(\hat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}\sigma_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
$=$ $\frac{c_{2k}}{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.27)
following (A.20), (A.21), (A.23), (A.24) and (A.25), as well as $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ under $H_{0}$.
It suffices to show $R_{N1}$ (in (A.7)) and $R_{N2}$ (in (A.8)) are $o_{p}(N^{-1/2})$ in this case. For $R_{N1}$, referring to the arguments in the proof of Theorem 4, here we have
$|R_{N1}|\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}\right)^{2}\equiv R_{N1}^{A}+R_{N1}^{B}.$
For $R_{N1}^{A}$, we obtain
$R_{N1}^{A}=\frac{(\hat{\mu}-\mu)^{2}}{N}\sum_{j=1}^{J}\frac{N_{j}}{\hat{\sigma}_{j}^{2}}\asymp\frac{1}{N}O_{p}\left(\frac{1}{N}\right)O_{p}(N)=O_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
by (A.20) and
$\sum_{j=1}^{J}\frac{N_{j}}{\hat{\sigma}_{j}^{2}}=\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}(1+o_{p}(1))=O_{p}(N).$
For $R_{N1}^{B}$, we have
$R_{N1}^{B}\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\left(\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)+\left(\hat{\mu}-\mu\right)^{4}\left(\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$+4\left(\hat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$\equiv$ $R_{N1}^{B,1}+\left(\hat{\mu}-\mu\right)^{4}R_{N1}^{B,2}+4\left(\hat{\mu}-\mu\right)^{2}R_{N1}^{B,3}.$
For $R_{N1}^{B,1}$, we rewrite it as
$R_{N1}^{B,1}=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\sum_{i_{3}=1}^{N_{j}}\left(e_{i_{1}j}^{2}-1\right)\left(e_{i_{2}j}^{2}-1\right)e_{i_{3}j}^{2}.$
Note that $\mathrm{E}[(e_{i_{1}j}^{2}-1)(e_{i_{2}j}^{2}-1)e_{i_{3}j}^{2}]=\mathrm{E}[(e^{2}-1)^{2}e^{2}]\mathbbm{1}(i_{1}=i_{2}=i_{3})+\mathrm{E}[(e^{2}-1)^{2}]\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$. Then,
$\mathrm{E}\left[R_{N1}^{B,1}\right]$ $=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(N_{j}\mathrm{E}\left[(e^{2}-1)^{2}e^{2}\right]+N_{j}(N_{j}-1)\mathrm{E}\left[(e^{2}-1)^{2}\right]\right)$
$=O\left(\frac{J}{N}\right),$
which implies $R_{N1}^{B,1}=o_{p}(N^{-1/2})$. Similarly, for $R_{N1}^{B,3}$, by $\mathrm{E}[e_{i_{1}j}e_{i_{2}j}e_{i_{3}j}^{2}]=\mathrm{E}[e^{4}]\mathbbm{1}(i_{1}=i_{2})+\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$, we have
$\mathrm{E}\left[R_{N1}^{B,3}\right]=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(N_{j}\mathrm{E}\left[e^{4}\right]+N_{j}(N_{j}-1)\right)=O\left(\frac{J}{N}\right),$
which implies $R_{N1}^{B,3}=o_{p}(N^{-1/2})$. For $R_{N1}^{B,2}$, straightforward calculation leads to
$\mathrm{E}\left[R_{N1}^{B,2}\right]=\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right]=\frac{1}{N}\sum_{j=1}^{J}N_{j}\sigma_{j}^{2}=O(1),$
then $R_{N1}^{B,2}=O_{p}(1)$. Along with (A.20), we obtain $R_{N1}^{B}=o_{p}(N^{-1/2})$. Hence $R_{N1}=o_{p}(N^{-1/2})$. On the other hand, for $R_{N2}$, we first note that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}\right)^{4}$
$\asymp\left(\hat{\mu}-\mu\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
$\asymp O_{p}\left(\frac{1}{N^{2}}\right)+o_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{N}\right),$
since
$\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
$\leq$ $\left\{\max_{1\leq j\leq J}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\right\}\left\{\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right\}$
$=$ $o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right)$
$=$ $o_{p}\left(\frac{1}{\sqrt{N}}\right),$
where we use $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}=o_{p}(1)$ from (A.21), and the last step can be obtained through similar arguments above. Besides, $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\tilde{\varepsilon}_{ij}^{2}/\tilde{\sigma}_{j}^{2}=O_{p}(1)$ can be derived from the procedure in (A.9) analogously. Then we have
$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}_{j}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
Therefore, (3.11) is validated. This completes the proof of Theorem 7. $\square$
Proof of Theorem 8 and 9.
The proof of Theorem 8 is analogous to that of Theorem 2. Under $H_{1}$, in order to obtain (3.14), we just repeat the procedure in the proof of Theorem 7, and replace $c_{1k}$ with $d_{1k}$ in (A.26), $c_{2k}$ with $d_{2k}$ in (A.27). From (3.5) we know that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}(1).$
Thus,
$\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)\to_{p}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}.$
Besides,
$\sqrt{N}\left(\tilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)$
$=$ $\sqrt{N}\left\{\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]\right$.
$\left.\qquad-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right\}$
$=$ $2a^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}$
$\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.28)
Under $H_{1}$, by CLT, we have
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right)-a_{1},\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)-a_{K}\right]^{\top}$
$\to_{d}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{W}_{j},$
where $\bm{W}_{1},\ldots,\bm{W}_{J}$ are independently distributed as
$\bm{W}_{j}\sim\mathcal{N}_{K}\left(0,\bm{\Lambda}_{K}^{(j)}\right),$
with $\bm{\Lambda}_{K}^{(j)}=\left(\lambda_{kl}^{(j)}\right)_{K\times K}$ given by
$\lambda_{kl}^{(j)}=$ $\mathrm{E}\left\{\left(\pi_{k}(Z)-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2k}}{2}\left(e^{2}-1\right)\right\}$
$\quad\left\{\left(\pi_{l}(Z)-\mathrm{E}\left[\pi_{l}(Z)\right]\right)-d_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2l}}{2}\left(e^{2}-1\right)\right\}$
$=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathrm{E}\left[e^{3}\right]$
$+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
Therefore,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Lambda}_{K}^{(j)}\right),$
and the asymptotic distribution of $\tilde{\Psi}_{K}^{2}$ in (3.15) can be derived from (A.28) through similar arguments above. This completes the proof of Theorem 8.
The proof of Theorem 9 can be referred to that of Theorem 3 due to the similar arguments and techniques used. To obtain the results in Theorem 9, following the proof of Theorem 8, it suffices to show that under $H_{1L}$, $\mathrm{E}[\pi_{k}(Z)]=\delta_{N}\Delta_{k}$, $\mathrm{E}[\pi_{k}(Z)\pi_{l}(Z)]=\delta_{kl}+o(1)$, $d_{1k}=c_{1k}+o(1)$, $d_{2k}=c_{2k}+o(1)$, $d_{3k}=c_{1k}+o(1)$, $d_{4k}=c_{2k}+o(1)$, $\mathrm{E}[e^{3}]=o(1)$, and $\mathrm{E}[e^{4}]=3+o(1)$, which can be found in the proof of Theorem 3. $\square$
Proof of Theorem 10.
We focus on the first case in Theorem 10, and the latter two cases can be verified analogously. We primarily follow well-established steps in the literature, for example, the proof of Theorem 2 in Ducharme and Lafaye de Micheaux, (2004), Lemma 1 in Kraus, (2007), and Propositions 6 and 7 in Duchesne et al., (2016). Note that $\{\hat{K}=k\}\subseteq\{N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\}$ for $k^{\prime}\neq k$, then we have
$\mathrm{P}\left(\hat{K}=k\right)\leq\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\right),\quad k^{\prime}\neq k.$ (A.29)
Under $H_{0}$, in order to show $\mathrm{P}(\hat{K}=1)\to 1$ as $N\to\infty$, we first show the contrary such that $\sum_{k=2}^{D}\mathrm{P}(\hat{K}=k)\to 0$. In fact, from (A.29) we have
$\sum_{k=2}^{D}\mathrm{P}\left(\hat{K}=k\right)$
$\leq$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{1}-\log N\right)$
$\leq$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq-\log N\right)$
$=$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N\right).$
From Corollary 1 we know that under $H_{0}$, $N\hat{\Psi}^{2}_{k}\to_{d}\chi^{2}_{k}$ for each $k=1,\ldots,D$. Thus for $k=2,\ldots,D$, $\mathrm{P}(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N)\to 0$ as $N\to\infty$. Therefore,
$\sum_{k=2}^{D}\mathrm{P}\left(\hat{K}=k\right)\leq\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N\right)\to 0,$
as $N\to\infty$, then $\mathrm{P}(\hat{K}=1)=1-\sum_{k=2}^{D}\mathrm{P}(\hat{K}=k)\to 1$, and $N\hat{\Psi}^{2}_{\hat{K}}\to_{d}\chi^{2}_{1}$ follows.
Under $H_{1}^{\prime}$, (A.29) yields that
$\mathrm{P}\left(\hat{K}=k\right)$ $\leq\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{K_{0}}-K_{0}\log N\right)$
$=\mathrm{P}\left(\hat{\Psi}^{2}_{k}-k\frac{\log N}{N}\geq\hat{\Psi}^{2}_{K_{0}}-K_{0}\frac{\log N}{N}\right).$
From Theorem 2 we know that as $N\to\infty$, $\hat{\Psi}^{2}_{k}\to_{p}0$ for $k=1,\ldots,K_{0}-1$, $\hat{\Psi}^{2}_{K_{0}}\to_{p}\bm{a}_{K_{0}}^{\top}\bm{\Sigma}_{K_{0}}^{-1}\bm{a}_{K_{0}}>0$ and $\log N/N\to 0$. Thus $\mathrm{P}(\hat{K}=k)\to 0$ for $k=1,\ldots,K_{0}-1$, and
$\lim_{N\to\infty}\mathrm{P}\left(\hat{K}\geq K_{0}\right)=1-\lim_{N\to\infty}\sum_{k=1}^{K_{0}-1}\mathrm{P}\left(\hat{K}=k\right)=1.$
Then, for any $x\in\mathbb{R}$, as $N\to\infty$,
$\mathrm{P}\left(N\hat{\Psi}^{2}_{\hat{K}}\leq x\right)$ $=\sum_{k=1}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\leq x,\hat{K}=k\right)$
$\leq\sum_{k=1}^{K_{0}-1}\mathrm{P}\left(\hat{K}=k\right)+\sum_{k=K_{0}}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\leq x\right)\to 0,$
since $\hat{\Psi}^{2}_{k}\to_{p}\bm{a}_{k}^{\top}\bm{\Sigma}_{k}^{-1}\bm{a}_{k}>0$ for $k=K_{0},\ldots,D$. Therefore, for any $x\in\mathbb{R}$,
$\lim_{N\to\infty}\mathrm{P}\left(N\hat{\Psi}^{2}_{\hat{K}}\leq x\right)=0,$
which means the test is consistent against the alternatives given by (4.2). $\square$
Intuition behind the revised limiting null distribution+ (4.3).
We mainly follow the arguments in Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007), and provide a heuristic derivation of the approximation $H(x)$ for the first case since the latter two cases can be derived analogously. Note that
$\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x\right)\\ =\mathrm{P}\left(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1\right)+\mathrm{P}\left(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2\right)+\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3\right).$ (A.30)
The third term on the right-hand side of (A.30), $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3)$, can be neglected under $H_{0}$ (Kallenberg and Ledwina, 1995b, ). The event $\{\hat{K}=1\}$ is approximated by $\{N\hat{\Psi}^{2}_{1}-\log N\geq N\hat{\Psi}^{2}_{2}-2\log N\}=\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\hat{K}=2\}$ is approximated by $\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})>\log N\}$.
We first investigate the asymptotic distribution of $(N\hat{\Psi}^{2}_{1},N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1}))^{\top}$. The limiting distributions of $N\hat{\Psi}^{2}_{1},N\hat{\Psi}^{2}_{2}$ are functions of a bivariate normal vector $(R_{1},R_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\bm{\Sigma}_{2})$. We generally denote elements of $\bm{\Sigma}_{2}$ as $\begin{pmatrix}a&b\\ b&c\end{pmatrix}$ and $\rho=b/\sqrt{ac}$. The distribution of $(R_{1},R_{2})^{\top}$ can be constructed through two independent standard normal variables $G_{1},G_{2}$ if we denote $\tilde{R}_{1}=\sqrt{a}(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})$ and $\tilde{R}_{2}=\sqrt{c}G_{2}$, then $(\tilde{R}_{1},\tilde{R}_{2})^{\top}\overset{d}{=}(R_{1},R_{2})^{\top}$. Thus, $N\hat{\Psi}^{2}_{1}\to_{d}R_{1}^{2}/a\overset{d}{=}\tilde{R}_{1}^{2}/a=(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})^{2}$. For $N\hat{\Psi}_{2}^{2}$, straightforward computations yield that
$N\hat{\Psi}_{2}^{2}$ $\to_{d}\left(\tilde{R}_{1},\tilde{R}_{2}\right)\bm{\Sigma}_{2}^{-1}\left(\tilde{R}_{1},\tilde{R}_{2}\right)^{\top}$
$=\frac{1}{ac-b^{2}}\left(\tilde{R}_{1},\tilde{R}_{2}\right)\begin{pmatrix}c&-b\\ -b&a\end{pmatrix}\left(\tilde{R}_{1},\tilde{R}_{2}\right)^{\top}$
$=\frac{1}{ac-b^{2}}\left(c\tilde{R}_{1}^{2}-2b\tilde{R}_{1}\tilde{R}_{2}+a\tilde{R}_{2}^{2}\right)$
$=G_{1}^{2}+G_{2}^{2}.$
Then $N(\hat{\Psi}_{2}^{2}-\hat{\Psi}_{1}^{2})\to_{d}(\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{2}$. Since $(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2},\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\mathbf{I}_{2})$, we obtain that $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ are both asymptotically $\chi_{1}^{2}$ distributed and are asymptotically independent.
Now we calculate the approximation $H(x)$. We will treat $H(x)$ separately for $x\leq\log N$, $\log N<x<2\log N$ and $x\geq 2\log N$. For $x\leq\log N$, note that
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,\hat{K}=2\right)\approx\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}>\log N\right)=0,$
by $N\hat{\Psi}^{2}_{1}\geq 0$. Thus, we have the approximation
$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)\approx\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]\equiv H(x),\quad x\leq\log N,$
from the asymptotic $\chi_{1}^{2}$ distribution and asymptotic independence of $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$. For $x\geq 2\log N$, we have
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,\hat{K}=2\right)$
$\approx$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$=$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)-\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right).$
As $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ is approximately $\chi_{1}^{2}$ distributed, we obtain
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)\approx 2\left(1-\Phi\left(\sqrt{\log N}\right)\right)\approx 2\frac{\phi\left(\sqrt{\log N}\right)}{\sqrt{\log N}}=\sqrt{\frac{2}{\pi N\log N}},$
by the fact that $t\phi(t)/(t+1)\leq 1-\Phi(t)\leq\phi(t)/t$ for $t\to\infty$. Besides,
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$\leq$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x\right)$
$\leq$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>2\log N\right)$
$\approx$ $\exp\left\{-\frac{1}{2}\times 2\log N\right\}$
$=$ $\frac{1}{N},$
where we use the distribution function of the limiting $\chi^{2}_{2}$ for $N\hat{\Psi}^{2}_{2}$. Hence $\mathrm{P}(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N)$ converges to 0 much slower than $\mathrm{P}(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N)$, and thus the latter can be neglected. Therefore, for $x\geq 2\log N$, we have the approximation
$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)+\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$\approx$ $\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]+2\left[1-\Phi\left(\sqrt{\log N}\right)\right]$
$\equiv$ $H(x),\quad x\geq 2\log N.$
For $\log N<x<2\log N$, Kallenberg and Ledwina, 1995b suggested the linearization as follows
$H(x)=H(\log N)+\frac{x-\log N}{\log N}[H(2\log N)-H(\log N)],\quad\log N<x<2\log N.$
This completes the derivation of the approximation (4.3) for the null distribution of the test statistic $N\hat{\Psi}^{2}_{\hat{K}}$. $\square$
Practical issues in simulation studies.
In our simulation study, we adopt the orthonormal Legendre polynomials on $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$. The properties of orthonormal Legendre polynomials lead to the following expressions for constants $c_{1k}$ and $c_{2k}$:
$c_{1k}$ $=\frac{\sqrt{2k+1}}{2^{k}}\,\sum_{j=0}^{[k/2]}(-1)^{j}\binom{k}{j}\binom{2(k-j)}{k}\int_{0}^{1}(2z-1)^{k-2j}\,\Phi^{-1}(z)\,\mathrm{d}z$
$=\sqrt{2k+1}\,\sum_{j=0}^{[k/2]}\left(\frac{-1}{4}\right)^{j}\binom{k}{j}\binom{2(k-j)}{k}\,\int_{-1/2}^{1/2}x^{k-2j}\,\Phi^{-1}\left(x+\frac{1}{2}\right)\,\mathrm{d}z,$
and
$c_{2k}=\sqrt{2k+1}\,\sum_{j=0}^{[k/2]}\left(\frac{-1}{4}\right)^{j}\binom{k}{j}\binom{2(k-j)}{k}\,\int_{-1/2}^{1/2}x^{k-2j}\,\left[\Phi^{-1}\left(x+\frac{1}{2}\right)\right]^{2}\,\mathrm{d}z.$
Note that $c_{2k}=0$ when $k$ is odd and $c_{1k}=0$ when $k$ is even. Therefore, in practice, we only compute the non-zero elements based on the above two formulae for $k=1,\ldots,K$ with $K\leq 5$.

### v3

@@
 Abstract
-The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models, yet it is rarely formally tested in practice.
-In this paper, we propose Neyman’s smooth tests for assessing the normality assumption across various types of ANOVA models.
-The proposed test statistics are constructed based on the Gaussian probability integral transformation of ANOVA residuals.
-Under the null hypothesis of normality, the test statistics are asymptotically Chi-square distributed, with degrees of freedom determined by the dimension of the smooth model (the number of orthonormal functions).
-A data-driven selection of the model dimension using a modified Schwarz’s criterion is also discussed.
-Simulation studies demonstrate the effectiveness of our proposed method.
+The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models, yet it is seldom subjected to formal testing in practice.
+In this paper, we develop Neyman’s smooth tests for assessing normality in a broad class of ANOVA models.
+The proposed test statistics are constructed via the Gaussian probability integral transformation of ANOVA residuals and are shown to follow an asymptotic Chi-square distribution under the null hypothesis, with degrees of freedom determined by the dimension of the smooth model.
+We further propose a data-driven selection of the model dimension based on a modified Schwarz’s criterion.
+Monte Carlo simulations demonstrate that the tests maintain the nominal size and achieve high power against a wide range of alternatives.
+Our framework thus provides a systematic and effective tool for formally validating the normality assumption in ANOVA models.
 Keywords: ANOVA; Estimation effect; Normality; Schwarz’s selection rule; Smooth test.
 1 Introduction
-Analysis of Variance (ANOVA) is an important and widely used method in exploratory and confirmatory data analysis (Gelman,, 2005), particularly for comparing group means and assessing the significance of factors in experimental designs.
-The theory of ANOVA models has been well established and developed in the literature, see, for example, Scheffé, (1959), Miller Jr, (1997), Wickens and Keppel, (2004), Dean et al., (2017), Hirotsu, (2017), Montgomery, (2017).
-A standard assumption in ANOVA models is that the random errors are normally distributed.
-Nonnormality may make the normal-based inference invalid.
-In terms of estimation, departures from normality generally undermine the validity of the normal-theory variance formulae for point estimates of variance components, because these formulae rely on theoretical results that represent the mean squares of random effects as constant multiples of Chi-square variables (Scheffé,, 1959).
-From the view of testing, the $F$-test employed in ANOVA is sensitive to nonnormality if the group sizes are unequal (Ali and Sharma,, 1996).
-The effects of nonnormality on size and power of the $F$-test are also extensively investigated in details (Pearson,, 1931; Gayen,, 1950; David and Johnson, 1951a, ; David and Johnson, 1951b, ; Srivastava,, 1959; Atiqullah,, 1962; Tiku,, 1964, 1971; Donaldson,, 1968; Janic-Wróblewska and Ledwina,, 2000).
-Therefore, it is crucial to evaluate the normality of ANOVA models.
-While a broad range of normality tests have been proposed for observed data and linear models, comparatively little effort has been devoted to systematically checking the normality assumption in ANOVA.
-A direct approach to normality diagnosis is to plot a normal probability plot of ANOVA residuals (Dean et al.,, 2017; Hirotsu,, 2017), which is intuitive but lacks a theoretical guarantee.
-To our knowledge, formal hypothesis testing procedures for assessing normality in ANOVA appear to be relatively scarce in the existing literature.
-Existing methods are merely a simple application of classical normality tests to ANOVA residuals, see, for instance, Bonett and Woodward, (1990), Bonett and Seier, (2002) and Hwang and Wei, (2006).
-Although applicable in practice, these methods do not precisely provide a rigorous assessment of normality within the ANOVA framework, nor do they account for the specific structure and constraints inherent in ANOVA models.
-In this article, following the spirit of Neyman, (1937), we propose a unified smooth test framework for assessing the normality assumption in various types of ANOVA models.
-Neyman’s smooth test has gained popularity across various scientific disciplines due to its well-known advantages in both theoretical and practical aspects.
-It has proven to be powerful in detecting against a broad class of alternatives; see, e.g., Bera and Ghosh, (2002), Thas, (2010), and Lehmann and Romano, (2022).
-Specifically, we first re-express the problem of testing for normality as testing for uniformity via the Gaussian probability integral transformation (PIT), and then construct the test statistics through the Gaussian PIT of ANOVA residuals.
-Under mild conditions for each model, the proposed test statistics are asymptotically Chi-square distributed under the null.
-Furthermore, we analyze the power properties of the tests under fixed and local alternatives.
-A modified Schwarz’s selection rule is proposed to determine the order of tests.
+Analysis of Variance (ANOVA) is a fundamental and widely used tool in both exploratory and confirmatory data analysis (Gelman,, 2005), particularly for comparing group means and assessing the significance of factors in experimental designs.
+The theory of ANOVA has been well established in the literature; see, for example, Scheffé, (1959), Miller Jr, (1997), Wickens and Keppel, (2004), Dean et al., (2017), Hirotsu, (2017), and Montgomery, (2017).
+A standard assumption in ANOVA is that the random errors are normally distributed.
+Violations of this assumption may invalidate normal-based inference.
+From an estimation perspective, departures from normality undermine the validity of variance component estimators, since the classical variance formulae rely on mean squares of random effects being constant multiples of Chi-square variables (Scheffé,, 1959).
+From a testing perspective, the $F$-test in ANOVA is sensitive to nonnormality, especially when group sizes are unequal (Ali and Sharma,, 1996).
+The impact of nonnormality on the size and power of the $F$-test has been extensively studied (Pearson,, 1931; Gayen,, 1950; David and Johnson, 1951a, ; David and Johnson, 1951b, ; Srivastava,, 1959; Atiqullah,, 1962; Tiku,, 1964; Donaldson,, 1968; Tiku,, 1971; Janic-Wróblewska and Ledwina,, 2000).
+These considerations highlight the importance of rigorously evaluating the normality assumption in ANOVA models.
+Although a variety of normality tests have been proposed for observed data and linear models, relatively little attention has been paid to systematically assessing normality in ANOVA.
+A straightforward diagnostic approach is to use the normal probability plots of ANOVA residuals (Dean et al.,, 2017; Hirotsu,, 2017), which are intuitive but lack theoretical justification.
+To our knowledge, formal hypothesis testing procedures designed explicitly for ANOVA are scarce in the literature.
+Most existing methods simply apply classical normality tests to ANOVA residuals, see, for example, Bonett and Woodward, (1990), Bonett and Seier, (2002), and Hwang and Wei, (2006).
+While applicable in practice, such approaches do not provide a rigorous assessment of normality within the ANOVA framework and fail to account for the structural constraints inherent in ANOVA models.
+In this article, following the spirit of Neyman, (1937), we propose a unified smooth test framework for assessing normality in various types of ANOVA models.
+Neyman’s smooth test is widely used in diverse scientific fields due to its theoretical soundness and practical effectiveness, with proven power against a broad class of alternatives; see, e.g., Bera and Ghosh, (2002), Thas, (2010), and Lehmann and Romano, (2022).
+Specifically, we reformulate the normality testing problem as one of uniformity via the Gaussian probability integral transform (PIT), and construct test statistics based on the Gaussian PIT of ANOVA residuals.
+Under mild conditions, the proposed statistics are asymptotically Chi-square distributed under the null hypothesis.
+We further analyze the power properties under both fixed and local alternatives.
+A modified Schwarz’s selection rule is proposed to determine the direction of tests.
 Monte Carlo experiments demonstrate the good performance of our methodology in finite samples.
-The advantages and contributions of our proposed tests are as follows.
-First, the proposed asymptotic Chi-square tests are simple to implement, in the sense that the critical values can be obtained directly from the limiting Chi-square distributions without bootstrapping.
-Second, we systematically investigate the properties of test statistics for three types of one-way fixed effects models with potentially diverging numbers of groups in a unified asymptotic framework.
-The parameter estimation effects in each model are well interpreted through our analysis procedure.
-The heterogeneity in group means and group variances does contribute to the specific form of test statistics and to the corresponding regularity conditions, but does not alter the convergence rate, which is only related to the total sample sizes under both the null and the alternatives.
-Lastly, we clarify the motivation and theoretical properties of data-driven smooth tests under ANOVA settings, and provide a revised limiting null distribution of the data-driven test statistics in finite samples.
-The rest of this paper is organized as follows.
-Section 2 presents the general testing framework.
-Section 3 establishes the theoretical results for smooth tests in three types of ANOVA models.
-In Section 4, a data-driven approach is proposed to determine the order of tests.
-Simulation results in Section 5 demonstrate good finite-sample performance of our proposed methodology.
-Some extensions are discussed in Section 6.
+The main contributions of this paper are threefold.
+First, the proposed asymptotic Chi-square tests are simple to implement: critical values are obtained directly from the limiting distributions without resorting to resampling.
+Second, we systematically investigate the properties of the test statistics in three types of one-way fixed effects models, potentially with a diverging number of groups, within a unified asymptotic framework.
+Our analysis clarifies how parameter estimation affects the test statistics: heterogeneity in group means and variances alters their explicit form and the corresponding regularity conditions, but the convergence rate depends only on the total sample size under both null and alternative hypotheses.
+Third, we provide theoretical justification for data-driven smooth tests in the ANOVA setting, including a revised limiting null distribution for the corresponding statistics in finite samples.
+The remainder of this paper is organized as follows.
+Section 2 introduces the general testing framework.
+Section 3 establishes theoretical results for three types of ANOVA models.
+Section 4 proposes a data-driven testing procedure.
+Section 5 reports simulation results.
+Section 6 discusses further extensions.
 Section 7 concludes.
-The proof of theoretical results and some practical issues are gathered in the supplementary material.
+Proofs and additional numerical results are provided in Appendix A.
 2 The testing framework
-Our problem of interest is whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$ (denoted as $\mathcal{N}(0,\sigma^{2})$) for some $\sigma>0$.
-We aim to test for the null hypothesis $H_{0}:e\sim\mathcal{N}(0,1)$ against the alternatives of nonnormality.
-To motivate our methodology, consider the transformed random variable $Z=\Phi(e)$, with $\Phi(\cdot)$ denoting the CDF of the standard normal distribution.
-Under the null hypothesis of normality, the CDF of $Z$ is given by
+Our problem of interest is to determine whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$, that is, $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ for some $\sigma>0$.
+Equivalently, this can be formulated as testing the null hypothesis about the standardized random error $H_{0}:e\sim\mathcal{N}(0,1)$ against alternatives of nonnormality.
+To motivate our methodology, consider the transformation $Z=\Phi(e)$, where $\Phi(\cdot)$ denotes the cumulative distribution function (CDF) of the standard normal distribution.
+Under the null hypothesis of normality, the CDF of the transformed random variable $Z$ is given by
 $G(z)=\mathrm{P}(Z\leq z)=\mathrm{P}\left(e\leq\Phi^{-1}(z)\right)=\Phi(\Phi^{-1}(z))=z.$
-Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1$ for $z\in[0,1]$.
-That is, $Z\sim\mathcal{U}[0,1]$ under $H_{0}$, where $\mathcal{U}[0,1]$ marks the uniform distribution on the interval $[0,1]$.
-Under the alternatives, the PDF $g(z)$ differs from 1.
-The distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ provides the basis for Neyman’s smooth test.
-Neyman, (1937) considered the following smooth alternative to the uniform density:
+Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1,z\in[0,1]$, that is, $Z\sim\mathcal{U}[0,1]$, where $\mathcal{U}[0,1]$ denotes the uniform distribution on the interval $[0,1]$.
+Under the alternative, the density $g(z)$ deviates from unity.
+This distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ provides the foundation for Neyman’s smooth test.
+In particular, Neyman, (1937) introduced the following smooth alternative to the uniform density:
 $h\left(z\right)=c\left(\bm{\theta}_{K}\right)\exp\left[\sum_{k=1}^{K}\theta_{k}\pi_{k}\left(z\right)\right],\quad 0\leq z\leq 1,$ (2.1)
-where $c(\bm{\theta}_{K})$ is a constant of integration depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ is an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
+where $c(\bm{\theta}_{K})$ is a a normalizing constant depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ denotes an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
 […]
-The null hypothesis that $Z\sim\mathcal{U}[0,1]$ can then be tested by testing $\theta_{1}=\theta_{2}=\ldots=\theta_{K}=0$ in (2.1).
-If we had independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ and $\sigma$ were known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
+The null hypothesis $Z\sim\mathcal{U}[0,1]$ can thus be assessed by testing $\theta_{1}=\theta_{2}=\ldots=\theta_{K}=0$ in (2.1).
+If independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ are available and $\sigma$ were known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
 $\Psi_{K}^{2}=\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}(Z_{i})\right)^{2},$ (2.3)
-where $K$ is some fixed positive integer and $Z_{i}=\Phi(e_{i})=\Phi(\varepsilon_{i}/\sigma)$.
-Under $H_{0}$, $n\Psi_{K}^{2}$ converges asymptotically to a Chi-square distribution with $K$ degrees of freedom (denoted as $\chi_{K}^{2}$), and the test has the local optimal properties of Rao’s score test.
-However, it is always the case that $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown.
-We may use some approximation $\{\hat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and some estimator $\hat{\sigma}$ for $\sigma$ (e.g., the residuals and the sample standard deviation, respectively).
-Then, consider the following approximation for $Z_{i}$:
-$\hat{Z}_{i}=\Phi(\hat{e}_{i})=\Phi\left(\frac{\hat{\varepsilon}_{i}}{\hat{\sigma}}\right),\quad i=1,\ldots,n.$
-With $\{\hat{Z}_{i}\}_{i=1}^{n}$ replacing $\{Z_{i}\}_{i=1}^{n}$, we can construct a feasible smooth test statistic of the following form:
-$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\hat{Z}_{i}\right)\right)^{2}.$
-Due to the replacement of $Z_{i}$ by $\hat{Z}_{i}$, the sequence $\{\hat{Z}_{i}\}_{i=1}^{n}$ may not be i.i.d.
-any longer.
+where $K$ is a fixed and given positive integer and $Z_{i}=\Phi(e_{i})=\Phi(\varepsilon_{i}/\sigma)$.
+Under $H_{0}$, the statistic $n\Psi_{K}^{2}$ converges in distribution to a Chi-square law with $K$ degrees of freedom, denoted by $\chi_{K}^{2}$.
+Moreover, the test inherits the local optimal properties of Rao’s score test.
+However, in practice, the sequence $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown.
+We instead rely on approximations $\{\hat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and an estimator $\hat{\sigma}$ for $\sigma$ (e.g., residuals and the sample standard deviation, respectively).
+Accordingly, we consider the following approximation for $Z_{i}$:
+$\hat{Z}_{i}=\Phi(\hat{e}_{i})=\Phi\left(\frac{\hat{\varepsilon}_{i}}{\hat{\sigma}}\right),\quad i=1,\ldots,n,$
+and construct a feasible test statistic of the following form:
+$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\hat{Z}_{i}\right)\right)^{2}.$ (2.4)
+The random variables $\{\hat{Z}_{i}\}_{i=1}^{n}$ are not i.i.d.
+any longer due to estimation effects caused by $\hat{\varepsilon}_{i}$ and $\hat{\sigma}$.
 Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\hat{Z}_{i})$.
-Compared with $\Psi_{K}^{2}$ in (2.3), $\hat{Z}_{i}$ has non-negligible estimation effects, as demonstrated in the theoretical results below.
-Consequently, the presence of estimation effects invalidates the above form of the test statistic and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution.
-This motivates us to study the estimation effects carefully and then derive the correct quadratic-form test statistics.
-It is also important to address the issue of sample sizes in ANOVA models.
-Throughout this section, to highlight the essence of smooth tests, we do not directly consider the ANOVA models but first examine the testing for a single random sequence of sample size $n$.
-This simplified setting captures the core ideas of the methodology and facilitates the derivation of key asymptotic properties.
-In the complete ANOVA framework, additional complications arise due to multiple sample sizes and the indexing structure, which require a more elaborate treatment.
-Besides, from a practical perspective, our method is expected to accommodate scenarios in which the number of groups grows with the total sample size.
-In subsequent analysis, we will systematically tackle these issues for specific ANOVA models.
+Compared with $\Psi_{K}^{2}$ in (2.3), the statistic based on $\hat{Z}_{i}$ suffers from non-negligible estimation effects, as demonstrated in the theoretical results below.
+Consequently, the presence of estimation effects invalidates the form of the “naive” statistic (2.4) and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution.
+This motivates a detailed analysis of estimation effects, leading to the correct quadratic-form test statistic.
+Another important issue concerns sample sizes in ANOVA models.
+To highlight the essence of smooth tests, we first focus on the simplified setting of a single random sequence with sample size $n$ throughout this section.
+This simplified setting captures the core methodological ideas and facilitates the derivation of key asymptotic properties.
+In contrast, the complete ANOVA setting involves additional complications arising from multiple sample sizes and the associated indexing structure, which demand more refined analysis.
+From a practical standpoint, our method is further designed to accommodate scenarios where the number of groups increases with the total sample size.
+These issues will be systematically investigated in subsequent analysis of specific ANOVA models.
 3 Testing for normality in one-way fixed effects models
-In this section, we establish smooth tests for normality in three types of one-way fixed effects models where either the group means or the group variances are equal.
-We examine the testing procedure and the theoretical properties separately for each case.
+In this section, we develop smooth tests for normality in three types of one-way fixed effects models, where each model assumes equality in at least one of the group means or the group variances.
+The testing procedure and its theoretical properties are investigated separately for each case.
 Let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observations of interest coming from $J$ groups, with $N_{j}$ being the sample sizes in group $j$ for $1\leq j\leq J$.
-Define $N=\sum_{j=1}^{J}N_{j}$ as the total sample size.
-In our theoretical framework, the number of groups $J$ is allowed to diverge as $N$ increases.
+The total sample size is defined as $N=\sum_{j=1}^{J}N_{j}$.
+In our theoretical framework, the number of groups, $J$, is allowed to diverge as $N$ increases.
 3.1 Same mean and same variance
-We first investigate the ANOVA models with equal group means and equal group variances. Consider the following model:
+We begin with the ANOVA models where both the group means and group variances are equal. Specifically, consider the model
 $Y_{ij}=\mu+\varepsilon_{ij}=\mu+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.1)
 where $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d.
-random errors with mean zero and variance one, $\mu$ is the common group mean, and $\sigma$ is the standard deviation of random errors.
+standardized random errors with mean zero and unit variance, $\mu$ denotes the common group mean, and $\sigma$ is the standard deviation of random errors $\{\varepsilon_{ij}\}_{i=1,j=1}^{N_{j},J}$.
 We also define $Z_{ij}=\Phi(e_{ij})$.
-We estimate $\mu$ and $\sigma^{2}$ in (3.1) by the sample mean $\hat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}\right)^{2}$, respectively.
+The parameters $\mu$ and $\sigma^{2}$ in (3.1) are estimated by the sample mean $\hat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}\right)^{2}$, respectively.
 Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu})/\hat{\sigma}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
-To derive the test statistic and its asymptotic distribution, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij}),k=1,\ldots,K$.
-The following two assumptions are imposed for our asymptotic analysis.
+To derive the feasible test statistic and its properties, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ for $k=1,\ldots,K$.
+For this purpose, we impose the following two assumptions.
 Assumption 3.1.
-$\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. with continuous CDF $F(x)$, PDF $f(x)$, mean zero, variance one, and finite fourth moments.
+$\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. with continuous CDF $F(x)$, PDF $f(x)$, mean zero, unit variance, and finite sixth moments.
 Assumption 3.2.
-For $k=1,\ldots,K$, $\pi_{k}(\cdot)$ are twice differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
-The two mild conditions are in line with similar assumptions adopted in the literature of Neyman’s smooth tests, including Bera et al., (2013) and Song and Xiao, (2022). Introduce two constants:
+For $k=1,\ldots,K$, $\pi_{k}(\cdot)$ are twice continuously differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
+The two mild assumptions are in line with similar ones adopted in the literature of Neyman’s smooth tests, including Bera et al., (2013) and Song and Xiao, (2022).
+For subsequent analysis, we introduce two constants:
 $c_{1k}=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z,\quad c_{2k}=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z,\quad k=1,\ldots,K.$
@@
 as $N\to\infty$.
-Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ is asymptotically equivalent to the sum of three terms: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$.
+Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ is equivalent to the sum of three terms after neglecting the higher-order term: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$.
 The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$.
-With the assistance of Theorem 1, by the central limit theorem (CLT), we can obtain the asymptotic normality of the $K$-dimensional vector
+With the assistance of Theorem 1, and by invoking the central limit theorem (CLT), we obtain the asymptotic normality of the $K$-dimensional vector
 $\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$ (3.2)
 $\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Sigma}_{K}\right),$
-where $\bm{\Sigma}_{K}=\left(\sigma_{kl}\right)_{K\times K}$ is given by
+where the asymptotic covariance matrix $\bm{\Sigma}_{K}=\left(\sigma_{kl}\right)_{K\times K}$ is given by
 $\sigma_{kl}$ $=\mathrm{E}\left[\left(\pi_{k}(Z)-c_{1k}e-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right)\left(\pi_{l}(Z)-c_{1l}e-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right)\right]$
 $=\delta_{kl}-c_{1k}c_{1l}-\frac{1}{2}c_{2k}c_{2l}.$
-Based on this fact, our proposed feasible smooth test statistic is defined as
+Based on this result, we define the feasible smooth test statistic as
 $\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\bm{\Sigma}_{K}^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.3)
@@
 as $N\to\infty$.
-Corollary 1 demonstrates an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\hat{\Psi}_{K}^{2}$.
+Corollary 1 establishes an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\hat{\Psi}_{K}^{2}$.
 Given the asymptotic significance level of $\alpha$, we reject the null hypothesis if
@@
 where $\chi_{K,1-\alpha}^{2}$ denotes the $(1-\alpha)$-th quantile of the $\chi_{K}^{2}$ distribution.
-Now we turn to the asymptotic behavior of $\hat{\Psi}_{K}^{2}$ under the fixed alternatives and a Pitman-type sequence of local alternatives.
+We now investigate the asymptotic behavior of $\hat{\Psi}_{K}^{2}$ under the fixed alternatives as well as under a Pitman-type sequence of local alternatives.
 For the fixed alternatives, we consider the following form:
 […]
-Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed, which leads to different asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ from those under $H_{0}$.
-To formally present the corresponding theoretical results under $H_{1}$, we need some notations.
+Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed.
+Consequently, the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ differs from that under $H_{0}$.
+To formally state the corresponding theoretical results under $H_{1}$, we introduce the following notations.
 Let
@@
 $d_{4k}=\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)x^{2}f(x)\mathrm{d}x,$
-for $k=1,\ldots,K$. Note that under $H_{0}$, $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ due to $f(x)=\phi(x)$. The following theorem demonstrates the asymptotic properties under $H_{1}$.
+for $k=1,\ldots,K$.
+Note that under $H_{0}$, the constants $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ since $f(x)=\phi(x)$.
+The following theorem characterizes the asymptotic properties of $\hat{\Psi}_{K}^{2}$ under the alternative hypothesis $H_{1}$.
 Theorem 2.
@@
 $-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
-From Theorem 2, one can see that under $H_{1}$ in (3.4), if there exists at least one $1\leq k\leq K$ such that
+From Theorem 2, it follows that under the alternative hypothesis $H_{1}$ in (3.4), if there exists at least one $1\leq k\leq K$ such that
 $a_{k}\equiv\mathrm{E}[\pi_{k}(Z)]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))f(x)\mathrm{d}x=\int_{0}^{1}\pi_{k}(z)\frac{f(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z\neq 0,$
-the smooth test statistic $N\hat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, and the asymptotic power of the test is 1. Specifically, the power function is given by
+then the smooth test statistic satisfies $N\hat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, implying that the asymptotic power of the test is $1$.
+Specifically, the power function is given by
 $1-\Phi\left(\left(4N\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)^{-1/2}\chi_{K,1-\alpha}^{2}-\left(N\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}/4\right)^{1/2}\right).$
-For further power analysis, we consider a Pitman-type sequence of local alternatives converging to the null hypothesis at a suitable rate, which is specified as follows:
+For further investigation of the test’s power, we consider a Pitman-type sequence of local alternatives that converges to the null hypothesis at an appropriate rate, which is specified as follows:
 $H_{1L}:F(x)=(1-\delta_{N})\Phi(x)+\delta_{N}Q(x),$ (3.7)
-where $Q\left(x\right)$ (which admits PDF $q\left(x\right)$) represents some distribution function that is different from $\Phi(x)$, and $\delta_{N}\to 0$ as $N\to\infty$. Denote
+where $Q\left(x\right)$ (which admits PDF $q\left(x\right)$) represents some distribution function that is different from $\Phi(x)$, and $\delta_{N}\to 0$ as $N\to\infty$. Define
 $\Delta_{k}=\int_{0}^{1}\pi_{k}(z)\frac{q(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z,\quad k=1,\ldots,K.$
 The following theorem presents the theoretical properties of $N\hat{\Psi}_{K}^{2}$ under the local alternatives.
-Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the local power of the test.
+Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the nontrivial local power of the test.
 Theorem 3.
@@
 where $\bm{\Delta}_{K}=(\Delta_{1},\ldots,\Delta_{K})^{\top}$, and $\chi^{2}_{K}\left(\tau\right)$ denotes the noncentral $\chi^{2}$ distribution with $K$ degrees of freedom and nonnegative noncentrality parameter $\tau$.
-It is worthwhile to mention that $N^{-1/2}$ is the critical order of magnitude for our test to have a nontrivial power against the Pitman sequence of local alternatives.
-It would be impossible to distinguish the alternatives from the null if $\delta_{N}=o(N^{-1/2})$.
-Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\hat{\Psi}_{K}^{2}$ has nontrivial asymptotic power against the local alternatives.
+It is worthwhile to mention that $N^{-1/2}$ is the critical order of magnitude for our test to achieve nontrivial power against the Pitman-type sequence of local alternatives.
+If $\delta_{N}=o(N^{-1/2})$, it would be impossible to distinguish the alternatives from the null.
+Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\hat{\Psi}_{K}^{2}$ attains nontrivial asymptotic power against the local alternatives because the noncentrality parameter $\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}>0$.
 Remark 1.
-We should emphasize that the proposed smooth test is not a “consistent” test in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives.
-Only under alternatives where $\mathrm{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power go to 1.
+We emphasize that the aforementioned smooth test is not “consistent” in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives.
+Only under alternatives where $\mathrm{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power converge to $1$.
 Otherwise, if $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathrm{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\hat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$.
-For example, if $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ such that $\mathrm{E}[e]=0$ and $\operatorname{Var}[e]=1$, and we take the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$, then for $Z=\Phi(e)$, we have
+For example, let $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ so that $\mathrm{E}[e]=0$ and $\operatorname{Var}[e]=1$, and consider the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$.
+For $Z=\Phi(e)$, we have
 $\mathrm{E}[Z]$ $=\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}\Phi(z)\mathrm{d}z$
@@
 $=\frac{1}{2},$
-which leads to $\mathrm{E}[\pi_{1}(Z)]=0$.
-In fact, smooth tests are neither directional nor omnibus; that is, they maintain decent power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see Bera et al., (2013) for details.
+which implies $\mathrm{E}[\pi_{1}(Z)]=0$.
+In fact, smooth tests are neither directional nor omnibus; that is, they maintain reasonable power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see Bera et al., (2013) for further discussion.
 3.2 Different means and same variance
-Next we allow the group means $\mu$ in model (3.1) to vary across the group index $j$, in the sense that
+Next, we consider ANOVA models in which the group means are allowed to differ, that is,
 $Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.9)
-where $\{\mu_{j}\}_{j=1}^{J}$ represents the potentially different means of each group.
-Let $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ be estimators of $\mu_{j},j=1,\ldots,J$ and $\sigma^{2}$, respectively.
+where $\{\mu_{j}\}_{j=1}^{J}$ denote the potentially distinct means of each group.
+Let $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ be the estimators of $\mu_{j}$ ($j=1,\ldots,J$) and $\sigma^{2}$, respectively.
 Correspondingly we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
-The following theorem presents the null limiting properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ under the settings for model (3.9).
+The following theorem characterizes the properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ under the model (3.9).
 Theorem 4.
@@
 $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
-as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\hat{\Psi}_{K}^{2}$ defined in (3.3), we still have
+as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\hat{\Psi}_{K}^{2}$ defined in (3.3), we have
 $N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}.$
-Theorem 4 is parallel with Theorem 1 and Corollary 1 except for extra conditions regarding the group sample sizes and the number of groups.
-These additional conditions are imposed for technical considerations and to account for the heterogeneity across group means from model (3.9).
-We illustrate the intuition behind the conditions as follows.
-First, we require that the sample size of each group, $N_{j}$, tends to infinity.
-For model (3.9) where $\mu_{j}$ is allowed to be distinct, only the observations in group $j$ can be used to estimate $\mu_{j}$, thus $N_{j}\to\infty$ is necessary for the consistency of $\hat{\mu}_{j}$.
-Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must be satisfied jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$, which rules out the case that the number of groups grows faster than the group sample sizes.
-In contrast, for model (3.1) in the previous section, only the total sample size $N\to\infty$ suffices to derive the asymptotic properties regardless of the magnitude of $N_{j}$s and $J$.
-This is because in model (3.1), all groups share the same mean $\mu$ and variance $\sigma^{2}$.
-Thus, the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are all i.i.d., and we can just pool them together for efficient estimation and inference.
-The following two theorems establish the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ (defined by (3.3)) under the alternatives.
-The conditions are the same as those of Theorem 4, and the results are parallel with those in Theorems 2 and 3.
+Theorem 4 parallels Theorem 1 and Corollary 1, except that it requires additional conditions on the group sample sizes and the number of groups.
+These conditions are imposed for technical reasons and to account for heterogeneity across group means in model (3.9).
+The intuition behind the conditions is illustrated as follows.
+First, the sample size of each group, $N_{j}$, must tend to infinity.
+In model (3.9), where $\mu_{j}$ can be distinct, only the observations within group $j$ can be used to estimate $\mu_{j}$, making $N_{j}\to\infty$ necessary for the consistency of $\hat{\mu}_{j}$.
+Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must hold jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$.
+This rules out scenarios where the number of groups grows faster than the sample sizes within each group.
+In contrast, for model (3.1) in the previous section, it suffices that the total sample size $N\to\infty$ to derive asymptotic properties, regardless of the magnitudes of $N_{j}$ and $J$.
+This is because in model (3.1), all groups share the same mean $\mu$ and variance $\sigma^{2}$, so the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d., allowing them to be pooled for efficient estimation and inference.
+The following two theorems establish formally the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ (defined in (3.3)) under the alternatives.
+The conditions are identical to those in Theorem 4, and the results parallel those in Theorems 2 and 3.
 Theorem 5.
@@
 3.3 Same mean and different variances
-Now we allow the group variances $\sigma$ in model (3.1) to be different among the groups, in the sense that
+We now extend model (3.1) by allowing the group variances to differ across groups, that is,
 $Y_{ij}=\mu+\varepsilon_{ij}=\mu+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.10)
-where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially different standard deviations of each group.
-This time, we adopt the estimators $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\hat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2},j=1,\ldots,J$, respectively.
-Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of variation across groups.
-Let $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu})/\hat{\sigma}_{j}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
-Define $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ regarding the relative proportion among the group sample sizes.
+where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially distinct standard deviations of each group.
+For estimation, we adopt $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\hat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2}$ ($j=1,\ldots,J$), respectively.
+Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of group variances.
+Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu})/\hat{\sigma}_{j}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
+Let $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ be quantities characterizing the relative proportions of the group sample sizes.
 Here we further impose the following conditions for model (3.10): (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$.
-They both imply that the sample sizes and the standard deviations of each group are of the same magnitude.
-Like the procedure in Sections 3.1 and 3.2 above, we start from the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ here, see the following theorem.
+These conditions imply that the group-specific sample sizes and standard deviations are of the same order of magnitude.
+Following the approaches in Sections 3.1 and 3.2, we first examine the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, which is summarized in the theorem below.
 Theorem 7.
@@
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$.
-Note that the decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4.
-The difference lies in the parameter estimation effect of $\hat{\mu}$ characterized by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right hand side of (3.11).
-Unlike the adoption of the sample mean for estimating $\mu$ in model (3.1), here we consider a new form of $\hat{\mu}$ for model (3.10), which leads to a different estimation effect term.
-Specially, when all groups share the same variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term can be simplified as $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect reduces to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$.
-In contrast, the estimation effect of the variances is still aggregated as the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
-The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$ can be derived from (3.11). Under the conditions of Theorem 7, by CLT,
+Note that the asymptotic decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4.
+The difference lies in the parameter estimation effect of $\hat{\mu}$, which is reflected by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right-hand side of (3.11).
+Unlike model (3.1), which uses the sample mean to estimate $\mu$, model (3.10) employs a different estimator, $\hat{\mu}$, resulting in this distinct estimation effect.
+In particular, when all groups share the same variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term reduces to $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect simplifies to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$.
+In contrast, the estimation effect of the variances always aggregates into the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
+The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$ follows directly from (3.11).
+Under the conditions of Theorem 7, an application of the CLT yields that
 $\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
@@
 $\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right).$
-By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ by $\hat{\sigma}_{j}$, $\hat{p}_{j}=N_{j}/N$ and $\hat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted as $\widehat{\bm{\Omega}}_{K}^{(j)}$.
-Correspondingly, the test statistic takes the form
+By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ with their sample analogues $\hat{\sigma}_{j}$, $\hat{p}_{j}=N_{j}/N$ and $\hat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted by $\widehat{\bm{\Omega}}_{K}^{(j)}$.
+Accordingly, the test statistic is defined as
 $\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}\hat{p}_{j}\widehat{\bm{\Omega}}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.12)
-We also define the infeasible version,
+We also introduce the infeasible version,
 $\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.13)
@@
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$.
-Corollary 2 demonstrates that under the null hypothesis, both the feasible test statistic $\hat{\Psi}_{K}^{2}$ (in (3.12)) and its infeasible version $\tilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed after multiplied by $N$, the total sample size.
-Although the test statistics considered here take a slightly different form from those in the earlier setting, the limiting Chi-square distribution is retained and the convergence rate remains $N$, thereby reinforcing the validity of our unified theoretical framework.
-The following two theorems establish the asymptotic behavior of $\tilde{\Psi}_{K}^{2}$ under the fixed and local alternatives, which is also analogous to that of $\hat{\Psi}_{K}^{2}$ in Sections 3.1 and 3.2.
+Corollary 2 shows that under the null hypothesis, both the feasible test statistic $\hat{\Psi}_{K}^{2}$ (defined in (3.12)) and its infeasible counterpart $\tilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed when multiplied by the total sample size $N$.
+Although these test statistics take a slightly different form from those in the previous Sections 3.1 and 3.2, the limiting Chi-square distribution is retained, and the convergence rate remains $N$, which further supports the validity of our unified theoretical framework.
+The following two theorems characterize the asymptotic behavior of $\tilde{\Psi}_{K}^{2}$ under both fixed and local alternatives, in a manner analogous to that of $\hat{\Psi}_{K}^{2}$ discussed in Sections 3.1 and 3.2.
 Theorem 8.
 Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right])-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
-$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
+$-\left.\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$. Furthermore,
@@
 Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right])-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
-$+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
+$-\left.\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$. Furthermore,
@@
 Remark 2.
-To conclude this section, we summarize the regularity conditions of the three models above in Table 1.
-On the one hand, one can find that more complex model structures lead to more restrictions on parameters and sample sizes.
-The heterogeneity across groups indeed poses additional challenges for theoretical analysis.
-On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can always be established for each of the three cases.
-This demonstrates that our proposed method possesses a unified theoretical framework while retaining the ability to capture the distinctive features of different ANOVA models.
+To conclude this section, we summarize the regularity conditions for the three models in Table 1.
+On one hand, more complex model structures naturally impose additional restrictions on parameters and group sample sizes, as the heterogeneity across groups introduces extra challenges for theoretical analysis.
+On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can be established for each model.
+This highlights that our proposed methodology provides a unified theoretical framework while remaining flexible enough to accommodate the distinctive features of different ANOVA settings.
 Table 1: The regularity conditions for the three models
@@
 4 Data-driven choice of $K$
-The methodology in Section 3 is based on the fact that the order of smooth alternatives (2.1), $K$, is a predetermined fixed integer before the tests are performed.
-In fact, the order $K$ has a key impact on the performance of smooth tests (Ledwina,, 1994; Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016).
-If $K$ is large, the alternatives we test against contain redundant components that do not contribute to the test statistic markedly but increase the number of degrees of freedom.
-This will make smooth tests more omnibus and lead to power dilution.
-If $K$ is small, as described in Section 3, it is more likely to happen that $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, which makes the tests exhibit no power.
-Therefore, it is crucial to select an appropriate $K$ before applying the test.
-In this section, we focus on the data-driven choice of $K$ for the three cases in Section 3.
-In the literature of Neyman’s smooth tests, the spirit of data-driven methods was first introduced by Ledwina, (1994) for the problem of testing for uniformity.
+The methodology in Section 3 relies on the fact that the order of smooth alternatives (2.1), $K$, is fixed before testing.
+The choice of $K$ critically affects the performance of smooth tests (Ledwina,, 1994; Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016).
+A large $K$ introduces redundant components in the alternatives that contribute little to the test statistic but inflate the degrees of freedom, leading to power dilution.
+Conversely, a small $K$ increases the risk that $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, rendering the test powerless.
+Hence, an appropriate choice of $K$ is crucial.
+In this section, we focus on data-driven selection of $K$ for the three cases in Section 3.
+In the literature of Neyman’s smooth tests, data-driven selection was first proposed by Ledwina, (1994) for testing uniformity.
 The data-driven procedure consists of two steps.
-First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to find a suitable dimension of the smooth model for the data.
+First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to determine a suitable dimension for the smooth model that best fits the data.
 Second, Neyman’s smooth test is performed within the selected model space, yielding a data-driven test statistic.
-Therefore, the data-driven smooth test serves as a kind of first selection followed by a more precise instrument, being Neyman’s test in the “right” direction.
-Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were shown in Kallenberg and Ledwina, 1995a and Kallenberg and Ledwina, 1995b , Inglot and Ledwina, (1996).
-Later, some modified versions of Schwarz’s selection rule were employed to address more complicated issues, such as testing for composite hypotheses (Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Inglot et al.,, 1997), testing for independence (Kallenberg and Ledwina,, 1999), two-sample testing problems (Janic-Wróblewska and Ledwina,, 2000), and so on.
-The modified selection rules utilize the smooth test statistics and circumvent the need to compute the maximized log-likelihood function of smooth models; thus, they are easier to implement.
-Motivated by the strategies mentioned above, we adopt a modified Schwarz’s rule to determine the order $K$ for our problem. The selection rule is specified as
+The data-driven smooth test thus serves to combine preliminary model selection with a more precise inferential procedure, being Neyman’s test in the “right” direction.
+Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were established in Kallenberg and Ledwina, 1995a and Kallenberg and Ledwina, 1995b , Inglot and Ledwina, (1996).
+Later, modifications of Schwarz’s rule addressed more complex problems, including testing for composite hypotheses (Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Inglot et al.,, 1997), testing for independence (Kallenberg and Ledwina,, 1999), two-sample testing problems (Janic-Wróblewska and Ledwina,, 2000), and so on.
+These modified rules leverage smooth test statistics directly, avoiding the computation of the maximized log-likelihood, and are easier to implement.
+Motivated by these strategies, we adopt a modified Schwarz’s rule to determine the order $K$ for smooth tests in ANOVA models. The selection rule is specified as follows:
 $\hat{K}=\min\left\{\mathop{\arg\max}\limits_{1\leq k\leq D}\left(N\hat{\Psi}_{k}^{2}-k\log N\right)\right\},$ (4.1)
-where $D$ is some fixed positive integer, and $\hat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) up to specific scenarios.
-The corresponding data-driven test statistic is $N\Psi_{\hat{K}}^{2}$, where $\hat{K}$ is given by (4.1).
+where $D$ is a fixed positive integer, and $\hat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) depending on the specific scenario.
+The resulting data-driven test statistic is $N\Psi_{\hat{K}}^{2}$, with $\hat{K}$ given by (4.1).
 The form of (4.1) is mainly inspired by Ducharme and Fontez, (2004), Ducharme and Lafaye de Micheaux, (2004), Kraus, (2007) and Duchesne et al., (2016) where the upper bound of selection $D$ is fixed.
-An alternative is to set $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000).
-Although the diverging upper bound helps to improve the consistency of smooth tests theoretically, the rate of divergence is required to be quite slow, and the empirical power levels off rapidly as $D$ increases in their simulation studies.
-Therefore, our proposed selection rule (4.1) with fixed $D$ is reasonable and simple to implement in practice.
-To establish the properties of $\hat{K}$ and $N\Psi_{\hat{K}}^{2}$, we need a revised version of the alternative hypotheses $H_{1}$ (3.4) and Assumption 3.2. Consider the following alternatives:
-$H_{1}^{\prime}:\mathrm{E}[\pi_{1}(Z)]=\ldots=\mathrm{E}[\pi_{K_{0}-1}(Z)]=0,\mathrm{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D.$ (4.2)
-In other words, $H_{1}^{\prime}$ encompasses a broader range of alternatives than $H_{1}$ due to the replacement of the fixed order $K$ with the larger $D$.
-Analogously, the condition on the orthonormal system is now broadened to include all functions up to order $D$; see the revised assumption below.
+An alternative is to let $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000).
+Although a diverging upper bound can theoretically improve the consistency of smooth tests, the required rate of divergence is slow, and simulation studies show that empirical power levels off rapidly as $D$ increases.
+Hence, our proposed selection rule (4.1) with fixed $D$ is both reasonable and easy to implement in practice.
+To establish the theoretical properties of $\hat{K}$ and $N\hat{\Psi}_{\hat{K}}^{2}$, we introduce a revised version of the alternative hypothesis $H_{1}$ (3.4) and Assumption 3.2. Consider
+$H_{1}^{\prime}:\mathrm{E}[\pi_{1}(Z)]=\ldots=\mathrm{E}[\pi_{K_{0}-1}(Z)]=0,\mathrm{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D,$ (4.2)
+which encompasses a broader range of alternatives than $H_{1}$ by replacing the fixed order $K$ with the larger $D$.
+Correspondingly, the condition on the orthonormal system is extended to include all functions up to order $D$, as follows.
 Assumption 4.1.
 For $k=1,\ldots,D$, $\pi_{k}(\cdot)$ are two times differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
-The following theorem demonstrates the properties of data-driven tests under the null and the revised alternatives.
+The following theorem establishes the properties of data-driven tests under the null and the revised alternatives.
 Theorem 10.
@@
 Theorem 10 presents the unified results of data-driven smooth tests for the three cases in Section 3.
-On the one hand, under the null hypothesis, the probability of $\{\hat{K}=1\}$ tends to $1$ asymptotically, which implies that the first-order smooth model fits $\{\hat{Z}_{ij}\}_{i=1,j=1}^{N_{j}.J}$ best among the $D$ candidate models.
-From the perspective of hypothesis testing, it is $\hat{\Psi}_{1}^{2}$ that is informative and adequate for diagnosing the uniformity/normality.
-On the other hand, the data-driven selection procedure helps identify the mechanism behind nonuniformity/nonnormality and leads to a more powerful test against a broader range of alternatives, $H_{1}^{\prime}$.
-These findings are parallel with existing results about data-driven smooth tests.
-However, in practice, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics is often inaccurate in finite samples (Kallenberg and Ledwina, 1995b, ; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016).
-To overcome this issue in our problem, provided the nested orthonormal system (i.e.
-$\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the instruction of Kallenberg and Ledwina, 1995b , Inglot et al., (1997), Kallenberg and Ledwina, (1997), Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007) and adopt a finite-sample approximation for the null distribution $N\hat{\Psi}_{\hat{K}}^{2}$:
+Under the null hypothesis, the probability of $\{\hat{K}=1\}$ tends to $1$ asymptotically, implying that the first-order smooth model provides the best fit to $\{\hat{Z}_{ij}\}_{i=1,j=1}^{N_{j},J}$ among the $D$ candidate models.
+From the perspective of hypothesis testing, this means that $\hat{\Psi}_{1}^{2}$ is informative and sufficient for assessing uniformity or normality.
+Meanwhile, the data-driven selection procedure allows for the identification of the mechanism behind nonuniformity (and also nonnormality) and enhances the test’s power against a broader class of alternatives, $H_{1}^{\prime}$.
+These findings are consistent with previous results on data-driven smooth tests.
+In practice, however, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics often performs poorly in finite samples (Kallenberg and Ledwina, 1995b, ; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016).
+To address this issue in our problem, and provided the nested orthonormal system (i.e.
+$\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the approach of Kallenberg and Ledwina, 1995b , Inglot et al., (1997), Kallenberg and Ledwina, (1997), Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007) and adopt the following finite-sample approximation for the null distribution of $N\hat{\Psi}_{\hat{K}}^{2}$:
 […]
 Such an approximation primarily accounts for the selection uncertainty of $\hat{K}$.
-In finite samples, under the null, the empirical probability of observing $\{\hat{K}=1\}$ may not be exactly one, as the event $\{\hat{K}=2\}$ can still occur with a small but negligible probability.
-Accordingly, (4.3) is obtained from $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x)\approx\mathrm{P}(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1)+\mathrm{P}(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2)$.
-The reliability and effectiveness of $H(x)$ are demonstrated through simulation studies in the following section.
+In finite samples, under the null, the empirical probability of the event $\{\hat{K}=1\}$ may not be exactly $1$, as $\{\hat{K}=2\}$ can occur with a small but negligible probability.
+Consequently, (4.3) is derived from the approximation $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x)\approx\mathrm{P}(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1)+\mathrm{P}(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2)$.
+The reliability and accuracy of $H(x)$ are further illustrated through simulation studies in the next section and the appendix.
 5 Simulation
 In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the proposed tests developed in Sections 3 and 4.
-Following Neyman, (1937) and Song and Xiao, (2022), we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$.
+Following Neyman, (1937), Bera et al., (2013) and Song and Xiao, (2022), we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$.
 Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of Duchesne et al., (2016), and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material.
-From a computational perspective, the test statistic $N\hat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in Duchesne et al., (2016), which leverages the closed-form of $\bm{\Sigma}_{K}^{-1}$ (see equations (12) and (13) therein).
-Since Section 3.1 is a special case of Section 3.2 or 3.3, we focus our evaluation on the settings described in Sections 3.2 and 3.3 through two numerical experiments.
-In the first experiment, we generate observations with different means and the same variances as follows:
-$Y_{ij}^{(0,1)}\sim\mathcal{N}\left(5j,4\right),\quad Y_{ij}^{(1,1)}\sim\chi_{2}^{2}+(5j-2),\quad i=1,\ldots,jm,\quad j=1,\ldots,5,$
-where $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$, $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$ and $H_{1}$ respectively. Note that
-$\mathrm{E}[Y_{ij}^{(0,1)}]=\mathrm{E}[Y_{ij}^{(1,1)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,1)}]=\operatorname{Var}[Y_{ij}^{(1,1)}]=4$.
+From a computational perspective, the test statistic $N\hat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in Duchesne et al., (2016), which leverages the closed form of $\bm{\Sigma}_{K}^{-1}$ [see equations (12) and (13) therein].
+We first study the tests for model (3.1) through two experiments. The observations are generated as follows.
+  • Experiment I:
+$Y_{ij}^{(0,1)}\sim\mathcal{N}\left(5,4\right),\quad Y_{ij}^{(1,1)}\sim\chi_{2}^{2}+3,\quad i=1,\ldots,jm,\quad j=1,\ldots,5.$
+  • Experiment II:
+$Y_{ij}^{(0,2)}\sim\mathcal{N}\left(8,1\right),\ Y_{ij}^{(1,2)}\sim\mathcal{U}\left[8-\sqrt{3},8+\sqrt{3}\right],\ i=1,\ldots,jm,\ j=1,\ldots,5.$
+Here $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$.
 The sample size parameter $m$ varies from 10 to 150 in increments of 10.
 We set the significance level $\alpha=5\%$ and perform $500$ replications.
-We evaluate the performance of the test statistic $N\hat{\Psi}_{K}^{2}$ introduced in Section 3.2, with $1\leq K\leq 5$, as well as the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ from Section 4, where $\hat{K}$ is determined by (4.1) with $D=5$.
+We evaluate the test statistic $N\hat{\Psi}_{K}^{2}$ from Section 3.1, with $1\leq K\leq 5$, and the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ from Section 4, where $\hat{K}$ is determined by (4.1) with $D=5$.
+Results of Experiment I are shown in Tables 2, 3, 4 and Figure 1.
 Table 2 reports the empirical rejection rates under $H_{0}$.
-For fixed values of $K=1,\ldots,5$, $N\hat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes, such as $m=10$ (corresponding to a total sample size of $N=150$).
-For the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$, the results indicate that the limiting $\chi^{2}_{1}$ distribution does not adequately approximate the null distribution, whereas the revised distribution $H(x)$ achieves better size control.
-The empirical frequency of $\hat{K}$ under $H_{0}$ and $H_{1}$ is presented in Tables 3 and 4, respectively.
-Under $H_{0}$, the frequency of $\{\hat{K}=1\}$ approaches 1 as $m$ increases, while under $H_{1}$, $\hat{K}$ is consistently greater than 1, which aligns with our theoretical expectations.
-Additionally, Figure 1 displays the sample means and corresponding error bars (sample mean $\pm$ sample standard deviation) of $\hat{K}$ across different values of $m$.
-Here, under our data-generating processes, under $H_{0}$, the sample mean of $\hat{K}$ remains close to 1 regardless of the sample size, and the variability diminishes as $m$ increases, indicating consistent model selection under the null.
-Under $H_{1}$, $\hat{K}$ increases steadily and converges to the maximum selection range $D=5$ as $m$ grows, with reduced variability.
-These findings align with our theoretical results and demonstrate the adaptiveness of the data-driven procedure.
-Table 2: Empirical rejection rates under $H_{0}$ in the first experiment, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
+For fixed $K$, $N\hat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes such as $m=10$ (corresponding to a total sample size of $N=150$).
+For the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$, the limiting $\chi^{2}_{1}$ distribution underestimates the variability under $H_{0}$, while the revised distribution $H(x)$ achieves better size control.
+Tables 3 and 4 present the empirical frequency of $\hat{K}$ under $H_{0}$ and $H_{1}$, respectively.
+Under $H_{0}$, the frequency of $\{\hat{K}=1\}$ approaches $1$ as $m$ increases, whereas under $H_{1}$, $\hat{K}$ is consistently greater than $1$, in line with theoretical expectations.
+Figure 1 displays the sample means of $\hat{K}$ across different values of $m$, with error bars representing one standard deviation around the mean.
+Under $H_{0}$, the sample mean of $\hat{K}$ stays close to $1$ with diminishing variability as $m$ grows, indicating consistent model selection.
+Under $H_{1}$, $\hat{K}$ increases steadily and approaches the maximum $D=5$, with reduced variability, demonstrating the adaptiveness of the data-driven procedure.
+The results of Experiment II are quite similar; see Tables 5, 6, and 7, as well as Figure 2.
+However, from Table 5, under $H_{1}$, the test statistic $N\hat{\Psi}_{1}^{2}$ fails to exhibit power because in this case, $\mathrm{E}[\pi_{1}(Z)]=\mathrm{E}[\pi_{1}(\Phi(Y_{ij}^{(1,2)}-8))]=0$.
+This illustrates that $N\hat{\Psi}_{K}^{2}$ may be inconsistent against certain alternatives, especially when $K$ is small.
+Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, $\{\hat{K}=1\}$ does not occur, consistent with Theorem 10 (3), which ensures that $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ under $H_{1}^{\prime}$ with $K_{0}=2$.
+Table 2: Empirical rejection rates under $H_{0}$ in Experiment I, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
+The empirical power under $H_{1}$ equals 1 across all settings.
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
+10 0.050 0.048 0.048 0.044 0.040 0.056 0.034
+20 0.046 0.044 0.034 0.040 0.048 0.066 0.046
+30 0.042 0.036 0.044 0.068 0.068 0.060 0.046
+40 0.058 0.040 0.036 0.058 0.052 0.058 0.050
+50 0.064 0.044 0.054 0.054 0.062 0.068 0.060
+60 0.054 0.034 0.038 0.044 0.050 0.070 0.062
+70 0.056 0.038 0.048 0.046 0.052 0.056 0.048
+80 0.038 0.050 0.072 0.042 0.048 0.058 0.048
+90 0.062 0.066 0.052 0.064 0.058 0.048 0.040
+100 0.034 0.050 0.050 0.060 0.054 0.052 0.038
+110 0.054 0.032 0.054 0.060 0.050 0.038 0.036
+120 0.052 0.052 0.038 0.044 0.052 0.052 0.050
+130 0.052 0.050 0.064 0.060 0.058 0.068 0.064
+140 0.062 0.040 0.026 0.046 0.030 0.038 0.030
+150 0.054 0.044 0.062 0.034 0.058 0.056 0.048
+Table 3: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment I
+$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
+10 0.982 0.018 0 0 0
+20 0.988 0.010 0 0 0.002
+30 0.986 0.012 0 0.002 0
+40 0.984 0.014 0.002 0 0
+50 0.988 0.012 0 0 0
+60 0.980 0.020 0 0 0
+70 0.994 0.006 0 0 0
+80 0.996 0.002 0.002 0 0
+90 0.996 0.002 0.002 0 0
+100 0.996 0.004 0 0 0
+110 0.994 0.006 0 0 0
+120 0.990 0.010 0 0 0
+130 0.994 0.006 0 0 0
+140 0.994 0.006 0 0 0
+150 0.994 0.006 0 0 0
+Table 4: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment I
+$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
+10 0.004 0 0 0.484 0.512
+20 0 0 0 0.326 0.674
+30 0 0 0 0.194 0.806
+40 0 0 0 0.090 0.910
+50 0 0 0 0.076 0.924
+60 0 0 0 0.030 0.970
+70 0 0 0 0.012 0.988
+80 0 0 0 0.010 0.990
+90 0 0 0 0.006 0.994
+100 0 0 0 0.004 0.996
+110 0 0 0 0.006 0.994
+120 0 0 0 0 1
+130 0 0 0 0 1
+140 0 0 0 0.002 0.998
+150 0 0 0 0 1
+Figure 1: The sample means and error bars of $\hat{K}$ in Experiment I
+Table 5: Empirical rejection rates in Experiment II, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
+$H_{0}$ $H_{1}$
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
+10 0.042 0.050 0.046 0.042 0.052 0.076 0.048 0.046 0.998 0.996 0.998 0.996 0.998 0.998
+20 0.034 0.042 0.026 0.050 0.060 0.070 0.054 0.018 1 1 1 1 1 1
+30 0.028 0.048 0.036 0.046 0.046 0.070 0.062 0.022 1 1 1 1 1 1
+40 0.054 0.054 0.046 0.050 0.060 0.062 0.052 0.042 1 1 1 1 1 1
+50 0.064 0.062 0.042 0.060 0.050 0.068 0.058 0.032 1 1 1 1 1 1
+60 0.032 0.054 0.060 0.040 0.040 0.078 0.064 0.026 1 1 1 1 1 1
+70 0.076 0.038 0.052 0.052 0.036 0.052 0.042 0.022 1 1 1 1 1 1
+80 0.040 0.046 0.034 0.048 0.058 0.074 0.068 0.028 1 1 1 1 1 1
+90 0.044 0.052 0.050 0.054 0.042 0.060 0.050 0.032 1 1 1 1 1 1
+100 0.038 0.068 0.050 0.046 0.046 0.064 0.056 0.028 1 1 1 1 1 1
+110 0.052 0.052 0.046 0.052 0.048 0.050 0.046 0.032 1 1 1 1 1 1
+120 0.052 0.042 0.038 0.034 0.052 0.040 0.036 0.030 1 1 1 1 1 1
+130 0.048 0.052 0.044 0.036 0.048 0.062 0.054 0.040 1 1 1 1 1 1
+140 0.062 0.058 0.056 0.038 0.038 0.046 0.042 0.044 1 1 1 1 1 1
+150 0.038 0.056 0.046 0.052 0.060 0.060 0.054 0.034 1 1 1 1 1 1
+Table 6: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment II
+$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
+10 0.968 0.026 0.004 0 0.002
+20 0.982 0.016 0.002 0 0
+30 0.978 0.020 0.002 0 0
+40 0.992 0.008 0 0 0
+50 0.990 0.006 0.004 0 0
+60 0.996 0.004 0 0 0
+70 0.994 0.006 0 0 0
+80 0.984 0.016 0 0 0
+90 0.996 0.004 0 0 0
+100 0.994 0.006 0 0 0
+110 0.996 0.004 0 0 0
+120 0.996 0.004 0 0 0
+130 0.994 0.006 0 0 0
+140 0.994 0.006 0 0 0
+150 0.988 0.012 0 0 0
+Table 7: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment II
+$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
+10 0.002 0.984 0.012 0 0.002
+20 0 0.994 0.004 0 0.002
+30 0 0.990 0.008 0 0.002
+40 0 0.996 0.004 0 0
+50 0 0.996 0.004 0 0
+60 0 0.994 0.004 0 0.002
+70 0 0.998 0.002 0 0
+80 0 1 0 0 0
+90 0 1 0 0 0
+100 0 0.996 0.004 0 0
+110 0 0.998 0.002 0 0
+120 0 0.994 0.006 0 0
+130 0 0.998 0.002 0 0
+140 0 0.998 0.002 0 0
+150 0 0.996 0.004 0 0
+Figure 2: The sample means and error bars of $\hat{K}$ in Experiment II
+The next two experiments are designed to evaluate the tests for models (3.9) and (3.10), respectively.
+All settings are inherited from the previous two experiments, except that the data-generating process is modified accordingly.
+  • Experiment III:
+$Y_{ij}^{(0,3)}\sim\mathcal{N}\left(5j,4\right),\quad Y_{ij}^{(1,3)}\sim\chi_{2}^{2}+(5j-2),\quad i=1,\ldots,jm,\quad j=1,\ldots,5.$
+  • Experiment IV:
+$Y_{ij}^{(0,4)}\sim\mathcal{N}\left(8,j^{2}\right),\ Y_{ij}^{(1,4)}\sim\mathcal{U}\left[8-\sqrt{3}j,8+\sqrt{3}j\right],\ i=1,\ldots,jm,\ j=1,\ldots,5.$
+Here $\{Y_{ij}^{(0,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$.
+Note that in Experiment III, $\mathrm{E}[Y_{ij}^{(0,3)}]=\mathrm{E}[Y_{ij}^{(1,3)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,3)}]=\operatorname{Var}[Y_{ij}^{(1,3)}]=4$, which corresponds to model (3.9); in Experiment IV, $\mathrm{E}[Y_{ij}^{(0,4)}]=\mathrm{E}[Y_{ij}^{(1,4)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,4)}]=\operatorname{Var}[Y_{ij}^{(1,4)}]=j^{2}$, which corresponds to model (3.10).
+The conclusions from Experiments III and IV are analogous to those from Experiments I and II, respectively (see Tables 813 and Figures 34), thus validating our unified testing framework.
+To examine the robustness of our findings with respect to the number of groups, we additionally consider a variant of Experiment III with $J=10$ (referred to as Experiment III).
+The results corroborate the main conclusions, as reported in A.
+Table 8: Empirical rejection rates under $H_{0}$ in Experiment III, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
 The empirical power under $H_{1}$ equals 1 across all settings.
@@
 $150$ 0.064 0.056 0.048 0.058 0.058 0.046 0.042
-Table 3: Empirical frequency of $\hat{K}$ under $H_{0}$ in the first experiment
+Table 9: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment III
 $m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
@@
 $150$ 0.996 0.004 0 0 0
-Table 4: Empirical frequency of $\hat{K}$ under $H_{1}$ in the first experiment
+Table 10: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment III
 $m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
@@
 $150$ 0 0 0 0 1
-Figure 1: The sample means and error bars of $\hat{K}$ in the first experiment
-Figure 2: The sample means and error bars of $\hat{K}$ in the second experiment
-In the second experiment, we generate observations with the same means and different variances as follows:
-$Y_{ij}^{(0,2)}\sim\mathcal{N}\left(8,j^{2}\right),\ Y_{ij}^{(1,2)}\sim\mathcal{U}\left[8-\sqrt{3}j,8+\sqrt{3}j\right],\ i=1,\ldots,jm,\ j=1,\ldots,5,$
-where $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$, $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$ and $H_{1}$ respectively. Note that
-$\mathrm{E}[Y_{ij}^{(0,2)}]=\mathrm{E}[Y_{ij}^{(1,2)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,2)}]=\operatorname{Var}[Y_{ij}^{(1,2)}]=j^{2}$.
-We adopt the same settings as in the first experiment.
-The results are analogous to those observed previously; see Tables 5, 6, and 7, as well as Figure 2.
-However, from Table 5, we observe that under $H_{1}$, the test statistic $N\hat{\Psi}_{1}^{2}$ fails to demonstrate power.
-This is because $\mathrm{E}[\pi_{1}(Z)]=\mathrm{E}[\pi_{1}(\Phi((Y_{ij}^{(1,2)}-8)/j))]=0$, indicating that $N\hat{\Psi}_{K}^{2}$ may lack consistency against certain directions of alternatives, particularly when $K$ is small.
-Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, the event $\{\hat{K}=1\}$ does not occur, which is consistent with Theorem 10 (3), where it is established that $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ under $H^{\prime}_{1}$ with $K_{0}=2$.
-To summarize the simulation results, we conclude that the proposed test statistic $N\hat{\Psi}_{K}^{2}$ maintains the nominal significance level well under $H_{0}$, and demonstrates high power under the alternatives unless $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$.
-The data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ using the approximated null distribution $H(x)$ performs reliably in controlling the Type  error and achieves excellent power performance.
-Table 5: Empirical rejection rates in the second experiment, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
+Figure 3: The sample means and error bars of $\hat{K}$ in Experiment III
+Figure 4: The sample means and error bars of $\hat{K}$ in Experiment IV
+In summary, the experiments demonstrate that the proposed tests in Section 3 maintain the nominal significance level well under the null, and exhibit high power under the alternatives, except when $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$.
+The data-driven tests with the approximated null distribution $H(x)$ reliably control the Type I error and achieve excellent power performance.
+Table 11: Empirical rejection rates in Experiment IV, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
 $H_{0}$ $H_{1}$
@@
 $150$ 0.056 0.064 0.054 0.058 0.060 0.054 0.052 0.028 1 1 1 1 1 1
-Table 6: Empirical frequency of $\hat{K}$ under $H_{0}$ in the second experiment
+Table 12: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment IV
 $m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
@@
 $150$ 0.998 0.002 0 0 0
-Table 7: Empirical frequency of $\hat{K}$ under $H_{1}$ in the second experiment
+Table 13: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment IV
 $m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
@@
 In this section, we briefly discuss smooth tests for normality of other ANOVA models beyond those in Section 3. The testing approach above is also applicable to these models.
-6.1 One-way fixed effects model with different group means and variances
-We first consider the one-way ANOVA model with different group means and different group variances, which is given by
+6.1 One-way fixed effects model with heterogeneous means and variances
+We first consider the one-way ANOVA model with potentially different group means and variances, given by
 $Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J.$ (6.1)
-Here $\{\mu_{j}\}_{j=1}^{J},\{\sigma_{j}\}_{j=1}^{J}$ represent the potentially different group means and group standard deviations, respectively.
-Due to the heterogeneity across groups in model (6.1), the estimation of each group mean and group variance only depends on observations in the corresponding group, in the sense that $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij},\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ for $j=1,\ldots,J$.
+Here $\{\mu_{j}\}_{j=1}^{J}$ and $\{\sigma_{j}\}_{j=1}^{J}$ denote the group-specific means and standard deviations.
+Owing to heterogeneity across groups, the estimators of each group mean and variance rely solely on the corresponding observations, in the sense that $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij},\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ for $j=1,\ldots,J$.
 Accordingly, we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}_{j}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
-To implement the smooth test, one can follow the steps in Section 3 and derive a formula like regarding the estimation effects in $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij}),k=1,\ldots,K$.
-Then the test statistic is constructed as a quadratic form of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$, normalized by the inverse of its asymptotic covariance matrix.
-It is expected that an asymptotic $\chi^{2}$ test can be obtained under mild conditions for model (6.1), which leads to a unified testing framework for one-way fixed effects models.
+To implement the smooth test, one may follow the procedure in Section 3 and derive an expression for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$, that accounts for the estimation effects.
+The test statistic is then constructed as a quadratic form of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$, normalized by the inverse of its asymptotic covariance matrix.
+Under mild conditions, an asymptotic $\chi^{2}$ test can be established for model (6.1), thereby extending the unified framework to this case.
 6.2 One-way random effects model
-Now we turn to the one-way random effects model. The observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are generated from following equation:
+Next, we consider the one-way random effects model:
 $Y_{ij}=\mu+A_{j}+\varepsilon_{ij}=\mu+A_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (6.2)
-where the $\{A_{j}\}_{j=1}^{J}$ are i.i.d.
-random effects independent of the random errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$, with $\mathrm{E}[A_{j}]=0$ and $\operatorname{Var}[A_{j}]=\sigma_{a}^{2}$.
-In this setting, we estimate $\sigma^{2}$ by
+where the random effects $\{A_{j}\}_{j=1}^{J}$ are i.i.d.
+with $\mathrm{E}[A_{j}]=0$, $\operatorname{Var}[A_{j}]=\sigma_{a}^{2}$, and are independent of the errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$.
+We estimate $\sigma^{2}$ by
 $\hat{\sigma}^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\bar{Y}_{\cdot j}\right)^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\hat{\varepsilon}_{ij}^{2},$
 where $\bar{Y}_{\cdot j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\varepsilon}_{ij}=Y_{ij}-\bar{Y}_{\cdot j}$.
-Note that $\bar{Y}_{\cdot j}$ coincides with $\hat{\mu}_{j}$ in Section 3.2, leading to an identical expression for $\hat{\sigma}^{2}$ here as that in Section 3.2, too.
-Therefore, the methodology in Section 3.2 can be directly applied to test the normality of $e_{ij}$ in (6.2), since the random effects $A_{j}$ are effectively eliminated by within-group centering via $Y_{ij}-\bar{Y}_{\cdot j}$.
-6.3 Two-way models
-Next, we extend the discussion to two-way ANOVA models. Suppose the observations
-$\{Y_{ij_{1}j_{2}}\}_{i=1,j_{1}=1,j_{2}=1}^{N_{j_{1}j_{2}},J_{1},J_{2}}$ are generated from the two-way fixed effects model:
-$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\left(\mu+\alpha_{j_{1}}+\beta_{j_{2}}+\gamma_{j_{1}j_{2}}\right)+\sigma e_{ij_{1}j_{2}},$ (6.3)
+Note that $\bar{Y}_{\cdot j}$ coincides with $\hat{\mu}_{j}$ in Section 3.2, so the expression for $\hat{\sigma}^{2}$ is identical to that in Section 3.2.
+Consequently, the methodology in Section 3.2 can be directly applied to test the normality of $e_{ij}$ in (6.2), since the random effects $A_{j}$ are removed by within-group centering.
+6.3 Two-way ANOVA models
+We now extend the discussion to two-way ANOVA models. Consider the two-way fixed effects model:
+$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\big(\mu+\alpha_{j_{1}}+\beta_{j_{2}}+\gamma_{j_{1}j_{2}}\big)+\sigma e_{ij_{1}j_{2}},$ (6.3)
 subject to the identifiability constraints
@@
 The theoretical framework in Section 3 remains valid in this setting.
-Similarly, for the two-way random effects model:
-$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\left(\mu+A_{j_{1}}+B_{j_{2}}+D_{j_{1}j_{2}}\right)+\sigma e_{ij_{1}j_{2}},$ (6.4)
+Similarly, consider the two-way random effects model:
+$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\big(\mu+A_{j_{1}}+B_{j_{2}}+D_{j_{1}j_{2}}\big)+\sigma e_{ij_{1}j_{2}},$ (6.4)
 where the i.i.d.
-zero-mean random effects $\{A_{j_{1}}\}_{j_{1}=1}^{J_{1}}$, $\{B_{j_{2}}\}_{j_{2}=1}^{J_{2}}$, $\{D_{j_{1}j_{2}}\}_{j_{1}=1,j_{2}=1}^{J_{1},J_{2}}$, and the random errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$, are mutually independent.
-To test the normality of $e_{ij}$ in (6.4), one can mimic the strategy used in Section 6.2 to eliminate the random effects first through within-cell centering, and then construct the test statistic based on the resulting residuals.
+random effects $\{A_{j_{1}}\}$, $\{B_{j_{2}}\}$, $\{D_{j_{1}j_{2}}\}$ and the errors $\{e_{ij}\}$ are mutually independent.
+To test the normality of $e_{ij}$ in (6.4), one can apply the strategy of Section 6.2, first eliminating the random effects by within-cell centering and then constructing the test statistic based on the residuals.
 To conclude, the methodology proposed in Section 3 can be seamlessly extended to more general ANOVA settings.
-This unified framework offers a flexible and effective approach to testing for normality in a wide range of linear models.
+This unified framework offers a flexible and effective approach to testing for normality across diverse linear models.
 7 Concluding remarks
 In this paper, we establish Neyman’s smooth tests for assessing the normality assumption in ANOVA models.
-For the proposed tests in three types of one-way fixed effects models, we demonstrate the asymptotic properties and validate the finite-sample performance through extensive simulations.
-Our approach provides a reliable tool for validating the normality assumption in diverse ANOVA settings.
+For three types of one-way fixed effects models, we derive the asymptotic properties of the proposed tests and validate their finite-sample performance through extensive simulations.
+Our results provide a rigorous and practical tool for evaluating normality in a broad range of ANOVA settings.
 Several directions remain open for future research.
-First, although we have developed tests within the ANOVA framework, the true data-generating mechanism may not always be known a priori.
-A promising approach is to conduct preliminary structural tests to identify the most suitable model before applying the proposed procedures.
-Second, the independence assumption on random errors can be relaxed to enable further theoretical analysis.
-For example, repeated clinical trials frequently exhibit autocorrelated errors (Ma et al.,, 2012; Schober and Vetter,, 2018; Langenberg et al.,, 2022).
-In such cases, our smooth test framework could potentially be adapted by incorporating estimated correlation structures.
-Finally, for the data-driven selection rule, it is intriguing to explore the case where the upper bound $D=D(N)\to\infty$ as $N\to\infty$ from the theoretical perspective.
+First, while the proposed tests are developed for specific ANOVA models, the true data-generating mechanism may not be known in practice.
+A natural extension is to combine our procedures with preliminary structural tests to identify the most appropriate ANOVA specification before applying the proposed methodology.
+Second, the assumption of independent random errors can be relaxed.
+For example, repeated clinical trials often involve correlated errors (Ma et al.,, 2012; Schober and Vetter,, 2018; Langenberg et al.,, 2022).
+In such cases, our smooth test framework may be adapted by incorporating estimated correlation structures.
+Finally, for the data-driven selection rule, it is of theoretical interest to explore the scenario where the upper bound $D=D(N)$ diverges with $N$, that is, $D\to\infty$ as $N\to\infty$.
 Appendix A Appendix
-We provide the proof of our main theoretical results and some practical issues in this Appendix.
+We provide the proof of our main theoretical results, along with some additional numerical results, in this Appendix.
 The symbol $\lesssim$” means that the left side is bounded by a positive constant times the right side.
@@
 The proof is similar to that of Theorems 1, 2, and 3.
-Note that in the case of different means and same variance we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$.
+Note that in the case of different means and same variance, we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$.
 To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ without distinction of the null and alternatives.
@@
 $\operatorname{Var}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\right]=O\left(\frac{1}{N}\right).$
-Now we start from (A.3). In the case of the same mean and different variances, (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
+Now we start from (A.3). In the case of the same mean and different variances, (A.3) and (A.4) still hold. The difference lies in the fact that the first term of (A.4) should be further expressed as
 $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(\frac{\varepsilon_{ij}}{\sigma_{j}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)$
@@
 which means the test is consistent against the alternatives given by (4.2). $\square$
-Intuition behind the revised limiting null distribution+ (4.3).
-We mainly follow the arguments in Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007), and provide a heuristic derivation of the approximation $H(x)$ for the first case since the latter two cases can be derived analogously.
+Intuition behind the revised limiting null distribution (4.3).
+We primarily follow the arguments in Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), and Kraus, (2007), and provide a heuristic derivation of the approximation $H(x)$ for the first case, as the latter two cases can be derived analogously.
 Note that
-$\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x\right)\\ =\mathrm{P}\left(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1\right)+\mathrm{P}\left(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2\right)+\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3\right).$ (A.30)
-The third term on the right-hand side of (A.30), $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3)$, can be neglected under $H_{0}$ (Kallenberg and Ledwina, 1995b, ).
+$\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x\right)=$ $\mathrm{P}\left(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1\right)+\mathrm{P}\left(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2\right)$
+$+\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3\right).$ (A.30)
+The third term on the right-hand side of (A), $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3)$, can be neglected under $H_{0}$ (Kallenberg and Ledwina, 1995b, ).
 The event $\{\hat{K}=1\}$ is approximated by $\{N\hat{\Psi}^{2}_{1}-\log N\geq N\hat{\Psi}^{2}_{2}-2\log N\}=\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\hat{K}=2\}$ is approximated by $\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})>\log N\}$.
@@
 by $N\hat{\Psi}^{2}_{1}\geq 0$. Thus, we have the approximation
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)\approx\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]\equiv H(x),\quad x\leq\log N,$
+$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)$
+$\approx$ $\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]\equiv H(x),\quad x\leq\log N,$
 from the asymptotic $\chi_{1}^{2}$ distribution and asymptotic independence of $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$. For $x\geq 2\log N$, we have
@@
 As $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ is approximately $\chi_{1}^{2}$ distributed, we obtain
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)\approx 2\left(1-\Phi\left(\sqrt{\log N}\right)\right)\approx 2\frac{\phi\left(\sqrt{\log N}\right)}{\sqrt{\log N}}=\sqrt{\frac{2}{\pi N\log N}},$
+$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
+$\approx$ $2\left(1-\Phi\left(\sqrt{\log N}\right)\right)\approx 2\frac{\phi\left(\sqrt{\log N}\right)}{\sqrt{\log N}}=\sqrt{\frac{2}{\pi N\log N}},$
 by the fact that $t\phi(t)/(t+1)\leq 1-\Phi(t)\leq\phi(t)/t$ for $t\to\infty$. Besides,
@@
 This completes the derivation of the approximation (4.3) for the null distribution of the test statistic $N\hat{\Psi}^{2}_{\hat{K}}$. $\square$
-Practical issues in simulation studies.
+Implement issues in simulation studies.
 In our simulation study, we adopt the orthonormal Legendre polynomials on $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$.
@@
 $c_{2k}=\sqrt{2k+1}\,\sum_{j=0}^{[k/2]}\left(\frac{-1}{4}\right)^{j}\binom{k}{j}\binom{2(k-j)}{k}\,\int_{-1/2}^{1/2}x^{k-2j}\,\left[\Phi^{-1}\left(x+\frac{1}{2}\right)\right]^{2}\,\mathrm{d}z.$
-Note that $c_{2k}=0$ when $k$ is odd and $c_{1k}=0$ when $k$ is even. Therefore, in practice, we only compute the non-zero elements based on the above two formulae for $k=1,\ldots,K$ with $K\leq 5$.
+Note that $c_{2k}=0$ when $k$ is odd and $c_{1k}=0$ when $k$ is even.
+Therefore, in our experiments, we only compute the non-zero elements based on the above two formulae for $k=1,\ldots,K$ with $K\leq 5$.
+Additional simulation results.
+The results of Experiment III are reported in Tables 14, 15, 16 and Figure 5.
+The main findings are broadly consistent with those in Experiment III, with two additional observations.
+First, Table 14 indicates that the limiting $\chi^{2}_{1}$ distribution for the data-driven test yields valid control of the Type I error.
+Second, the empirical distribution of $\hat{K}$ exhibits increased stability under both the null and the alternative hypotheses, with instances where it degenerates to a single value with probability one.
+Both phenomena are due to the increase in total sample size resulting from the larger number of groups, which is consistent with the theoretical results.
+Table 14: Empirical rejection rates under $H_{0}$ in Experiment III, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
+The empirical power under $H_{1}$ equals 1 across all settings.
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
+10 0.028 0.046 0.062 0.036 0.054 0.060 0.048
+20 0.046 0.066 0.078 0.052 0.068 0.056 0.052
+30 0.050 0.056 0.044 0.040 0.048 0.044 0.042
+40 0.054 0.054 0.042 0.040 0.044 0.058 0.052
+50 0.052 0.062 0.052 0.056 0.054 0.048 0.042
+60 0.062 0.050 0.060 0.048 0.048 0.052 0.046
+70 0.068 0.052 0.052 0.054 0.044 0.058 0.054
+80 0.052 0.034 0.044 0.062 0.052 0.048 0.046
+90 0.048 0.048 0.034 0.050 0.058 0.042 0.040
+100 0.064 0.048 0.052 0.048 0.052 0.058 0.054
+110 0.044 0.048 0.058 0.058 0.032 0.060 0.056
+120 0.052 0.058 0.066 0.044 0.052 0.080 0.078
+130 0.046 0.050 0.038 0.058 0.062 0.058 0.052
+140 0.058 0.046 0.032 0.054 0.046 0.056 0.050
+150 0.046 0.040 0.060 0.050 0.054 0.038 0.038
+Table 15: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment III
+$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
+10 0.994 0.006 0 0 0
+20 1 0 0 0 0
+30 0.992 0.008 0 0 0
+40 0.994 0.006 0 0 0
+50 0.988 0.012 0 0 0
+60 0.994 0.006 0 0 0
+70 0.998 0.002 0 0 0
+80 0.992 0.008 0 0 0
+90 0.996 0.004 0 0 0
+100 0.994 0.006 0 0 0
+110 0.996 0.004 0 0 0
+120 0.998 0.002 0 0 0
+130 0.990 0.010 0 0 0
+140 0.994 0.006 0 0 0
+150 0.998 0.002 0 0 0
+Table 16: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment III
+$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
+10 0 0 0 0.120 0.880
+20 0 0 0 0.024 0.976
+30 0 0 0 0 1
+40 0 0 0 0 1
+50 0 0 0 0 1
+60 0 0 0 0 1
+70 0 0 0 0 1
+80 0 0 0 0 1
+90 0 0 0 0 1
+100 0 0 0 0 1
+110 0 0 0 0 1
+120 0 0 0 0 1
+130 0 0 0 0 1
+140 0 0 0 0 1
+150 0 0 0 0 1
+Figure 5: The sample means and error bars of $\hat{K}$ in Experiment III

### v4

@@
-Smooth Tests for Normality in ANOVA
+Data-driven Smooth Tests for Normality in ANOVA When the Number of Groups is Large
 Abstract
-The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models, yet it is seldom subjected to formal testing in practice.
-In this paper, we develop Neyman’s smooth tests for assessing normality in a broad class of ANOVA models.
-The proposed test statistics are constructed via the Gaussian probability integral transformation of ANOVA residuals and are shown to follow an asymptotic Chi-square distribution under the null hypothesis, with degrees of freedom determined by the dimension of the smooth model.
-We further propose a data-driven selection of the model dimension based on a modified Schwarz’s criterion.
-Monte Carlo simulations demonstrate that the tests maintain the nominal size and achieve high power against a wide range of alternatives.
-Our framework thus provides a systematic and effective tool for formally validating the normality assumption in ANOVA models.
-Keywords: ANOVA; Estimation effect; Normality; Schwarz’s selection rule; Smooth test.
+The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models.
+However, it is rarely subjected to formal testing in practice, and theoretically justified procedures are largely unavailable, especially when the number of groups diverges.
+In this paper, we develop Neyman’s smooth tests for assessing normality in a broad class of ANOVA models, allowing the number of groups to diverge.
+The proposed test statistics are constructed via the Gaussian probability integral transformation of ANOVA residuals.
+We show that using residuals induces non-negligible parameter estimation effects, whose structure depends on the underlying ANOVA model and plays a crucial role in shaping the form of the test statistics and their asymptotic behavior.
+Under the null hypothesis of normality, the resulting statistics follow an asymptotic Chi-square distribution, with degrees of freedom determined by the order of the smooth test (i.e., the number of components included in the smooth test).
+We further propose a modified Schwarz’s selection rule to automatically determine the order, thereby yielding fully data-driven smooth tests that require no additional tuning parameters.
+Simulation studies and a real-data example indicate that the proposed tests perform well in practice and are readily applicable.
+Keywords: ANOVA; Estimation effects; Normality; Schwarz’s selection rule; Smooth tests
 1 Introduction
-Analysis of Variance (ANOVA) is a fundamental and widely used tool in both exploratory and confirmatory data analysis (Gelman,, 2005), particularly for comparing group means and assessing the significance of factors in experimental designs.
-The theory of ANOVA has been well established in the literature; see, for example, Scheffé, (1959), Miller Jr, (1997), Wickens and Keppel, (2004), Dean et al., (2017), Hirotsu, (2017), and Montgomery, (2017).
+Analysis of Variance (ANOVA) is a fundamental and widely used tool in both exploratory and confirmatory data analysis (gelman2005analysis), particularly for comparing group means and assessing the significance of factors in experimental designs.
+The theory of ANOVA has been well established in the literature; see, for example, scheffe1959anova, miller1997beyond, wickens2004design, dean2017design, hirotsu2017advanced, and montgomery2017design.
 A standard assumption in ANOVA is that the random errors are normally distributed.
 Violations of this assumption may invalidate normal-based inference.
-From an estimation perspective, departures from normality undermine the validity of variance component estimators, since the classical variance formulae rely on mean squares of random effects being constant multiples of Chi-square variables (Scheffé,, 1959).
-From a testing perspective, the $F$-test in ANOVA is sensitive to nonnormality, especially when group sizes are unequal (Ali and Sharma,, 1996).
-The impact of nonnormality on the size and power of the $F$-test has been extensively studied (Pearson,, 1931; Gayen,, 1950; David and Johnson, 1951a, ; David and Johnson, 1951b, ; Srivastava,, 1959; Atiqullah,, 1962; Tiku,, 1964; Donaldson,, 1968; Tiku,, 1971; Janic-Wróblewska and Ledwina,, 2000).
+From an estimation perspective, departures from normality undermine the validity of variance component estimators, since the classical variance formulae rely on mean squares of random effects being constant multiples of Chi-square variables (scheffe1959anova).
+From a testing perspective, the $F$-test in ANOVA is sensitive to nonnormality, especially when group sample sizes are unbalanced (ali1996robustness) or the number of groups is large (akritas2000asymptotics; akritas2004heteroscedastic; wang2006two).
+The impact of nonnormality on the size and power of the $F$-test has been extensively studied (pearson1931analysis; gayen1950distribution; david1951method; david1951effect; srivastava1959effect; atiqullah1962estimation; tiku1964approximating; donaldson1968robustness; tiku1971power; akritas2000asymptotics; janic2000data).
 These considerations highlight the importance of rigorously evaluating the normality assumption in ANOVA models.
-Although a variety of normality tests have been proposed for observed data and linear models, relatively little attention has been paid to systematically assessing normality in ANOVA.
-A straightforward diagnostic approach is to use the normal probability plots of ANOVA residuals (Dean et al.,, 2017; Hirotsu,, 2017), which are intuitive but lack theoretical justification.
-To our knowledge, formal hypothesis testing procedures designed explicitly for ANOVA are scarce in the literature.
-Most existing methods simply apply classical normality tests to ANOVA residuals, see, for example, Bonett and Woodward, (1990), Bonett and Seier, (2002), and Hwang and Wei, (2006).
-While applicable in practice, such approaches do not provide a rigorous assessment of normality within the ANOVA framework and fail to account for the structural constraints inherent in ANOVA models.
-In this article, following the spirit of Neyman, (1937), we propose a unified smooth test framework for assessing normality in various types of ANOVA models.
-Neyman’s smooth test is widely used in diverse scientific fields due to its theoretical soundness and practical effectiveness, with proven power against a broad class of alternatives; see, e.g., Bera and Ghosh, (2002), Thas, (2010), and Lehmann and Romano, (2022).
-Specifically, we reformulate the normality testing problem as one of uniformity via the Gaussian probability integral transform (PIT), and construct test statistics based on the Gaussian PIT of ANOVA residuals.
-Under mild conditions, the proposed statistics are asymptotically Chi-square distributed under the null hypothesis.
-We further analyze the power properties under both fixed and local alternatives.
-A modified Schwarz’s selection rule is proposed to determine the direction of tests.
-Monte Carlo experiments demonstrate the good performance of our methodology in finite samples.
-The main contributions of this paper are threefold.
-First, the proposed asymptotic Chi-square tests are simple to implement: critical values are obtained directly from the limiting distributions without resorting to resampling.
-Second, we systematically investigate the properties of the test statistics in three types of one-way fixed effects models, potentially with a diverging number of groups, within a unified asymptotic framework.
-Our analysis clarifies how parameter estimation affects the test statistics: heterogeneity in group means and variances alters their explicit form and the corresponding regularity conditions, but the convergence rate depends only on the total sample size under both null and alternative hypotheses.
-Third, we provide theoretical justification for data-driven smooth tests in the ANOVA setting, including a revised limiting null distribution for the corresponding statistics in finite samples.
+Although a variety of normality tests have been proposed for observed data and linear models (see, for example, dagostino1986goodness and bonett2002test), relatively little attention has been paid to the systematic assessment of normality in ANOVA.
+A straightforward diagnostic approach is to use the normal probability plots of ANOVA residuals (dean2017design; hirotsu2017advanced), which is intuitive but lacks theoretical justification.
+To our knowledge, formal normality testing procedures explicitly designed for ANOVA are scarce in the literature.
+Most existing methods instead apply classical normality tests—such as the Shapiro–Wilk test (shapiro1965analysis; shapiro1972approximate), the Jarque–Bera test (jarque1987test), and the Kolmogorov–Smirnov test (in particular, the Lilliefors version) (lilliefors1967kolmogorov)—directly to ANOVA residuals; see, for example, bonett1990testing; bonett2002test; hwang2006novel.
+While convenient in practice, such approaches suffer from two fundamental limitations.
+First, they ignore the estimation effect induced by fitting the ANOVA model, which may alter the distributional properties of the residuals.
+Second, they lack theoretical guarantees, especially in ANOVA models with a diverging number of groups, under which these procedures may become invalid.
+Consequently, they do not provide a rigorous assessment of normality within the ANOVA framework and fail to account for the inherent structural constraints of ANOVA.
+In this article, following the spirit of neyman1937smooth, we propose a unified Neyman’s smooth test framework for assessing the normality of random errors in various types of ANOVA models where the number of groups is allowed to diverge.
+Neyman’s smooth tests are widely used in diverse scientific fields due to their theoretical soundness and practical effectiveness; see, e.g., bera_neymans_2002, Chapters 4 and 10 of thasComparingDistributions2010, and Section 16.4 of Lehmann_Romano_2022.
+Specifically, we reformulate the normality testing problem as a uniformity problem via the Gaussian probability integral transform (PIT) and construct test statistics based on the Gaussian PIT of the ANOVA residuals.
+The use of residuals introduces parameter estimation effects whose structure depends on the underlying model and is therefore intrinsic to the ANOVA framework.
+Consequently, the resulting statistic takes the form of a quadratic expression involving the inverse of a covariance matrix, whose structure is determined by these estimation effects.
+Given mild conditions, the proposed test statistics are asymptotically Chi-square distributed under the null hypothesis of normality.
+We also analyze the power properties under both fixed and local alternatives.
+Furthermore, a modified Schwarz’s selection rule is proposed to determine the test directions (i.e., the order of the test), yielding fully data-driven smooth tests that do not require additional tuning parameters.
+Numerical results demonstrate the good performance of our proposed methodology in finite samples.
+The main contributions of this paper are twofold.
+First, we systematically investigate the effects of parameter estimation on ANOVA residuals-based normality testing.
+We show that heterogeneity in group means and variances alters the explicit form of the test statistics and the associated regularity conditions, while the convergence rate depends solely on the total sample size under both the null and alternative hypotheses.
+Second, we provide rigorous theoretical justification for data-driven smooth tests in the ANOVA setting.
+In particular, we establish a revised limiting null distribution for the corresponding statistics that remains valid in finite samples.
+As a result, the proposed tests are simple to implement, with critical values obtained directly from the limiting distributions without resorting to resampling.
+Overall, the proposed framework accommodates settings with a potentially diverging number of groups and thus overcomes key limitations of many classical methods.
 The remainder of this paper is organized as follows.
@@
 Section 5 reports simulation results.
-Section 6 discusses further extensions.
+Section 6 provides an empirical application.
 Section 7 concludes.
-Proofs and additional numerical results are provided in Appendix A.
+Proofs and additional numerical results are provided in the Online Appendix.
 2 The testing framework
-Our problem of interest is to determine whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$, that is, $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ for some $\sigma>0$.
+Our problem of interest is to assess whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$, that is, $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ for some $\sigma>0$.
 Equivalently, this can be formulated as testing the null hypothesis about the standardized random error $H_{0}:e\sim\mathcal{N}(0,1)$ against alternatives of nonnormality.
@@
 Under the null hypothesis of normality, the CDF of the transformed random variable $Z$ is given by
-$G(z)=\mathrm{P}(Z\leq z)=\mathrm{P}\left(e\leq\Phi^{-1}(z)\right)=\Phi(\Phi^{-1}(z))=z.$
+$G(z)=\mathbb{P}(Z\leq z)=\mathbb{P}\left(e\leq\Phi^{-1}(z)\right)=\Phi(\Phi^{-1}(z))=z.$
 Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1,z\in[0,1]$, that is, $Z\sim\mathcal{U}[0,1]$, where $\mathcal{U}[0,1]$ denotes the uniform distribution on the interval $[0,1]$.
 Under the alternative, the density $g(z)$ deviates from unity.
-This distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ provides the foundation for Neyman’s smooth test.
-In particular, Neyman, (1937) introduced the following smooth alternative to the uniform density:
+This distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ forms the basis of Neyman’s smooth test.
+In particular, neyman1937smooth introduced the following smooth alternative to the uniform density:
 $h\left(z\right)=c\left(\bm{\theta}_{K}\right)\exp\left[\sum_{k=1}^{K}\theta_{k}\pi_{k}\left(z\right)\right],\quad 0\leq z\leq 1,$ (2.1)
-where $c(\bm{\theta}_{K})$ is a a normalizing constant depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ denotes an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
+where $c(\bm{\theta}_{K})$ is a normalizing constant depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ denotes an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
 […]
 The null hypothesis $Z\sim\mathcal{U}[0,1]$ can thus be assessed by testing $\theta_{1}=\theta_{2}=\ldots=\theta_{K}=0$ in (2.1).
-If independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ are available and $\sigma$ were known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
+If independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ are available and $\sigma$ is also known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
 $\Psi_{K}^{2}=\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}(Z_{i})\right)^{2},$ (2.3)
@@
 However, in practice, the sequence $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown.
-We instead rely on approximations $\{\hat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and an estimator $\hat{\sigma}$ for $\sigma$ (e.g., residuals and the sample standard deviation, respectively).
+We instead rely on approximations $\{\widehat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and an estimator $\widehat{\sigma}$ for $\sigma$ (e.g., residuals and the sample standard deviation, respectively).
 Accordingly, we consider the following approximation for $Z_{i}$:
-$\hat{Z}_{i}=\Phi(\hat{e}_{i})=\Phi\left(\frac{\hat{\varepsilon}_{i}}{\hat{\sigma}}\right),\quad i=1,\ldots,n,$
+$\widehat{Z}_{i}=\Phi(\widehat{e}_{i})=\Phi\left(\frac{\widehat{\varepsilon}_{i}}{\widehat{\sigma}}\right),\quad i=1,\ldots,n,$
 and construct a feasible test statistic of the following form:
-$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\hat{Z}_{i}\right)\right)^{2}.$ (2.4)
-The random variables $\{\hat{Z}_{i}\}_{i=1}^{n}$ are not i.i.d.
-any longer due to estimation effects caused by $\hat{\varepsilon}_{i}$ and $\hat{\sigma}$.
-Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\hat{Z}_{i})$.
-Compared with $\Psi_{K}^{2}$ in (2.3), the statistic based on $\hat{Z}_{i}$ suffers from non-negligible estimation effects, as demonstrated in the theoretical results below.
+$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\widehat{Z}_{i}\right)\right)^{2}.$ (2.4)
+The random variables $\{\widehat{Z}_{i}\}_{i=1}^{n}$ are not i.i.d.
+any longer due to estimation effects caused by $\widehat{\varepsilon}_{i}$ and $\widehat{\sigma}$.
+Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\widehat{Z}_{i})$.
+Compared with $\Psi_{K}^{2}$ in (2.3), the statistic based on $\widehat{Z}_{i}$ suffers from non-negligible estimation effects, as demonstrated in the theoretical results below.
 Consequently, the presence of estimation effects invalidates the form of the “naive” statistic (2.4) and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution.
-This motivates a detailed analysis of estimation effects, leading to the correct quadratic-form test statistic.
+This motivates a detailed analysis of the effects of estimation that yields the correct quadratic-form test statistic.
 Another important issue concerns sample sizes in ANOVA models.
@@
 This simplified setting captures the core methodological ideas and facilitates the derivation of key asymptotic properties.
-In contrast, the complete ANOVA setting involves additional complications arising from multiple sample sizes and the associated indexing structure, which demand more refined analysis.
-From a practical standpoint, our method is further designed to accommodate scenarios where the number of groups increases with the total sample size.
+In contrast, the complete ANOVA setting introduces additional complications due to multiple sample sizes and the associated indexing structure, necessitating a more refined analysis.
+From a practical standpoint, our method is designed to accommodate scenarios in which the number of groups increases with the total sample size.
 These issues will be systematically investigated in subsequent analysis of specific ANOVA models.
 3 Testing for normality in one-way fixed effects models
-In this section, we develop smooth tests for normality in three types of one-way fixed effects models, where each model assumes equality in at least one of the group means or the group variances.
-The testing procedure and its theoretical properties are investigated separately for each case.
-Let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observations of interest coming from $J$ groups, with $N_{j}$ being the sample sizes in group $j$ for $1\leq j\leq J$.
-The total sample size is defined as $N=\sum_{j=1}^{J}N_{j}$.
-In our theoretical framework, the number of groups, $J$, is allowed to diverge as $N$ increases.
-3.1 Same mean and same variance
-We begin with the ANOVA models where both the group means and group variances are equal. Specifically, consider the model
+In this section, we develop smooth tests for normality in three types of one-way fixed effects models, each imposing homogeneity on either the group means or the group variances.
+The testing procedures and their theoretical properties are studied separately for each case.
+Formally, let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observed responses from $J$ groups, where $N_{j}$ is the sample size of group $j$ for $1\leq j\leq J$.
+The total sample size is $N=\sum_{j=1}^{J}N_{j}$.
+In our asymptotic framework, the number of groups $J$ is allowed to diverge as $N\to\infty$.
+3.1 Common group means and common group variances
+We begin with ANOVA models in which both the group means and group variances are equal. Specifically, consider the model
 $Y_{ij}=\mu+\varepsilon_{ij}=\mu+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.1)
@@
 We also define $Z_{ij}=\Phi(e_{ij})$.
-The parameters $\mu$ and $\sigma^{2}$ in (3.1) are estimated by the sample mean $\hat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}\right)^{2}$, respectively.
-Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu})/\hat{\sigma}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
-To derive the feasible test statistic and its properties, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ for $k=1,\ldots,K$.
+The parameters $\mu$ and $\sigma^{2}$ in model (3.1) are estimated by the sample mean $\widehat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\widehat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\widehat{\mu}\right)^{2}$, respectively.
+Define $\widehat{\varepsilon}_{ij}=Y_{ij}-\widehat{\mu}$, $\widehat{e}_{ij}=\widehat{\varepsilon}_{ij}/\widehat{\sigma}=(Y_{ij}-\widehat{\mu})/\widehat{\sigma}$, and $\widehat{Z}_{ij}=\Phi(\widehat{e}_{ij})$.
+To derive the feasible test statistic and its properties, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ for $k=1,\ldots,K$.
 For this purpose, we impose the following two assumptions.
@@
 For $k=1,\ldots,K$, $\pi_{k}(\cdot)$ are twice continuously differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
-The two mild assumptions are in line with similar ones adopted in the literature of Neyman’s smooth tests, including Bera et al., (2013) and Song and Xiao, (2022).
+The two mild assumptions are in line with similar ones adopted in the literature of Neyman’s smooth tests, including bera2013smooth and song2022smooth.
 For subsequent analysis, we introduce two constants:
@@
 Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 as $N\to\infty$.
-Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ is equivalent to the sum of three terms after neglecting the higher-order term: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$.
-The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$.
+Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ is equivalent to the sum of three terms after neglecting the higher-order ones: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$.
+The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$, $k=1,\ldots,K$.
 With the assistance of Theorem 1, and by invoking the central limit theorem (CLT), we obtain the asymptotic normality of the $K$-dimensional vector
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$ (3.2)
-$\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Sigma}_{K}\right),$
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\widehat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\widehat{Z}_{ij}\right)\right]^{\top}$ (3.2)
+$\overset{d}{\to}\mathcal{N}_{K}\left(\bm{0},\bm{\Sigma}_{K}\right),$
 where the asymptotic covariance matrix $\bm{\Sigma}_{K}=\left(\sigma_{kl}\right)_{K\times K}$ is given by
-$\sigma_{kl}$ $=\mathrm{E}\left[\left(\pi_{k}(Z)-c_{1k}e-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right)\left(\pi_{l}(Z)-c_{1l}e-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right)\right]$
+$\sigma_{kl}$ $=\mathbb{E}\left[\left(\pi_{k}(Z)-c_{1k}e-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right)\left(\pi_{l}(Z)-c_{1l}e-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right)\right]$
 $=\delta_{kl}-c_{1k}c_{1l}-\frac{1}{2}c_{2k}c_{2l}.$
 Based on this result, we define the feasible smooth test statistic as
-$\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\bm{\Sigma}_{K}^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.3)
-The following corollary establishes the limiting null distribution of $\hat{\Psi}_{K}^{2}$ defined in (3.3).
+$\widehat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)^{\top}\bm{\Sigma}_{K}^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right).$ (3.3)
+The following corollary establishes the limiting null distribution of $\widehat{\Psi}_{K}^{2}$ defined in (3.3).
 Corollary 1.
 Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$,
-$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2},$
+$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2},$
 as $N\to\infty$.
-Corollary 1 establishes an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\hat{\Psi}_{K}^{2}$.
+Corollary 1 establishes an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\widehat{\Psi}_{K}^{2}$.
 Given the asymptotic significance level of $\alpha$, we reject the null hypothesis if
-$N\hat{\Psi}_{K}^{2}>\chi_{K,1-\alpha}^{2},$
+$N\widehat{\Psi}_{K}^{2}>\chi_{K,1-\alpha}^{2},$
 where $\chi_{K,1-\alpha}^{2}$ denotes the $(1-\alpha)$-th quantile of the $\chi_{K}^{2}$ distribution.
-We now investigate the asymptotic behavior of $\hat{\Psi}_{K}^{2}$ under the fixed alternatives as well as under a Pitman-type sequence of local alternatives.
+We now investigate the asymptotic behavior of $\widehat{\Psi}_{K}^{2}$ under the fixed alternatives as well as under a Pitman-type sequence of local alternatives.
 For the fixed alternatives, we consider the following form:
@@
 Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed.
-Consequently, the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ differs from that under $H_{0}$.
+Consequently, the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ and $\widehat{\Psi}_{K}^{2}$ differs from that under $H_{0}$.
 To formally state the corresponding theoretical results under $H_{1}$, we introduce the following notations.
 Let
-$d_{1k}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)f(x)\mathrm{d}x,$
-$d_{2k}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)xf(x)\mathrm{d}x,$
-$d_{3k}=\mathrm{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)xf(x)\mathrm{d}x,$
+$d_{1k}=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)f(x)\mathrm{d}x,$
+$d_{2k}=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)xf(x)\mathrm{d}x,$
+$d_{3k}=\mathbb{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)xf(x)\mathrm{d}x,$
 and
-$d_{4k}=\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)x^{2}f(x)\mathrm{d}x,$
-for $k=1,\ldots,K$.
+$d_{4k}=\mathbb{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)x^{2}f(x)\mathrm{d}x,$
+for $k=1,\ldots,K$, where $\phi$ denotes the PDF of standard normal distribution.
 Note that under $H_{0}$, the constants $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ since $f(x)=\phi(x)$.
-The following theorem characterizes the asymptotic properties of $\hat{\Psi}_{K}^{2}$ under the alternative hypothesis $H_{1}$.
+The following theorem characterizes the asymptotic properties of $\widehat{\Psi}_{K}^{2}$ under the alternative hypothesis $H_{1}$.
 Theorem 2.
 Suppose Assumptions 3.1 and 3.2 hold. Then, under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.5)
-$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.5)
+$+\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 as $N\to\infty$. Furthermore,
-$\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K},$
+$\widehat{\Psi}_{K}^{2}\overset{p}{\to}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K},$
 and
-$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right),$ (3.6)
-where $\bm{a}_{K}\equiv(a_{1},\ldots,a_{K})^{\top}=(\mathrm{E}[\pi_{1}(Z)],\ldots,\mathrm{E}[\pi_{K}(Z)])^{\top}$ and $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ is given by
-$\xi_{kl}=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
-$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
-From Theorem 2, it follows that under the alternative hypothesis $H_{1}$ in (3.4), if there exists at least one $1\leq k\leq K$ such that
-$a_{k}\equiv\mathrm{E}[\pi_{k}(Z)]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))f(x)\mathrm{d}x=\int_{0}^{1}\pi_{k}(z)\frac{f(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z\neq 0,$
-then the smooth test statistic satisfies $N\hat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, implying that the asymptotic power of the test is $1$.
+$\sqrt{N}\left(\widehat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right),$ (3.6)
+where $\bm{a}_{K}\equiv(a_{1},\ldots,a_{K})^{\top}=(\mathbb{E}[\pi_{1}(Z)],\ldots,\mathbb{E}[\pi_{K}(Z)])^{\top}$ and $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ is given by
+$\xi_{kl}=$ $\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
+$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathbb{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathbb{E}\left[e^{4}\right]-1\right].$
+The decomposition (3.5) exhibits different properties from that in Theorem 1.
+Specifically, the first term on the right-hand side $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$ now has a nonzero expectation under $H_{1}$ given by (3.4), in contrast to its zero mean under $H_{0}$, which constitutes the primary source of power for the proposed test.
+The latter two terms, $d_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$ and $d_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$, represent the estimation effects of $\widehat{\mu}$ and $\widehat{\sigma}^{2}$ under $H_{1}$, respectively.
+Moreover, when $H_{0}$ is true, $d_{1k}=c_{1k}$ and $d_{2k}=c_{2k}$, then the decomposition (3.5) coincides with that in Theorem 1.
+From Theorem 2, it follows that under $H_{1}$ given by (3.4), if there exists at least one $1\leq k\leq K$ such that
+$a_{k}\equiv\mathbb{E}[\pi_{k}(Z)]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))f(x)\mathrm{d}x=\int_{0}^{1}\pi_{k}(z)\frac{f(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z\neq 0,$
+then the smooth test statistic satisfies $N\widehat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, implying that the asymptotic power of the test is $1$.
 Specifically, the power function is given by
@@
 $\Delta_{k}=\int_{0}^{1}\pi_{k}(z)\frac{q(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z,\quad k=1,\ldots,K.$
-The following theorem presents the theoretical properties of $N\hat{\Psi}_{K}^{2}$ under the local alternatives.
+The following theorem presents the theoretical properties of $N\widehat{\Psi}_{K}^{2}$ under the local alternatives.
 Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the nontrivial local power of the test.
@@
 Suppose Assumptions 3.1 and 3.2 hold. Then, under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.8)
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.8)
 $+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 as $N\to\infty$. Furthermore,
-$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
+$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
 where $\bm{\Delta}_{K}=(\Delta_{1},\ldots,\Delta_{K})^{\top}$, and $\chi^{2}_{K}\left(\tau\right)$ denotes the noncentral $\chi^{2}$ distribution with $K$ degrees of freedom and nonnegative noncentrality parameter $\tau$.
-It is worthwhile to mention that $N^{-1/2}$ is the critical order of magnitude for our test to achieve nontrivial power against the Pitman-type sequence of local alternatives.
-If $\delta_{N}=o(N^{-1/2})$, it would be impossible to distinguish the alternatives from the null.
-Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\hat{\Psi}_{K}^{2}$ attains nontrivial asymptotic power against the local alternatives because the noncentrality parameter $\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}>0$.
+The asymptotic decomposition (3.8) is closely related to those obtained under $H_{0}$ and $H_{1}$, while exhibiting a key difference.
+In particular, compared with the decomposition under $H_{0}$, the leading term $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$ becomes noncentered under $H_{1L}$, with mean shift $\mathbb{E}[\pi_{k}(Z)]=\delta_{N}\Delta_{k}$, which introduces an additional deterministic drift term of order $\delta_{N}$.
+The two estimation effect terms $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$ and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ coincide with those under $H_{0}$.
+When $\delta_{N}=N^{-1/2}$, the drift term $\delta_{N}\Delta_{k}$ is of the same order as the stochastic fluctuations, leading to a nondegenerate limit distribution.
+Consequently, the statistic $N\widehat{\Psi}_{K}^{2}$ converges to a noncentral Chi-square distribution.
+This also shows that $N^{-1/2}$ is the critical rate for detecting Pitman-type local alternatives: if $\delta_{N}=o(N^{-1/2})$, the alternatives become asymptotically indistinguishable from the null.
+Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\widehat{\Psi}_{K}^{2}$ attains nontrivial asymptotic power against the local alternatives because the noncentrality parameter $\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}>0$.
 Remark 1.
 We emphasize that the aforementioned smooth test is not “consistent” in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives.
-Only under alternatives where $\mathrm{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power converge to $1$.
-Otherwise, if $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathrm{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\hat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$.
-For example, let $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ so that $\mathrm{E}[e]=0$ and $\operatorname{Var}[e]=1$, and consider the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$.
+Only under alternatives where $\mathbb{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power converge to $1$.
+Otherwise, if $\mathbb{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathbb{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\widehat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$.
+For example, let $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ so that $\mathbb{E}[e]=0$ and $\operatorname{Var}[e]=1$, and consider the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$.
 For $Z=\Phi(e)$, we have
-$\mathrm{E}[Z]$ $=\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}\Phi(z)\mathrm{d}z$
+$\mathbb{E}[Z]$ $=\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}\Phi(z)\mathrm{d}z$
 $=\frac{1}{2\sqrt{3}}z\Phi(z)\Big|_{-\sqrt{3}}^{\sqrt{3}}-\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}z\mathrm{d}\Phi(z)$
@@
 $=\frac{1}{2},$
-which implies $\mathrm{E}[\pi_{1}(Z)]=0$.
-In fact, smooth tests are neither directional nor omnibus; that is, they maintain reasonable power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see Bera et al., (2013) for further discussion.
-3.2 Different means and same variance
-Next, we consider ANOVA models in which the group means are allowed to differ, that is,
+which implies $\mathbb{E}[\pi_{1}(Z)]=0$ and thus no power under this choice.
+In fact, smooth tests are neither directional nor omnibus; that is, they maintain reasonable power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see bera2013smooth for further discussion.
+3.2 Heterogeneous group means with common variances
+Next, we consider ANOVA models with heterogeneous group means while maintaining common variances:
 $Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.9)
 where $\{\mu_{j}\}_{j=1}^{J}$ denote the potentially distinct means of each group.
-Let $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ be the estimators of $\mu_{j}$ ($j=1,\ldots,J$) and $\sigma^{2}$, respectively.
-Correspondingly we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
-The following theorem characterizes the properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ under the model (3.9).
+Let $\widehat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\widehat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\widehat{\mu}_{j}\right)^{2}$ be the estimators of $\mu_{j}$ ($j=1,\ldots,J$) and $\sigma^{2}$, respectively.
+Correspondingly we define $\widehat{\varepsilon}_{ij}=Y_{ij}-\widehat{\mu}_{j}$, $\widehat{e}_{ij}=\widehat{\varepsilon}_{ij}/\widehat{\sigma}=(Y_{ij}-\widehat{\mu}_{j})/\widehat{\sigma}$, and $\widehat{Z}_{ij}=\Phi(\widehat{e}_{ij})$.
+The following theorem characterizes the properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ under the model (3.9).
 Theorem 4.
 Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
-as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\hat{\Psi}_{K}^{2}$ defined in (3.3), we have
-$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}.$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
+as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\widehat{\Psi}_{K}^{2}$ defined in (3.3), we have
+$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}.$
 Theorem 4 parallels Theorem 1 and Corollary 1, except that it requires additional conditions on the group sample sizes and the number of groups.
@@
 First, the sample size of each group, $N_{j}$, must tend to infinity.
-In model (3.9), where $\mu_{j}$ can be distinct, only the observations within group $j$ can be used to estimate $\mu_{j}$, making $N_{j}\to\infty$ necessary for the consistency of $\hat{\mu}_{j}$.
+In model (3.9), where $\mu_{j}$ can be distinct, only the observations within group $j$ can be used to estimate $\mu_{j}$, making $N_{j}\to\infty$ necessary for the consistency of $\widehat{\mu}_{j}$.
 Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must hold jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$.
@@
 In contrast, for model (3.1) in the previous section, it suffices that the total sample size $N\to\infty$ to derive asymptotic properties, regardless of the magnitudes of $N_{j}$ and $J$.
-This is because in model (3.1), all groups share the same mean $\mu$ and variance $\sigma^{2}$, so the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d., allowing them to be pooled for efficient estimation and inference.
-The following two theorems establish formally the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ (defined in (3.3)) under the alternatives.
+This is because in model (3.1), all groups share the common mean $\mu$ and common variance $\sigma^{2}$, so the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d., allowing them to be pooled for efficient estimation and inference.
+The following two theorems establish formally the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ and $\widehat{\Psi}_{K}^{2}$ (defined in (3.3)) under the alternatives.
 The conditions are identical to those in Theorem 4, and the results parallel those in Theorems 2 and 3.
@@
 Suppose Assumptions 3.1 and 3.2 hold. Then, under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
-$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
+$+\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore,
-$\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K},$
+$\widehat{\Psi}_{K}^{2}\overset{p}{\to}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K},$
 and
-$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right).$
+$\sqrt{N}\left(\widehat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right).$
 Theorem 6.
 Suppose Assumptions 3.1 and 3.2 hold. Then, under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
 $+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore,
-$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right).$
+$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right).$
 Theorems 4, 5, and 6 provide an asymptotic $\chi^{2}$ test for the normality of model (3.9) and demonstrate its power properties.
@@
 When the group means in (3.9) are all equal in the sense that $\mu_{1}=\ldots=\mu_{J}=\mu$, the methodology in this section can be reduced to that in Section 3.1 for model (3.1), which validates the unified inferential framework and reflects the effects of group heterogeneity in means.
-3.3 Same mean and different variances
-We now extend model (3.1) by allowing the group variances to differ across groups, that is,
+3.3 Common group means with heterogeneous variances
+We now extend model (3.1) to allow for heterogeneous group variances, that is,
 $Y_{ij}=\mu+\varepsilon_{ij}=\mu+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.10)
 where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially distinct standard deviations of each group.
-For estimation, we adopt $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\hat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2}$ ($j=1,\ldots,J$), respectively.
+For estimation, we adopt $\widehat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\widehat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\widehat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2}$ ($j=1,\ldots,J$), respectively.
 Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of group variances.
-Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu})/\hat{\sigma}_{j}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
+Define $\widehat{\varepsilon}_{ij}=Y_{ij}-\widehat{\mu}$, $\widehat{e}_{ij}=\widehat{\varepsilon}_{ij}/\widehat{\sigma}_{j}=(Y_{ij}-\widehat{\mu})/\widehat{\sigma}_{j}$ and $\widehat{Z}_{ij}=\Phi(\widehat{e}_{ij})$.
 Let $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ be quantities characterizing the relative proportions of the group sample sizes.
@@
 These conditions imply that the group-specific sample sizes and standard deviations are of the same order of magnitude.
-Following the approaches in Sections 3.1 and 3.2, we first examine the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, which is summarized in the theorem below.
+Following the approach in Sections 3.1 and 3.2, we first examine the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$, which is summarized in the theorem below.
 Theorem 7.
 Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.11)
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.11)
 $+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$.
-Note that the asymptotic decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4.
-The difference lies in the parameter estimation effect of $\hat{\mu}$, which is reflected by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right-hand side of (3.11).
-Unlike model (3.1), which uses the sample mean to estimate $\mu$, model (3.10) employs a different estimator, $\hat{\mu}$, resulting in this distinct estimation effect.
-In particular, when all groups share the same variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term reduces to $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect simplifies to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$.
+Note that the asymptotic decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4.
+The difference lies in the parameter estimation effect of $\widehat{\mu}$, which is reflected by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right-hand side of (3.11).
+Unlike model (3.1), which uses the sample mean to estimate $\mu$, model (3.10) employs a different estimator, $\widehat{\mu}$, resulting in this distinct estimation effect.
+In particular, when all groups share the common variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term reduces to $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect simplifies to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$.
 In contrast, the estimation effect of the variances always aggregates into the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
-The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$ follows directly from (3.11).
+The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\widehat{Z}_{ij})$ follows directly from (3.11).
 Under the conditions of Theorem 7, an application of the CLT yields that
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
-$\to_{d}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{Z}_{j},$
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\widehat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\widehat{Z}_{ij}\right)\right]^{\top}$
+$\overset{d}{\to}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{Z}_{j},$
 where $\bm{Z}_{1},\ldots,\bm{Z}_{J}$ are independently distributed as
@@
 with $\bm{\Omega}_{K}^{(j)}=\left(\omega_{kl}^{(j)}\right)_{K\times K}$ given by
-$\omega_{kl}^{(j)}$ $=\mathrm{E}\left\{\pi_{k}(Z)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right\}$
+$\omega_{kl}^{(j)}$ $=\mathbb{E}\left\{\pi_{k}(Z)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right\}$
 $\qquad\left\{\pi_{l}(Z)-c_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right\}$
@@
 Therefore, under $H_{0}$, we have
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right).$
-By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ with their sample analogues $\hat{\sigma}_{j}$, $\hat{p}_{j}=N_{j}/N$ and $\hat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted by $\widehat{\bm{\Omega}}_{K}^{(j)}$.
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\overset{d}{\to}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right).$
+By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ with their sample analogues $\widehat{\sigma}_{j}$, $\widehat{p}_{j}=N_{j}/N$ and $\widehat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted by $\widehat{\bm{\Omega}}_{K}^{(j)}$.
 Accordingly, the test statistic is defined as
-$\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}\hat{p}_{j}\widehat{\bm{\Omega}}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.12)
+$\widehat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}\widehat{p}_{j}\widehat{\bm{\Omega}}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right).$ (3.12)
 We also introduce the infeasible version,
-$\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.13)
+$\widetilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right).$ (3.13)
 Corollary 2.
@@
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$.
-Corollary 2 shows that under the null hypothesis, both the feasible test statistic $\hat{\Psi}_{K}^{2}$ (defined in (3.12)) and its infeasible counterpart $\tilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed when multiplied by the total sample size $N$.
+Corollary 2 shows that under the null hypothesis, both the feasible test statistic $\widehat{\Psi}_{K}^{2}$ (defined in (3.12)) and its infeasible counterpart $\widetilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed when multiplied by the total sample size $N$.
 Although these test statistics take a slightly different form from those in the previous Sections 3.1 and 3.2, the limiting Chi-square distribution is retained, and the convergence rate remains $N$, which further supports the validity of our unified theoretical framework.
-The following two theorems characterize the asymptotic behavior of $\tilde{\Psi}_{K}^{2}$ under both fixed and local alternatives, in a manner analogous to that of $\hat{\Psi}_{K}^{2}$ discussed in Sections 3.1 and 3.2.
+The following two theorems characterize the asymptotic behavior of $\widetilde{\Psi}_{K}^{2}$ under both fixed and local alternatives, in a manner analogous to that of $\widehat{\Psi}_{K}^{2}$ discussed in Sections 3.1 and 3.2.
 Theorem 8.
 Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
-$-\left.\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
+$-\left.\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$. Furthermore,
-$\tilde{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K},$
+$\widetilde{\Psi}_{K}^{2}\overset{p}{\to}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K},$
 and
-$\sqrt{N}\left(\tilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Upsilon}_{K}\bm{a}_{K}\right),$ (3.15)
+$\sqrt{N}\left(\widetilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Upsilon}_{K}\bm{a}_{K}\right),$ (3.15)
 where
@@
 and $\bm{\Lambda}_{K}^{(j)}=\left(\lambda_{kl}^{(j)}\right)_{K\times K}$ is given by
-$\lambda_{kl}^{(j)}=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
-$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathrm{E}\left[e^{3}\right]$
-$+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
+$\lambda_{kl}^{(j)}=$ $\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
+$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathbb{E}\left[e^{3}\right]$
+$+\frac{d_{2k}d_{2l}}{4}\left[\mathbb{E}\left[e^{4}\right]-1\right].$
 Theorem 9.
 Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
 $-\left.\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$. Furthermore,
-$N\tilde{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{\Delta}_{K}\right).$
+$N\widetilde{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{\Delta}_{K}\right).$
 Remark 2.
-To conclude this section, we summarize the regularity conditions for the three models in Table 1.
-On one hand, more complex model structures naturally impose additional restrictions on parameters and group sample sizes, as the heterogeneity across groups introduces extra challenges for theoretical analysis.
+To conclude this section, we summarize the regularity conditions on sample size structure and cross-group heterogeneity for the three models (3.1), (3.9), and (3.10) in Table 1.
+On the one hand, more complex model structures naturally impose additional restrictions on within-group variability and group sample sizes, as heterogeneity across groups introduces additional challenges for theoretical analysis.
 On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can be established for each model.
-This highlights that our proposed methodology provides a unified theoretical framework while remaining flexible enough to accommodate the distinctive features of different ANOVA settings.
-Table 1: The regularity conditions for the three models
+This highlights that our proposed methodology provides a unified theoretical framework while remaining flexible enough to accommodate the distinctive features of various ANOVA settings.
+Table 1: Regularity conditions on sample sizes and group heterogeneity for the three one-way models
 Model Conditions
-Same mean & same variance (3.1) $N\to\infty$
-Different means & same variance (3.9)
-$\min\{N_{1},\ldots,N_{J}\}\to\infty$,
-$J=o(N^{1/2})$, $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$
-Same mean & different variances (3.10)
-$\min\{N_{1},\ldots,N_{J}\}\to\infty$,
-$J=o(N^{1/2})$ , $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$,
-$\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$,
-$\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$
+Model (3.1) $N\to\infty$
+Model (3.9) $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$, $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$
+Model (3.10) $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ , $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$, $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$, $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$
 4 Data-driven choice of $K$
 The methodology in Section 3 relies on the fact that the order of smooth alternatives (2.1), $K$, is fixed before testing.
-The choice of $K$ critically affects the performance of smooth tests (Ledwina,, 1994; Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016).
-A large $K$ introduces redundant components in the alternatives that contribute little to the test statistic but inflate the degrees of freedom, leading to power dilution.
-Conversely, a small $K$ increases the risk that $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, rendering the test powerless.
+The choice of $K$ critically affects the performance of smooth tests (ledwina1994data; kallenberg1995consistency; kallenberg1995data; inglot1996asymptotic; inglot1997data; kallenberg1997dataA; kallenberg1997dataB; kallenberg1999data; janic2000data; ducharme2004smooth; ducharme2004goodness; kraus2007data; duchesne2016estimating).
+A large $K$ introduces redundant components into the alternatives that contribute little to the test statistic but inflate the degrees of freedom, thereby diluting power.
+Conversely, a small $K$ increases the risk that $\mathbb{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, rendering the test powerless.
 Hence, an appropriate choice of $K$ is crucial.
-In this section, we focus on data-driven selection of $K$ for the three cases in Section 3.
-In the literature of Neyman’s smooth tests, data-driven selection was first proposed by Ledwina, (1994) for testing uniformity.
-The data-driven procedure consists of two steps.
-First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to determine a suitable dimension for the smooth model that best fits the data.
+In this section, we focus on the data-driven selection of $K$ for the three tests developed in Section 3.
+In Neyman’s smooth test literature, the data-driven selection of $K$ was first proposed by ledwina1994data for testing uniformity.
+The data-driven testing procedure consists of two steps.
+First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to determine the dimension of the smooth model that best fits the data.
 Second, Neyman’s smooth test is performed within the selected model space, yielding a data-driven test statistic.
-The data-driven smooth test thus serves to combine preliminary model selection with a more precise inferential procedure, being Neyman’s test in the “right” direction.
-Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were established in Kallenberg and Ledwina, 1995a and Kallenberg and Ledwina, 1995b , Inglot and Ledwina, (1996).
-Later, modifications of Schwarz’s rule addressed more complex problems, including testing for composite hypotheses (Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Inglot et al.,, 1997), testing for independence (Kallenberg and Ledwina,, 1999), two-sample testing problems (Janic-Wróblewska and Ledwina,, 2000), and so on.
+The data-driven smooth test thus combines preliminary model selection with a more precise inferential procedure, serving as Neyman’s test in the “right” direction.
+Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were established in kallenberg1995consistency and kallenberg1995data, inglot1996asymptotic.
+Later, modifications of Schwarz’s rule addressed more complex problems, including testing composite hypotheses (kallenberg1997dataA; kallenberg1997dataB; inglot1997data), testing independence (kallenberg1999data), and two-sample testing (janic2000data), among others.
 These modified rules leverage smooth test statistics directly, avoiding the computation of the maximized log-likelihood, and are easier to implement.
 Motivated by these strategies, we adopt a modified Schwarz’s rule to determine the order $K$ for smooth tests in ANOVA models. The selection rule is specified as follows:
-$\hat{K}=\min\left\{\mathop{\arg\max}\limits_{1\leq k\leq D}\left(N\hat{\Psi}_{k}^{2}-k\log N\right)\right\},$ (4.1)
-where $D$ is a fixed positive integer, and $\hat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) depending on the specific scenario.
-The resulting data-driven test statistic is $N\Psi_{\hat{K}}^{2}$, with $\hat{K}$ given by (4.1).
-The form of (4.1) is mainly inspired by Ducharme and Fontez, (2004), Ducharme and Lafaye de Micheaux, (2004), Kraus, (2007) and Duchesne et al., (2016) where the upper bound of selection $D$ is fixed.
-An alternative is to let $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000).
-Although a diverging upper bound can theoretically improve the consistency of smooth tests, the required rate of divergence is slow, and simulation studies show that empirical power levels off rapidly as $D$ increases.
+$\widehat{K}=\min\left\{\mathop{\arg\max}\limits_{1\leq k\leq D}\left(N\widehat{\Psi}_{k}^{2}-k\log N\right)\right\},$ (4.1)
+where $D$ is a fixed positive integer, and $\widehat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) depending on the specific scenario.
+The resulting data-driven test statistic is $N\widehat{\Psi}_{\widehat{K}}^{2}$, with $\widehat{K}$ given by (4.1).
+The form of (4.1) is mainly inspired by ducharme2004smooth, ducharme2004goodness, kraus2007data and duchesne2016estimating where the upper bound of selection $D$ is fixed.
+An alternative is to let $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, kallenberg1995consistency; kallenberg1995data; inglot1996asymptotic; inglot1997data; kallenberg1997dataA; kallenberg1997dataB; kallenberg1999data; janic2000data).
+Although a diverging upper bound can, in principle, improve the consistency of smooth tests, the required rate of divergence is slow, and simulation studies show that empirical power levels off rapidly as $D$ increases.
 Hence, our proposed selection rule (4.1) with fixed $D$ is both reasonable and easy to implement in practice.
-To establish the theoretical properties of $\hat{K}$ and $N\hat{\Psi}_{\hat{K}}^{2}$, we introduce a revised version of the alternative hypothesis $H_{1}$ (3.4) and Assumption 3.2. Consider
-$H_{1}^{\prime}:\mathrm{E}[\pi_{1}(Z)]=\ldots=\mathrm{E}[\pi_{K_{0}-1}(Z)]=0,\mathrm{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D,$ (4.2)
+To establish the theoretical properties of $\widehat{K}$ and $N\widehat{\Psi}_{\widehat{K}}^{2}$, we introduce a revised version of the alternative hypothesis $H_{1}$ (3.4) and Assumption 3.2.
+Consider
+$H_{1}^{\prime}:\mathbb{E}[\pi_{1}(Z)]=\ldots=\mathbb{E}[\pi_{K_{0}-1}(Z)]=0,\mathbb{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D,$ (4.2)
 which encompasses a broader range of alternatives than $H_{1}$ by replacing the fixed order $K$ with the larger $D$.
@@
   • (1).
-    For model (3.1) in Section 3.1, as $N\to\infty$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$ and $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\hat{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
+    For model (3.1) in Section 3.1, as $N\to\infty$, under $H_{0}:e\sim\Phi(x)$, $\mathbb{P}(\widehat{K}=1)\to 1$ and $N\widehat{\Psi}_{\widehat{K}}^{2}\overset{d}{\to}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathbb{P}(\widehat{K}\geq K_{0})\to 1$ and $\mathbb{P}(N\widehat{\Psi}^{2}_{\widehat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
   • (2).
-    For model (3.9) in Section 3.2, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$ and $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\hat{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
+    For model (3.9) in Section 3.2, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$, under $H_{0}:e\sim\Phi(x)$, $\mathbb{P}(\widehat{K}=1)\to 1$ and $N\widehat{\Psi}_{\widehat{K}}^{2}\overset{d}{\to}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathbb{P}(\widehat{K}\geq K_{0})\to 1$ and $\mathbb{P}(N\widehat{\Psi}^{2}_{\widehat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
   • (3).
-    For model (3.10) in Section 3.3, with additional conditions such that (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$, $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$ and $N\tilde{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\tilde{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
+    For model (3.10) in Section 3.3, with additional conditions such that (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, under $H_{0}:e\sim\Phi(x)$, $\mathbb{P}(\widehat{K}=1)\to 1$, $N\widehat{\Psi}_{\widehat{K}}^{2}\overset{d}{\to}\chi_{1}^{2}$ and $N\widetilde{\Psi}_{\widehat{K}}^{2}\overset{d}{\to}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathbb{P}(\widehat{K}\geq K_{0})\to 1$ and $\mathbb{P}(N\widetilde{\Psi}^{2}_{\widehat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
 Theorem 10 presents the unified results of data-driven smooth tests for the three cases in Section 3.
-Under the null hypothesis, the probability of $\{\hat{K}=1\}$ tends to $1$ asymptotically, implying that the first-order smooth model provides the best fit to $\{\hat{Z}_{ij}\}_{i=1,j=1}^{N_{j},J}$ among the $D$ candidate models.
-From the perspective of hypothesis testing, this means that $\hat{\Psi}_{1}^{2}$ is informative and sufficient for assessing uniformity or normality.
-Meanwhile, the data-driven selection procedure allows for the identification of the mechanism behind nonuniformity (and also nonnormality) and enhances the test’s power against a broader class of alternatives, $H_{1}^{\prime}$.
+Under the null hypothesis, the probability of $\{\widehat{K}=1\}$ tends to $1$ asymptotically, implying that the first-order smooth model provides the best fit to $\{\widehat{Z}_{ij}\}_{i=1,j=1}^{N_{j},J}$ among the $D$ candidate models.
+From the perspective of hypothesis testing, this means that $\widehat{\Psi}_{1}^{2}$ is informative and sufficient for assessing uniformity or normality.
+Meanwhile, the data-driven selection procedure allows identification of the mechanism underlying nonuniformity (equivalently, nonnormality) and enhances the test’s power against a broader class of alternatives, $H_{1}^{\prime}$.
 These findings are consistent with previous results on data-driven smooth tests.
-In practice, however, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics often performs poorly in finite samples (Kallenberg and Ledwina, 1995b, ; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016).
+In practice, however, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics often performs poorly in finite samples (kallenberg1995data; inglot1997data; kallenberg1997dataA; kallenberg1997dataB; kallenberg1999data; janic2000data; ducharme2004smooth; ducharme2004goodness; kraus2007data; duchesne2016estimating).
 To address this issue in our problem, and provided the nested orthonormal system (i.e.
-$\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the approach of Kallenberg and Ledwina, 1995b , Inglot et al., (1997), Kallenberg and Ledwina, (1997), Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007) and adopt the following finite-sample approximation for the null distribution of $N\hat{\Psi}_{\hat{K}}^{2}$:
+$\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the approach of kallenberg1995data, inglot1997data, kallenberg1997dataA, kallenberg1999data, janic2000data, kraus2007data and adopt the following finite-sample approximation for the null distribution of $N\widehat{\Psi}_{\widehat{K}}^{2}$:
 […]
-Such an approximation primarily accounts for the selection uncertainty of $\hat{K}$.
-In finite samples, under the null, the empirical probability of the event $\{\hat{K}=1\}$ may not be exactly $1$, as $\{\hat{K}=2\}$ can occur with a small but negligible probability.
-Consequently, (4.3) is derived from the approximation $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x)\approx\mathrm{P}(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1)+\mathrm{P}(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2)$.
-The reliability and accuracy of $H(x)$ are further illustrated through simulation studies in the next section and the appendix.
-5 Simulation
-In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the proposed tests developed in Sections 3 and 4.
-Following Neyman, (1937), Bera et al., (2013) and Song and Xiao, (2022), we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$.
-Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of Duchesne et al., (2016), and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material.
-From a computational perspective, the test statistic $N\hat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in Duchesne et al., (2016), which leverages the closed form of $\bm{\Sigma}_{K}^{-1}$ [see equations (12) and (13) therein].
+Such an approximation primarily accounts for the selection uncertainty of $\widehat{K}$.
+In finite samples, under the null, the empirical probability of the event $\{\widehat{K}=1\}$ may not be exactly $1$, as $\{\widehat{K}=2\}$ can occur with a small but non-negligible probability.
+Consequently, (4.3) is derived from the approximation $\mathbb{P}(N\widehat{\Psi}_{\widehat{K}}^{2}\leq x)\approx\mathbb{P}(N\widehat{\Psi}_{1}^{2}\leq x,\widehat{K}=1)+\mathbb{P}(N\widehat{\Psi}_{2}^{2}\leq x,\widehat{K}=2)$.
+The reliability and accuracy of $H(x)$ are further illustrated through simulation studies in the next section and in the Online Appendix.
+5 Simulations
+In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the smooth tests developed in Sections 3 and 4.
+Following neyman1937smooth, bera2013smooth, song2022smooth, and beutner2025two, we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=1}^{\infty}$.
+Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of duchesne2016estimating, and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material.
+From a computational perspective, the test statistic $N\widehat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in duchesne2016estimating, which leverages the closed form of $\bm{\Sigma}_{K}^{-1}$ (see equations (12) and (13) therein).
+For comparison, we include three classical normalty tests in the literature—Shapiro–Wilk (SW), Jarque–Bera (JB), and Kolmogorov–Smirnov (KS) tests—applied to the standardized residuals $\{\widehat{e}_{ij}\}_{i=1,j=1}^{N_{j},J}$ under the corresponding model, without accounting for the parameter estimation effects.
 We first study the tests for model (3.1) through two experiments. The observations are generated as follows.
@@
 Here $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$.
-The sample size parameter $m$ varies from 10 to 150 in increments of 10.
-We set the significance level $\alpha=5\%$ and perform $500$ replications.
-We evaluate the test statistic $N\hat{\Psi}_{K}^{2}$ from Section 3.1, with $1\leq K\leq 5$, and the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ from Section 4, where $\hat{K}$ is determined by (4.1) with $D=5$.
-Results of Experiment I are shown in Tables 2, 3, 4 and Figure 1.
-Table 2 reports the empirical rejection rates under $H_{0}$.
-For fixed $K$, $N\hat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes such as $m=10$ (corresponding to a total sample size of $N=150$).
-For the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$, the limiting $\chi^{2}_{1}$ distribution underestimates the variability under $H_{0}$, while the revised distribution $H(x)$ achieves better size control.
-Tables 3 and 4 present the empirical frequency of $\hat{K}$ under $H_{0}$ and $H_{1}$, respectively.
-Under $H_{0}$, the frequency of $\{\hat{K}=1\}$ approaches $1$ as $m$ increases, whereas under $H_{1}$, $\hat{K}$ is consistently greater than $1$, in line with theoretical expectations.
-Figure 1 displays the sample means of $\hat{K}$ across different values of $m$, with error bars representing one standard deviation around the mean.
-Under $H_{0}$, the sample mean of $\hat{K}$ stays close to $1$ with diminishing variability as $m$ grows, indicating consistent model selection.
-Under $H_{1}$, $\hat{K}$ increases steadily and approaches the maximum $D=5$, with reduced variability, demonstrating the adaptiveness of the data-driven procedure.
-The results of Experiment II are quite similar; see Tables 5, 6, and 7, as well as Figure 2.
-However, from Table 5, under $H_{1}$, the test statistic $N\hat{\Psi}_{1}^{2}$ fails to exhibit power because in this case, $\mathrm{E}[\pi_{1}(Z)]=\mathrm{E}[\pi_{1}(\Phi(Y_{ij}^{(1,2)}-8))]=0$.
-This illustrates that $N\hat{\Psi}_{K}^{2}$ may be inconsistent against certain alternatives, especially when $K$ is small.
-Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, $\{\hat{K}=1\}$ does not occur, consistent with Theorem 10 (3), which ensures that $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ under $H_{1}^{\prime}$ with $K_{0}=2$.
-Table 2: Empirical rejection rates under $H_{0}$ in Experiment I, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
-The empirical power under $H_{1}$ equals 1 across all settings.
-$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
-10 0.050 0.048 0.048 0.044 0.040 0.056 0.034
-20 0.046 0.044 0.034 0.040 0.048 0.066 0.046
-30 0.042 0.036 0.044 0.068 0.068 0.060 0.046
-40 0.058 0.040 0.036 0.058 0.052 0.058 0.050
-50 0.064 0.044 0.054 0.054 0.062 0.068 0.060
-60 0.054 0.034 0.038 0.044 0.050 0.070 0.062
-70 0.056 0.038 0.048 0.046 0.052 0.056 0.048
-80 0.038 0.050 0.072 0.042 0.048 0.058 0.048
-90 0.062 0.066 0.052 0.064 0.058 0.048 0.040
-100 0.034 0.050 0.050 0.060 0.054 0.052 0.038
-110 0.054 0.032 0.054 0.060 0.050 0.038 0.036
-120 0.052 0.052 0.038 0.044 0.052 0.052 0.050
-130 0.052 0.050 0.064 0.060 0.058 0.068 0.064
-140 0.062 0.040 0.026 0.046 0.030 0.038 0.030
-150 0.054 0.044 0.062 0.034 0.058 0.056 0.048
-Table 3: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment I
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-10 0.982 0.018 0 0 0
-20 0.988 0.010 0 0 0.002
-30 0.986 0.012 0 0.002 0
-40 0.984 0.014 0.002 0 0
-50 0.988 0.012 0 0 0
-60 0.980 0.020 0 0 0
-70 0.994 0.006 0 0 0
-80 0.996 0.002 0.002 0 0
-90 0.996 0.002 0.002 0 0
-100 0.996 0.004 0 0 0
-110 0.994 0.006 0 0 0
-120 0.990 0.010 0 0 0
-130 0.994 0.006 0 0 0
-140 0.994 0.006 0 0 0
-150 0.994 0.006 0 0 0
-Table 4: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment I
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-10 0.004 0 0 0.484 0.512
-20 0 0 0 0.326 0.674
-30 0 0 0 0.194 0.806
-40 0 0 0 0.090 0.910
-50 0 0 0 0.076 0.924
-60 0 0 0 0.030 0.970
-70 0 0 0 0.012 0.988
-80 0 0 0 0.010 0.990
-90 0 0 0 0.006 0.994
-100 0 0 0 0.004 0.996
-110 0 0 0 0.006 0.994
-120 0 0 0 0 1
-130 0 0 0 0 1
-140 0 0 0 0.002 0.998
-150 0 0 0 0 1
-Figure 1: The sample means and error bars of $\hat{K}$ in Experiment I
-Table 5: Empirical rejection rates in Experiment II, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
-$H_{0}$ $H_{1}$
-$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
-10 0.042 0.050 0.046 0.042 0.052 0.076 0.048 0.046 0.998 0.996 0.998 0.996 0.998 0.998
-20 0.034 0.042 0.026 0.050 0.060 0.070 0.054 0.018 1 1 1 1 1 1
-30 0.028 0.048 0.036 0.046 0.046 0.070 0.062 0.022 1 1 1 1 1 1
-40 0.054 0.054 0.046 0.050 0.060 0.062 0.052 0.042 1 1 1 1 1 1
-50 0.064 0.062 0.042 0.060 0.050 0.068 0.058 0.032 1 1 1 1 1 1
-60 0.032 0.054 0.060 0.040 0.040 0.078 0.064 0.026 1 1 1 1 1 1
-70 0.076 0.038 0.052 0.052 0.036 0.052 0.042 0.022 1 1 1 1 1 1
-80 0.040 0.046 0.034 0.048 0.058 0.074 0.068 0.028 1 1 1 1 1 1
-90 0.044 0.052 0.050 0.054 0.042 0.060 0.050 0.032 1 1 1 1 1 1
-100 0.038 0.068 0.050 0.046 0.046 0.064 0.056 0.028 1 1 1 1 1 1
-110 0.052 0.052 0.046 0.052 0.048 0.050 0.046 0.032 1 1 1 1 1 1
-120 0.052 0.042 0.038 0.034 0.052 0.040 0.036 0.030 1 1 1 1 1 1
-130 0.048 0.052 0.044 0.036 0.048 0.062 0.054 0.040 1 1 1 1 1 1
-140 0.062 0.058 0.056 0.038 0.038 0.046 0.042 0.044 1 1 1 1 1 1
-150 0.038 0.056 0.046 0.052 0.060 0.060 0.054 0.034 1 1 1 1 1 1
-Table 6: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment II
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-10 0.968 0.026 0.004 0 0.002
-20 0.982 0.016 0.002 0 0
-30 0.978 0.020 0.002 0 0
-40 0.992 0.008 0 0 0
-50 0.990 0.006 0.004 0 0
-60 0.996 0.004 0 0 0
-70 0.994 0.006 0 0 0
-80 0.984 0.016 0 0 0
-90 0.996 0.004 0 0 0
-100 0.994 0.006 0 0 0
-110 0.996 0.004 0 0 0
-120 0.996 0.004 0 0 0
-130 0.994 0.006 0 0 0
-140 0.994 0.006 0 0 0
-150 0.988 0.012 0 0 0
-Table 7: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment II
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-10 0.002 0.984 0.012 0 0.002
-20 0 0.994 0.004 0 0.002
-30 0 0.990 0.008 0 0.002
-40 0 0.996 0.004 0 0
-50 0 0.996 0.004 0 0
-60 0 0.994 0.004 0 0.002
-70 0 0.998 0.002 0 0
-80 0 1 0 0 0
-90 0 1 0 0 0
-100 0 0.996 0.004 0 0
-110 0 0.998 0.002 0 0
-120 0 0.994 0.006 0 0
-130 0 0.998 0.002 0 0
-140 0 0.998 0.002 0 0
-150 0 0.996 0.004 0 0
-Figure 2: The sample means and error bars of $\hat{K}$ in Experiment II
-The next two experiments are designed to evaluate the tests for models (3.9) and (3.10), respectively.
-All settings are inherited from the previous two experiments, except that the data-generating process is modified accordingly.
+For the proposed smooth tests, we evaluate the test statistic $N\widehat{\Psi}_{K}^{2}$ from Section 3.1 with $1\leq K\leq 5$, as well as the data-driven statistic $N\widehat{\Psi}_{\widehat{K}}^{2}$ from Section 4, where $\widehat{K}$ is determined by (4.1) with $D=5$.
+The sample size parameter $m$ ranges from 10 to 150 in increments of 10, and the significance level is set at $\alpha=5\%$ with $500$ replications performed.
+Results of Experiment I are presented in Tables 2, 3, and Figure 1.
+Table 2 reports the empirical rejection rates under $H_{0}$ for the proposed smooth tests and the classical normality tests.
+The smooth tests comprise both fixed-$K$ versions ($K=1,\ldots,5$) and data-driven procedures, implemented using the limiting $\chi^{2}_{1}$ null distribution ($\widehat{K}\&\chi^{2}_{1}$) or the approximated null distribution $H(x)$ ($\widehat{K}\&H(x)$).
+For fixed $K$, under $H_{0}$, the statistic $N\widehat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes such as $m=10$ (corresponding to a total sample size of $N=150$).
+For the data-driven test statistic $N\widehat{\Psi}_{\widehat{K}}^{2}$, the limiting $\chi^{2}_{1}$ distribution fails to control the rejection rate under $H_{0}$, whereas the approximated distribution $H(x)$ achieves better size control, yielding rejection rates close to the nominal $5\%$ level.
+Under $H_{1}$, all versions of the proposed smooth tests exhibit high power across the considered settings.
+As for the classical methods, although the Shapiro–Wilk and Jarque–Bera tests are applied directly to the standardized residuals without accounting for the estimation effects, their empirical rejection rates under $H_{0}$ remain close to the nominal $5\%$ level.
+Under $H_{1}$, both tests achieve high power, indicating that they still provide reasonably reliable numerical performance despite the lack of formal adjustment for the estimation effect.
+In contrast, under $H_{0}$, the Kolmogorov–Smirnov test exhibits almost zero rejection across all settings, indicating the practical failure of the test.
+Table 3 presents the empirical frequency of $\widehat{K}$ under $H_{0}$ and $H_{1}$, respectively.
+Under $H_{0}$, the frequency of $\{\widehat{K}=1\}$ approaches $1$ as $m$ increases, whereas under $H_{1}$, $\widehat{K}$ is consistently greater than $1$, in line with theoretical expectations.
+Figure 1 displays the sample means of $\widehat{K}$ across different values of $m$, with error bars representing one sample standard deviation around the sample means.
+Under $H_{0}$, the sample mean of $\widehat{K}$ stays close to $1$ with diminishing variability as $m$ grows, indicating consistent model selection.
+Under $H_{1}$, $\widehat{K}$ increases steadily and approaches the maximum $D=5$, with reduced variability, demonstrating the adaptiveness of the data-driven procedure.
+The results of Experiment II are quite similar; see Tables 4, 5, 6, and Figure 2.
+However, from Table 6, under $H_{1}$, the test statistic $N\widehat{\Psi}_{1}^{2}$ fails to exhibit power as a result of $\mathbb{E}[\pi_{1}(Z)]=\mathbb{E}[\pi_{1}(\Phi(Y_{ij}^{(1,2)}-8))]=0$ in this case.
+This illustrates that $N\widehat{\Psi}_{K}^{2}$ may be inconsistent against certain alternatives, especially when $K$ is small.
+Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, $\{\widehat{K}=1\}$ does not occur, consistent with Theorem 10 (1), which ensures that $\mathbb{P}(\widehat{K}\geq K_{0})\to 1$ under $H_{1}^{\prime}$ with $K_{0}=2$.
+Table 2: Empirical rejection rates under $H_{0}$ in Experiment I. The empirical power under $H_{1}$ equals 1 for all sample sizes and all considered tests.
+Smooth tests SW JB KS
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
+10 0.050 0.048 0.048 0.044 0.040 0.056 0.034 0.054 0.044 0
+20 0.046 0.044 0.034 0.040 0.048 0.066 0.046 0.054 0.042 0
+30 0.042 0.036 0.044 0.068 0.068 0.060 0.046 0.040 0.032 0.002
+40 0.058 0.040 0.036 0.058 0.052 0.058 0.050 0.044 0.038 0
+50 0.064 0.044 0.054 0.054 0.062 0.068 0.060 0.036 0.030 0
+60 0.054 0.034 0.038 0.044 0.050 0.070 0.062 0.056 0.058 0
+70 0.056 0.038 0.048 0.046 0.052 0.056 0.048 0.072 0.062 0
+80 0.038 0.050 0.072 0.042 0.048 0.058 0.048 0.046 0.038 0
+90 0.062 0.066 0.052 0.064 0.058 0.048 0.040 0.028 0.038 0
+100 0.034 0.050 0.050 0.060 0.054 0.052 0.038 0.038 0.040 0
+110 0.054 0.032 0.054 0.060 0.050 0.038 0.036 0.042 0.044 0
+120 0.052 0.052 0.038 0.044 0.052 0.052 0.050 0.050 0.046 0
+130 0.052 0.050 0.064 0.060 0.058 0.068 0.064 0.050 0.060 0
+140 0.062 0.040 0.026 0.046 0.030 0.038 0.030 0.068 0.050 0
+150 0.054 0.044 0.062 0.034 0.058 0.056 0.048 0.064 0.066 0
+Table 3: Empirical frequency of $\widehat{K}$ in Experiment I
+Under $H_{0}$ Under $H_{1}$
+$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
+10 0.982 0.018 0 0 0 0.004 0 0 0.484 0.512
+20 0.988 0.010 0 0 0.002 0 0 0 0.326 0.674
+30 0.986 0.012 0 0.002 0 0 0 0 0.194 0.806
+40 0.984 0.014 0.002 0 0 0 0 0 0.090 0.910
+50 0.988 0.012 0 0 0 0 0 0 0.076 0.924
+60 0.980 0.020 0 0 0 0 0 0 0.030 0.970
+70 0.994 0.006 0 0 0 0 0 0 0.012 0.988
+80 0.996 0.002 0.002 0 0 0 0 0 0.010 0.990
+90 0.996 0.002 0.002 0 0 0 0 0 0.006 0.994
+100 0.996 0.004 0 0 0 0 0 0 0.004 0.996
+110 0.994 0.006 0 0 0 0 0 0 0.006 0.994
+120 0.990 0.010 0 0 0 0 0 0 0 1
+130 0.994 0.006 0 0 0 0 0 0 0 1
+140 0.994 0.006 0 0 0 0 0 0 0.002 0.998
+150 0.994 0.006 0 0 0 0 0 0 0 1
+Figure 1: The sample means and error bars of $\widehat{K}$ in Experiment I
+Table 4: Empirical rejection rates under $H_{0}$ in Experiment II
+Smooth tests SW JB KS
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
+10 0.042 0.050 0.046 0.042 0.052 0.076 0.048 0.050 0.042 0
+20 0.034 0.042 0.026 0.050 0.060 0.070 0.054 0.032 0.042 0
+30 0.028 0.048 0.036 0.046 0.046 0.070 0.062 0.064 0.052 0.002
+40 0.054 0.054 0.046 0.050 0.060 0.062 0.052 0.044 0.048 0
+50 0.064 0.062 0.042 0.060 0.050 0.068 0.058 0.042 0.044 0
+60 0.032 0.054 0.060 0.040 0.040 0.078 0.064 0.056 0.062 0
+70 0.076 0.038 0.052 0.052 0.036 0.052 0.042 0.054 0.054 0
+80 0.040 0.046 0.034 0.048 0.058 0.074 0.068 0.044 0.044 0
+90 0.044 0.052 0.050 0.054 0.042 0.060 0.050 0.052 0.052 0
+100 0.038 0.068 0.050 0.046 0.046 0.064 0.056 0.044 0.038 0
+110 0.052 0.052 0.046 0.052 0.048 0.050 0.046 0.046 0.060 0
+120 0.052 0.042 0.038 0.034 0.052 0.040 0.036 0.044 0.052 0
+130 0.048 0.052 0.044 0.036 0.048 0.062 0.054 0.054 0.038 0
+140 0.062 0.058 0.056 0.038 0.038 0.046 0.042 0.032 0.044 0
+150 0.038 0.056 0.046 0.052 0.060 0.060 0.054 0.050 0.050 0
+Table 5: Empirical rejection rates under $H_{1}$ in Experiment II
+Smooth tests SW JB KS
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
+10 0.046 0.998 0.996 0.998 0.996 0.998 0.998 1 0.990 0.824
+20 0.018 1 1 1 1 1 1 1 1 0.994
+30 0.022 1 1 1 1 1 1 1 1 1
+40 0.042 1 1 1 1 1 1 1 1 1
+50 0.032 1 1 1 1 1 1 1 1 1
+60 0.026 1 1 1 1 1 1 1 1 1
+70 0.022 1 1 1 1 1 1 1 1 1
+80 0.028 1 1 1 1 1 1 1 1 1
+90 0.032 1 1 1 1 1 1 1 1 1
+100 0.028 1 1 1 1 1 1 1 1 1
+110 0.032 1 1 1 1 1 1 1 1 1
+120 0.030 1 1 1 1 1 1 1 1 1
+130 0.040 1 1 1 1 1 1 1 1 1
+140 0.044 1 1 1 1 1 1 1 1 1
+150 0.034 1 1 1 1 1 1 1 1 1
+Table 6: Empirical frequency of $\widehat{K}$ in Experiment II
+Under $H_{0}$ Under $H_{1}$
+$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
+10 0.968 0.026 0.004 0 0.002 0.002 0.984 0.012 0 0.002
+20 0.982 0.016 0.002 0 0 0 0.994 0.004 0 0.002
+30 0.978 0.020 0.002 0 0 0 0.990 0.008 0 0.002
+40 0.992 0.008 0 0 0 0 0.996 0.004 0 0
+50 0.990 0.006 0.004 0 0 0 0.996 0.004 0 0
+60 0.996 0.004 0 0 0 0 0.994 0.004 0 0.002
+70 0.994 0.006 0 0 0 0 0.998 0.002 0 0
+80 0.984 0.016 0 0 0 0 1 0 0 0
+90 0.996 0.004 0 0 0 0 1 0 0 0
+100 0.994 0.006 0 0 0 0 0.996 0.004 0 0
+110 0.996 0.004 0 0 0 0 0.998 0.002 0 0
+120 0.996 0.004 0 0 0 0 0.994 0.006 0 0
+130 0.994 0.006 0 0 0 0 0.998 0.002 0 0
+140 0.994 0.006 0 0 0 0 0.998 0.002 0 0
+150 0.988 0.012 0 0 0 0 0.996 0.004 0 0
+Figure 2: The sample means and error bars of $\widehat{K}$ in Experiment II
+The next two experiments, focusing on models (3.9) and (3.10), follow simulation settings similar to the previous ones. The data-generating process for each model is specified as follows.
   • Experiment III:
@@
 Here $\{Y_{ij}^{(0,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$.
-Note that in Experiment III, $\mathrm{E}[Y_{ij}^{(0,3)}]=\mathrm{E}[Y_{ij}^{(1,3)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,3)}]=\operatorname{Var}[Y_{ij}^{(1,3)}]=4$, which corresponds to model (3.9); in Experiment IV, $\mathrm{E}[Y_{ij}^{(0,4)}]=\mathrm{E}[Y_{ij}^{(1,4)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,4)}]=\operatorname{Var}[Y_{ij}^{(1,4)}]=j^{2}$, which corresponds to model (3.10).
-The conclusions from Experiments III and IV are analogous to those from Experiments I and II, respectively (see Tables 813 and Figures 34), thus validating our unified testing framework.
-To examine the robustness of our findings with respect to the number of groups, we additionally consider a variant of Experiment III with $J=10$ (referred to as Experiment III).
-The results corroborate the main conclusions, as reported in A.
-Table 8: Empirical rejection rates under $H_{0}$ in Experiment III, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
-The empirical power under $H_{1}$ equals 1 across all settings.
-$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
-$10$ 0.050 0.046 0.048 0.048 0.052 0.076 0.050
-2020 0.072 0.060 0.050 0.046 0.050 0.054 0.048
-$30$ 0.054 0.046 0.054 0.038 0.056 0.074 0.062
-$40$ 0.046 0.048 0.060 0.054 0.048 0.054 0.050
-$50$ 0.044 0.060 0.052 0.052 0.048 0.040 0.038
-$60$ 0.066 0.044 0.048 0.052 0.058 0.048 0.042
-$70$ 0.044 0.048 0.038 0.046 0.048 0.072 0.064
-$80$ 0.050 0.044 0.048 0.038 0.042 0.074 0.068
-$90$ 0.040 0.060 0.046 0.054 0.046 0.048 0.042
-$100$ 0.062 0.056 0.042 0.050 0.048 0.058 0.052
-$110$ 0.060 0.046 0.066 0.052 0.050 0.068 0.054
-$120$ 0.034 0.044 0.038 0.038 0.036 0.058 0.056
-$130$ 0.038 0.050 0.042 0.052 0.050 0.058 0.048
-$140$ 0.038 0.050 0.048 0.042 0.044 0.044 0.040
-$150$ 0.064 0.056 0.048 0.058 0.058 0.046 0.042
-Table 9: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment III
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-$10$ 0.976 0.024 0 0 0
-2020 0.974 0.018 0.006 0 0.002
-$30$ 0.986 0.014 0 0 0
-$40$ 0.990 0.010 0 0 0
-$50$ 0.998 0.002 0 0 0
-$60$ 0.988 0.012 0 0 0
-$70$ 0.992 0.008 0 0 0
-$80$ 0.990 0.008 0.002 0 0
-$90$ 0.988 0.012 0 0 0
-$100$ 0.996 0.004 0 0 0
-$110$ 0.994 0.004 0.002 0 0
-$120$ 0.992 0.006 0.002 0 0
-$130$ 0.998 0.002 0 0 0
-$140$ 0.992 0.008 0 0 0
-$150$ 0.996 0.004 0 0 0
-Table 10: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment III
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-$10$ 0.032 0.004 0 0.418 0.546
-2020 0 0 0 0.314 0.686
-$30$ 0 0 0 0.174 0.826
-$40$ 0 0 0 0.098 0.902
-$50$ 0 0 0 0.064 0.936
-$60$ 0 0 0 0.036 0.964
-$70$ 0 0 0 0.022 0.978
-$80$ 0 0 0 0.018 0.982
-$90$ 0 0 0 0.008 0.992
-$100$ 0 0 0 0.008 0.992
-$110$ 0 0 0 0.002 0.998
-$120$ 0 0 0 0 1
-$130$ 0 0 0 0 1
-$140$ 0 0 0 0.002 0.998
-$150$ 0 0 0 0 1
-Figure 3: The sample means and error bars of $\hat{K}$ in Experiment III
-Figure 4: The sample means and error bars of $\hat{K}$ in Experiment IV
-In summary, the experiments demonstrate that the proposed tests in Section 3 maintain the nominal significance level well under the null, and exhibit high power under the alternatives, except when $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$.
+Note that in Experiment III, $\mathbb{E}[Y_{ij}^{(0,3)}]=\mathbb{E}[Y_{ij}^{(1,3)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,3)}]=\operatorname{Var}[Y_{ij}^{(1,3)}]=4$, which corresponds to model (3.9); in Experiment IV, $\mathbb{E}[Y_{ij}^{(0,4)}]=\mathbb{E}[Y_{ij}^{(1,4)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,4)}]=\operatorname{Var}[Y_{ij}^{(1,4)}]=j^{2}$, which corresponds to model (3.10).
+For the smooth tests, the results of Experiments III and IV are broadly consistent with those of Experiments I and II (see Tables 711 and Figures 34), further supporting the validity of our unified testing framework.
+For the other three classical methods, a notable phenomenon arises in Experiment IV: under $H_{0}$, the Jarque–Bera test exhibits undersizing, suggesting that it is not well suited for models with heterogeneous group variances.
+In contrast, our proposed test remains well-calibrated and maintains stable size control.
+To further assess robustness with respect to the number of groups $J$, we also consider a variant of Experiment III with $J=10$ (denoted as Experiment III).
+The results corroborate our main conclusions; see Section A.10 of the Online Appendix for details.
+In summary, the experiments demonstrate that the proposed smooth tests in Section 3 maintain the nominal significance level well under the null, and exhibit high power under the alternatives except when $\mathbb{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$.
 The data-driven tests with the approximated null distribution $H(x)$ reliably control the Type I error and achieve excellent power performance.
-Table 11: Empirical rejection rates in Experiment IV, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
-$H_{0}$ $H_{1}$
-$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
-$10$ 0.034 0.044 0.040 0.030 0.028 0.062 0.042 0.040 0.998 0.998 0.996 0.994 0.998 0.998
-2020 0.062 0.064 0.060 0.056 0.058 0.060 0.050 0.034 1 1 1 1 1 1
-$30$ 0.040 0.040 0.038 0.042 0.048 0.060 0.044 0.034 1 1 1 1 1 1
-$40$ 0.038 0.042 0.044 0.028 0.032 0.052 0.040 0.024 1 1 1 1 1 1
-$50$ 0.044 0.042 0.042 0.042 0.044 0.060 0.048 0.036 1 1 1 1 1 1
-$60$ 0.062 0.080 0.052 0.064 0.074 0.082 0.066 0.016 1 1 1 1 1 1
-$70$ 0.062 0.066 0.060 0.050 0.062 0.052 0.042 0.050 1 1 1 1 1 1
-$80$ 0.046 0.048 0.042 0.058 0.036 0.040 0.038 0.042 1 1 1 1 1 1
-$90$ 0.060 0.058 0.050 0.052 0.046 0.052 0.042 0.032 1 1 1 1 1 1
-$100$ 0.052 0.038 0.046 0.058 0.042 0.054 0.048 0.018 1 1 1 1 1 1
-$110$ 0.040 0.034 0.038 0.032 0.040 0.052 0.046 0.032 1 1 1 1 1 1
-$120$ 0.056 0.044 0.066 0.046 0.046 0.066 0.058 0.040 1 1 1 1 1 1
-$130$ 0.052 0.058 0.054 0.056 0.050 0.058 0.052 0.028 1 1 1 1 1 1
-$140$ 0.050 0.060 0.054 0.042 0.040 0.074 0.068 0.028 1 1 1 1 1 1
-$150$ 0.056 0.064 0.054 0.058 0.060 0.054 0.052 0.028 1 1 1 1 1 1
-Table 12: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment IV
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-$10$ 0.972 0.024 0.004 0 0
-2020 0.986 0.012 0 0.002 0
-$30$ 0.980 0.018 0.002 0 0
-$40$ 0.992 0.008 0 0 0
-$50$ 0.990 0.010 0 0 0
-$60$ 0.996 0.004 0 0 0
-$70$ 0.994 0.006 0 0 0
-$80$ 0.996 0.004 0 0 0
-$90$ 0.996 0.004 0 0 0
-$100$ 0.990 0.010 0 0 0
-$110$ 0.994 0.006 0 0 0
-$120$ 0.992 0.006 0.002 0 0
-$130$ 0.992 0.008 0 0 0
-$140$ 0.986 0.014 0 0 0
-$150$ 0.998 0.002 0 0 0
-Table 13: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment IV
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-$10$ 0 0.990 0.008 0.002 0
-2020 0 0.982 0.018 0 0
-$30$ 0 0.990 0.008 0 0.002
-$40$ 0 0.994 0.006 0 0
-$50$ 0 0.996 0.004 0 0
-$60$ 0 0.996 0.004 0 0
-$70$ 0 0.994 0.006 0 0
-$80$ 0 0.996 0.004 0 0
-$90$ 0 1 0 0 0
-$100$ 0 0.998 0.002 0 0
-$110$ 0 0.996 0.004 0 0
-$120$ 0 1 0 0 0
-$130$ 0 1 0 0 0
-$140$ 0 0.988 0.012 0 0
-$150$ 0 1 0 0 0
-6 Discussion of other ANOVA models
-In this section, we briefly discuss smooth tests for normality of other ANOVA models beyond those in Section 3. The testing approach above is also applicable to these models.
-6.1 One-way fixed effects model with heterogeneous means and variances
-We first consider the one-way ANOVA model with potentially different group means and variances, given by
-$Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J.$ (6.1)
-Here $\{\mu_{j}\}_{j=1}^{J}$ and $\{\sigma_{j}\}_{j=1}^{J}$ denote the group-specific means and standard deviations.
-Owing to heterogeneity across groups, the estimators of each group mean and variance rely solely on the corresponding observations, in the sense that $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij},\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ for $j=1,\ldots,J$.
-Accordingly, we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}_{j}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$.
-To implement the smooth test, one may follow the procedure in Section 3 and derive an expression for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$, that accounts for the estimation effects.
-The test statistic is then constructed as a quadratic form of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$, normalized by the inverse of its asymptotic covariance matrix.
-Under mild conditions, an asymptotic $\chi^{2}$ test can be established for model (6.1), thereby extending the unified framework to this case.
-6.2 One-way random effects model
-Next, we consider the one-way random effects model:
-$Y_{ij}=\mu+A_{j}+\varepsilon_{ij}=\mu+A_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (6.2)
-where the random effects $\{A_{j}\}_{j=1}^{J}$ are i.i.d.
-with $\mathrm{E}[A_{j}]=0$, $\operatorname{Var}[A_{j}]=\sigma_{a}^{2}$, and are independent of the errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$.
-We estimate $\sigma^{2}$ by
-$\hat{\sigma}^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\bar{Y}_{\cdot j}\right)^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\hat{\varepsilon}_{ij}^{2},$
-where $\bar{Y}_{\cdot j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\varepsilon}_{ij}=Y_{ij}-\bar{Y}_{\cdot j}$.
-Note that $\bar{Y}_{\cdot j}$ coincides with $\hat{\mu}_{j}$ in Section 3.2, so the expression for $\hat{\sigma}^{2}$ is identical to that in Section 3.2.
-Consequently, the methodology in Section 3.2 can be directly applied to test the normality of $e_{ij}$ in (6.2), since the random effects $A_{j}$ are removed by within-group centering.
-6.3 Two-way ANOVA models
-We now extend the discussion to two-way ANOVA models. Consider the two-way fixed effects model:
-$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\big(\mu+\alpha_{j_{1}}+\beta_{j_{2}}+\gamma_{j_{1}j_{2}}\big)+\sigma e_{ij_{1}j_{2}},$ (6.3)
-subject to the identifiability constraints
-$\sum_{j_{1}=1}^{J_{1}}\alpha_{j_{1}}=\sum_{j_{2}=1}^{J_{2}}\beta_{j_{2}}=\sum_{j_{1}=1}^{J_{1}}\gamma_{j_{1}j_{2}}=\sum_{j_{2}=1}^{J_{2}}\gamma_{j_{1}j_{2}}=0.$
-Inspired by the methodology developed in Section 3, we can construct analogous estimators for $\mu$, $\{\alpha_{j_{1}}\}_{j_{1}=1}^{J_{1}}$, $\{\beta_{j_{2}}\}_{j_{2}=2}^{J_{2}}$, $\{\gamma_{j_{1}j_{2}}\}_{j_{1}=1,j_{2}=1}^{J_{1},J_{2}}$, and $\sigma^{2}$, and derive the corresponding test statistic.
-The theoretical framework in Section 3 remains valid in this setting.
-Similarly, consider the two-way random effects model:
-$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\big(\mu+A_{j_{1}}+B_{j_{2}}+D_{j_{1}j_{2}}\big)+\sigma e_{ij_{1}j_{2}},$ (6.4)
-where the i.i.d.
-random effects $\{A_{j_{1}}\}$, $\{B_{j_{2}}\}$, $\{D_{j_{1}j_{2}}\}$ and the errors $\{e_{ij}\}$ are mutually independent.
-To test the normality of $e_{ij}$ in (6.4), one can apply the strategy of Section 6.2, first eliminating the random effects by within-cell centering and then constructing the test statistic based on the residuals.
-To conclude, the methodology proposed in Section 3 can be seamlessly extended to more general ANOVA settings.
-This unified framework offers a flexible and effective approach to testing for normality across diverse linear models.
+In contrast with the three classical methods, our approach is supported by solid theoretical guarantees and demonstrates uniformly stable and competitive performance across all settings considered.
+Table 7: Empirical rejection rates under $H_{0}$ in Experiment III. The empirical power under $H_{1}$ equals 1 for all sample sizes and all considered tests.
+Smooth tests SW JB KS
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
+10 0.050 0.046 0.048 0.048 0.052 0.076 0.050 0.046 0.038 0
+20 0.072 0.060 0.050 0.046 0.050 0.054 0.048 0.064 0.048 0
+30 0.054 0.046 0.054 0.038 0.056 0.074 0.062 0.064 0.062 0
+40 0.046 0.048 0.060 0.054 0.048 0.054 0.050 0.050 0.046 0
+50 0.044 0.060 0.052 0.052 0.048 0.040 0.038 0.042 0.050 0
+60 0.066 0.044 0.048 0.052 0.058 0.048 0.042 0.048 0.052 0
+70 0.044 0.048 0.038 0.046 0.048 0.072 0.064 0.052 0.044 0
+80 0.050 0.044 0.048 0.038 0.042 0.074 0.068 0.042 0.044 0
+90 0.040 0.060 0.046 0.054 0.046 0.048 0.042 0.058 0.046 0
+100 0.062 0.056 0.042 0.050 0.048 0.058 0.052 0.048 0.044 0
+110 0.060 0.046 0.066 0.052 0.050 0.068 0.054 0.058 0.060 0
+120 0.034 0.044 0.038 0.038 0.036 0.058 0.056 0.046 0.052 0
+130 0.038 0.050 0.042 0.052 0.050 0.058 0.048 0.052 0.056 0
+140 0.038 0.050 0.048 0.042 0.044 0.044 0.040 0.058 0.068 0
+150 0.064 0.056 0.048 0.058 0.058 0.046 0.042 0.046 0.048 0
+Table 8: Empirical frequency of $\widehat{K}$ in Experiment III
+Under $H_{0}$ Under $H_{1}$
+$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
+10 0.976 0.024 0 0 0 0.032 0.004 0 0.418 0.546
+20 0.974 0.018 0.006 0 0.002 0 0 0 0.314 0.686
+30 0.986 0.014 0 0 0 0 0 0 0.174 0.826
+40 0.990 0.010 0 0 0 0 0 0 0.098 0.902
+50 0.998 0.002 0 0 0 0 0 0 0.064 0.936
+60 0.988 0.012 0 0 0 0 0 0 0.036 0.964
+70 0.992 0.008 0 0 0 0 0 0 0.022 0.978
+80 0.990 0.008 0.002 0 0 0 0 0 0.018 0.982
+90 0.988 0.012 0 0 0 0 0 0 0.008 0.992
+100 0.996 0.004 0 0 0 0 0 0 0.008 0.992
+110 0.994 0.004 0.002 0 0 0 0 0 0.002 0.998
+120 0.992 0.006 0.002 0 0 0 0 0 0 1
+130 0.998 0.002 0 0 0 0 0 0 0 1
+140 0.992 0.008 0 0 0 0 0 0 0.002 0.998
+150 0.996 0.004 0 0 0 0 0 0 0 1
+Figure 3: The sample means and error bars of $\widehat{K}$ in Experiment III
+Figure 4: The sample means and error bars of $\widehat{K}$ in Experiment IV
+Table 9: Empirical rejection rates under $H_{0}$ in Experiment IV
+Smooth tests SW JB KS
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
+10 0.034 0.044 0.040 0.030 0.028 0.062 0.042 0.046 0.008 0
+20 0.062 0.064 0.060 0.056 0.058 0.060 0.050 0.038 0.026 0.002
+30 0.040 0.040 0.038 0.042 0.048 0.060 0.044 0.042 0.032 0
+40 0.038 0.042 0.044 0.028 0.032 0.052 0.040 0.042 0.034 0
+50 0.044 0.042 0.042 0.042 0.044 0.060 0.048 0.052 0.048 0
+60 0.062 0.080 0.052 0.064 0.074 0.082 0.066 0.056 0.044 0
+70 0.062 0.066 0.060 0.050 0.062 0.052 0.042 0.040 0.046 0
+80 0.046 0.048 0.042 0.058 0.036 0.040 0.038 0.044 0.046 0
+90 0.060 0.058 0.050 0.052 0.046 0.052 0.042 0.048 0.046 0
+100 0.052 0.038 0.046 0.058 0.042 0.054 0.048 0.062 0.046 0
+110 0.040 0.034 0.038 0.032 0.040 0.052 0.046 0.042 0.034 0
+120 0.056 0.044 0.066 0.046 0.046 0.066 0.058 0.052 0.038 0
+130 0.052 0.058 0.054 0.056 0.050 0.058 0.052 0.068 0.050 0
+140 0.050 0.060 0.054 0.042 0.040 0.074 0.068 0.062 0.058 0
+150 0.056 0.064 0.054 0.058 0.060 0.054 0.052 0.042 0.046 0
+Table 10: Empirical rejection rates under $H_{1}$ in Experiment IV
+Smooth tests SW JB KS
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
+10 0.040 0.998 0.998 0.996 0.994 0.998 0.998 0.994 0.896 0.036
+20 0.034 1 1 1 1 1 1 1 1 0.394
+30 0.034 1 1 1 1 1 1 1 1 0.834
+40 0.024 1 1 1 1 1 1 1 1 0.978
+50 0.036 1 1 1 1 1 1 1 1 1
+60 0.016 1 1 1 1 1 1 1 1 1
+70 0.050 1 1 1 1 1 1 1 1 1
+80 0.042 1 1 1 1 1 1 1 1 1
+90 0.032 1 1 1 1 1 1 1 1 1
+100 0.018 1 1 1 1 1 1 1 1 1
+110 0.032 1 1 1 1 1 1 1 1 1
+120 0.040 1 1 1 1 1 1 1 1 1
+130 0.028 1 1 1 1 1 1 1 1 1
+140 0.028 1 1 1 1 1 1 1 1 1
+150 0.028 1 1 1 1 1 1 1 1 1
+Table 11: Empirical frequency of $\widehat{K}$ in Experiment IV
+Under $H_{0}$ Under $H_{1}$
+$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
+10 0.972 0.024 0.004 0 0 0 0.990 0.008 0.002 0
+20 0.986 0.012 0 0.002 0 0 0.982 0.018 0 0
+30 0.980 0.018 0.002 0 0 0 0.990 0.008 0 0.002
+40 0.992 0.008 0 0 0 0 0.994 0.006 0 0
+50 0.990 0.010 0 0 0 0 0.996 0.004 0 0
+60 0.996 0.004 0 0 0 0 0.996 0.004 0 0
+70 0.994 0.006 0 0 0 0 0.994 0.006 0 0
+80 0.996 0.004 0 0 0 0 0.996 0.004 0 0
+90 0.996 0.004 0 0 0 0 1 0 0 0
+100 0.990 0.010 0 0 0 0 0.998 0.002 0 0
+110 0.994 0.006 0 0 0 0 0.996 0.004 0 0
+120 0.992 0.006 0.002 0 0 0 1 0 0 0
+130 0.992 0.008 0 0 0 0 1 0 0 0
+140 0.986 0.014 0 0 0 0 0.988 0.012 0 0
+150 0.998 0.002 0 0 0 0 1 0 0 0
+6 An empirical application
+In this section, we illustrate the proposed smooth tests using data from the OECD Programme for International Student Assessment (PISA) 2018.
+PISA is a large-scale international assessment coordinated by the Organisation for Economic Co-operation and Development (OECD) that evaluates educational systems by measuring students’ learning outcomes and school characteristics across countries (schleicher2019pisa).
+The dataset used in this study is publicly available at https://www.kaggle.com/datasets/dilaraahan/pisa-2018-school-questionnaire.
+Since PISA data are widely used to compare educational resources across countries, statistical inference for group means is commonly employed, the validity of which depends crucially on distributional assumptions such as normality.
+However, educational indicators, including school size and student–teacher ratios, often exhibit substantial cross-country heterogeneity and dispersion, and prior studies have documented skewed distributions in school-level resource measures (hanushek2011economics; chandir2022student).
+Deviations from normality may therefore affect variance estimation and the reliability of ANOVA-based inference, making it important to formally assess this assumption.
+We consider two response variables, STRATIO (student–teacher ratio) and SCHSIZE (school size), measured at the school level, and use CNTRYID (country identifier) to define the grouping structure.
+Table 12 summarizes key sample characteristics after removing missing values, including the total number of observations $N$, the number of groups $J$, and the minimum group size $\min_{j}N_{j}$.
+The resulting group structure satisfies the conditions required by the asymptotic framework in Section 3 (see Table 1), thereby supporting the application of the proposed models and normality tests.
+For each response variable (STRATIO and SCHSIZE) under the grouping structure defined by CNTRYID, we fit the ANOVA models (3.1), (3.9), and (3.10), and apply the corresponding smooth tests for normality.
+For each model, we consider two implementations of the smooth test: one with fixed $K=4$, and one based on the data-driven procedure with $D=5$.
+Table 12: Summary of the response and grouping variables and sample characteristics in the PISA 2018 dataset
+Response variable Grouping variable NN $J$ $\min_{j}N_{j}$
+STRATIO CNTRYID 18042 76 39
+SCHSIZE CNTRYID 18321 76 42
+The results show that, for both response variables, STRATIO and SCHSIZE, the null hypothesis of normality is rejected at the $0.1\%$ significance level across all three model specifications and both implementations of the proposed tests.
+This uniform rejection pattern indicates strong robustness of the findings.
+Overall, there is compelling evidence that the distributions of STRATIO and SCHSIZE deviate substantially from normality.
+These deviations have important implications for PISA data, where ANOVA models are commonly used to compare educational resources across countries.
+For variables such as student–teacher ratios and school size, pronounced skewness or heavy-tailed behavior likely reflects substantial heterogeneity in educational systems, with a small number of countries or schools exhibiting extreme values.
+In such settings, classical $F$-tests may be sensitive to these distributional features, potentially leading to distorted inference when comparing group means.
+Consequently, ANOVA-based comparisons of educational resources should be interpreted with caution, and more robust approaches, such as transformations, rank-based methods, or other distribution-free procedures, may be preferable in practice.
+More broadly, these findings underscore the importance of formally assessing distributional assumptions prior to conducting inference in cross-country educational studies.
 7 Concluding remarks
-In this paper, we establish Neyman’s smooth tests for assessing the normality assumption in ANOVA models.
-For three types of one-way fixed effects models, we derive the asymptotic properties of the proposed tests and validate their finite-sample performance through extensive simulations.
+In this paper, we propose and examine data-driven Neyman’s smooth tests for assessing the normality assumption in ANOVA models with a potentially diverging number of groups.
+For three types of one-way fixed effects models, we derive the asymptotic properties of the proposed tests and validate their finite-sample performance through extensive numerical studies.
 Our results provide a rigorous and practical tool for evaluating normality in a broad range of ANOVA settings.
@@
 Second, the assumption of independent random errors can be relaxed.
-For example, repeated clinical trials often involve correlated errors (Ma et al.,, 2012; Schober and Vetter,, 2018; Langenberg et al.,, 2022).
+For example, repeated clinical trials often involve correlated errors (ma2012beyond; schober2018repeated; langenberg2022repeated).
 In such cases, our smooth test framework may be adapted by incorporating estimated correlation structures.
-Finally, for the data-driven selection rule, it is of theoretical interest to explore the scenario where the upper bound $D=D(N)$ diverges with $N$, that is, $D\to\infty$ as $N\to\infty$.
-Appendix A Appendix
-We provide the proof of our main theoretical results, along with some additional numerical results, in this Appendix.
-The symbol $\lesssim$” means that the left side is bounded by a positive constant times the right side.
-The symbol $\asymp$” means that both sides are asymptotically equivalent.
-Proof of Theorem 1.
-We first study the estimation effects of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}-\sigma^{2}$. Note that
-$\hat{\mu}-\mu=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\mu\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{\sigma}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.1)
-Under $H_{0}$, $e_{ij}\sim\mathcal{N}(0,1)i.i.d$., then $\sqrt{N}(\hat{\mu}-\mu)\sim\mathcal{N}(0,\sigma^{2})$, and $\hat{\mu}-\mu=O_{p}(N^{-1/2})$ as $N\to\infty$. Besides,
-$\hat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}\right)^{2}-\sigma^{2}\right)$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\left(\hat{\mu}-\mu\right)^{2}$
+Finally, for the selection rule (4.1), it is of theoretical interest to explore the scenario where the upper bound $D=D(N)$ diverges slowly with $N$, that is, $D\to\infty$ as $N\to\infty$.
+A.1 Proof of Theorem 1
+We first study the estimation effects of $\widehat{\mu}-\mu$ and $\widehat{\sigma}^{2}-\sigma^{2}$. Note that
+$\widehat{\mu}-\mu=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\mu\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{\sigma}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.1)
+Under $H_{0}$, $e_{ij}\sim\mathcal{N}(0,1)i.i.d$., then $\sqrt{N}(\widehat{\mu}-\mu)\sim\mathcal{N}(0,\sigma^{2})$, and $\widehat{\mu}-\mu=O_{p}(N^{-1/2})$ as $N\to\infty$. Besides,
+$\widehat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\widehat{\mu}\right)^{2}-\sigma^{2}\right)$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\left(\widehat{\mu}-\mu\right)^{2}$
 $=$ $\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+O_{p}\left(\frac{1}{N}\right).$ (A.2)
-Under $H_{0}$, by CLT, $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\mathcal{N}(0,2\sigma^{4})$ as $N\to\infty$, then $\hat{\sigma}^{2}-\sigma^{2}=O_{p}(N^{-1/2})$.
-Now we deal with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$. By second order Taylor expansion of $\pi_{k}(\hat{Z}_{ij})$ with respect to $\hat{Z}_{ij}$ at $Z_{ij}$, we obtain
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left(\hat{Z}_{ij}-Z_{ij}\right)$
-$+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\tilde{Z}_{ij}\right)\left(\hat{Z}_{ij}-Z_{ij}\right)^{2},$ (A.3)
+Under $H_{0}$, by CLT, $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})\overset{d}{\to}\mathcal{N}(0,2\sigma^{4})$ as $N\to\infty$, then $\widehat{\sigma}^{2}-\sigma^{2}=O_{p}(N^{-1/2})$.
+Now we deal with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$.
+By second order Taylor expansion of $\pi_{k}(\widehat{Z}_{ij})$ with respect to $\widehat{Z}_{ij}$ at $Z_{ij}$, we obtain
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left(\widehat{Z}_{ij}-Z_{ij}\right)$
+$+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\widetilde{Z}_{ij}\right)\left(\widehat{Z}_{ij}-Z_{ij}\right)^{2},$ (A.3)
 […]
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2},$ (A.4)
-where $\tilde{\varepsilon}_{ij}/\tilde{\sigma}$ an intermediate point between $^{\hat{\varepsilon}_{ij}}/{\hat{\sigma}}$ and $\varepsilon_{ij}/\sigma$, whose exact value may vary across different scenarios.
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2},$ (A.4)
+where $\widetilde{\varepsilon}_{ij}/\widetilde{\sigma}$ an intermediate point between $^{\widehat{\varepsilon}_{ij}}/{\widehat{\sigma}}$ and $\varepsilon_{ij}/\sigma$, whose exact value may vary across different scenarios.
 We focus on the first term in (A.4). Straightforward calculation leads to
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}$
-$=$ $-\frac{\hat{\mu}-\mu}{\hat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\frac{\hat{\sigma}^{2}-\sigma^{2}}{\hat{\sigma}(\hat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
-$=$ $-\frac{\hat{\mu}-\mu}{\sigma}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]-\frac{\hat{\sigma}^{2}-\sigma^{2}}{2\sigma^{2}}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]$
-$+o_{p}\left(\hat{\mu}-\mu\right)+o_{p}\left(\hat{\sigma}^{2}-\sigma^{2}\right),$ (A.5)
-where the last step follows due to the law of large numbers (LLN) for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)]+o_{p}(1)$ and $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)e]+o_{p}(1)$ as $N\to\infty$, as well as the consistency of $\hat{\sigma}$ to $\sigma$ from the arguments above.
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}$
+$=$ $-\frac{\widehat{\mu}-\mu}{\widehat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\frac{\widehat{\sigma}^{2}-\sigma^{2}}{\widehat{\sigma}(\widehat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
+$=$ $-\frac{\widehat{\mu}-\mu}{\sigma}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]-\frac{\widehat{\sigma}^{2}-\sigma^{2}}{2\sigma^{2}}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]$
+$+o_{p}\left(\widehat{\mu}-\mu\right)+o_{p}\left(\widehat{\sigma}^{2}-\sigma^{2}\right),$ (A.5)
+where the last step follows due to the law of large numbers (LLN) for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})=\mathbb{E}[\dot{\pi}_{k}(Z)\phi(e)]+o_{p}(1)$ and $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}=\mathbb{E}[\dot{\pi}_{k}(Z)\phi(e)e]+o_{p}(1)$ as $N\to\infty$, as well as the consistency of $\widehat{\sigma}$ to $\sigma$ from the arguments above.
 Under $H_{0}$, through integration by parts, we have
-$\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z=c_{1k},$
+$\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z=c_{1k},$
 and
-$\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{0}^{1}\pi_{k}(z)\left[\left(\Phi^{-1}(z)\right)^{2}-1\right]\mathrm{d}z=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z=c_{2k},$
+$\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{0}^{1}\pi_{k}(z)\left[\left(\Phi^{-1}(z)\right)^{2}-1\right]\mathrm{d}z=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z=c_{2k},$
 where we use the fact that $\mathrm{d}\Phi^{-1}(z)=\mathrm{d}z/\phi(\Phi^{-1}(z))$ and $\int_{0}^{1}\pi_{k}(z)\mathrm{d}z=0$. Thus, along with (A.1) and (A.2), (A.5) can be re-expressed as
@@
 It suffices to show that
-$R_{N1}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\tilde{Z}_{ij}\right)\left(\hat{Z}_{ij}-Z_{ij}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.7)
+$R_{N1}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\widetilde{Z}_{ij}\right)\left(\widehat{Z}_{ij}-Z_{ij}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.7)
 and
-$R_{N2}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.8)
+$R_{N2}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.8)
 For $R_{N1}$, by the boundedness of $\ddot{\pi}_{k}$ (in Assumption 3.2) and $\phi$,
-$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left[\Phi\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}\right)-\Phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\right]^{2}$
-$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\phi^{2}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
-$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
-$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}.$
+$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left[\Phi\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}\right)-\Phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\right]^{2}$
+$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\phi^{2}\left(\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
+$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
+$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{2}.$
 By the inequality $(a-b)^{2}\leq 2(a^{2}+b^{2})$ for any $a,b\in\mathbb{R}$,
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
-$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
-$\asymp$ $\left(\hat{\mu}-\mu\right)^{2}+(\hat{\sigma}^{2}-\sigma^{2})^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{2}$
+$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{2}$
+$\asymp$ $\left(\widehat{\mu}-\mu\right)^{2}+(\widehat{\sigma}^{2}-\sigma^{2})^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
 $=$ $O_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{N}\right)\left(\sigma^{2}+o_{p}(1)\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 where the last step is the result of (A.1), (A.2) and LLN for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}=\sigma^{2}+o_{p}(1)$. For $R_{N2}$, we have
-$|R_{N2}|$ $=\left|\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left[-\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\phi\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\right]\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}\right|$
-$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left|\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right|\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
-$\leq\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}\right)^{1/2}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}\right)^{1/2},$
+$|R_{N2}|$ $=\left|\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left[-\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\phi\left(\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right)\right]\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}\right|$
+$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left|\frac{\widetilde{\varepsilon}_{ij}}{\widetilde{\sigma}}\right|\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
+$\leq\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\widetilde{\varepsilon}_{ij}^{2}}{\widetilde{\sigma}^{2}}\right)^{1/2}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}\right)^{1/2},$
 where the first “==” is due to $\dot{\phi}(x)=-x\phi(x)$, and the last $\leq$” is due to Cauchy–Schwarz inequality. Similar arguments as above yield that
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}$
-$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}\right]$
-$\asymp$ $\left(\hat{\mu}-\mu\right)^{4}+\left(\hat{\sigma}^{2}-\sigma^{2}\right)^{4}\left(\mathrm{E}\left[e^{4}\right]+o_{p}(1)\right)=O_{p}\left(\frac{1}{N^{2}}\right),$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{4}$
+$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{4}\right]$
+$\asymp$ $\left(\widehat{\mu}-\mu\right)^{4}+\left(\widehat{\sigma}^{2}-\sigma^{2}\right)^{4}\left(\mathbb{E}\left[e^{4}\right]+o_{p}(1)\right)=O_{p}\left(\frac{1}{N^{2}}\right),$
 and
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}$ $\leq\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{{\varepsilon}_{ij}}{{\sigma}}\right)^{2}$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\widetilde{\varepsilon}_{ij}^{2}}{\widetilde{\sigma}^{2}}$ $\leq\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{{\varepsilon}_{ij}}{{\sigma}}\right)^{2}$
 $=1+\left(1+o_{p}(1)\right)=O_{p}(1).$ (A.9)
 Thus, $|R_{N2}|=O_{p}(N^{-1})=o_{p}(N^{-1/2})$. Hence (A.7) and (A.8) are verified. Combining (A.3), (A.4), (A.5), (A.6), (A.7) and (A.8), we complete the proof of Theorem 1. $\square$
-Proof of Theorem 2.
+A.2 Proof of Theorem 2
 The proof of Theorem 2 is similar to that of Theorem 1.
-Under $H_{1}$, (A.1) still holds, and $\sqrt{N}(\hat{\mu}-\mu)\to_{d}\mathcal{N}(0,\sigma^{2})$ since $e_{ij}\sim F\neq\Phi$.
-Analogously, (A.2) still holds, and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\sigma^{2}\mathcal{N}(0,\mathrm{E}[e^{4}]-1)$.
+Under $H_{1}$, (A.1) still holds, and $\sqrt{N}(\widehat{\mu}-\mu)\overset{d}{\to}\mathcal{N}(0,\sigma^{2})$ since $e_{ij}\sim F\neq\Phi$.
+Analogously, (A.2) still holds, and $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})\overset{d}{\to}\sigma^{2}\mathcal{N}(0,\mathbb{E}[e^{4}]-1)$.
 We start from (A.3). Under $H_{1}$, the arguments and results in (A.3), (A.4) and (A.5) still hold. Unlike (A.6), here in the context of $H_{1}$ the last line of (A.5) should be expressed as
@@
 From (3.5), we know that
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}(1),$
-as $N\to\infty$, which implies $\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$. Furthermore,
-$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)$
-$=$ $\sqrt{N}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\bm{\Sigma}_{K}^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}(1),$
+as $N\to\infty$, which implies $\widehat{\Psi}_{K}^{2}\overset{p}{\to}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$. Furthermore,
+$\sqrt{N}\left(\widehat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)$
+$=$ $\sqrt{N}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\bm{\Sigma}_{K}^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right]$
 $-\sqrt{N}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$
-$=$ $2a^{\top}\bm{\Sigma}_{K}^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}\bm{\Sigma}_{K}^{-1}$
-$\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.11)
+$=$ $2a^{\top}\bm{\Sigma}_{K}^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}\bm{\Sigma}_{K}^{-1}$
+$\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.11)
 The following part demonstrates the asymptotic normality of
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\hat{Z}_{ij})-\bm{a}_{K}\right).$
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\widehat{Z}_{ij})-\bm{a}_{K}\right).$
 Indeed, the $k$-th component in it is
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\pi_{k}(\hat{Z}_{ij})-a_{k}\right)=\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\zeta_{k}(e_{ij})-\mathrm{E}\left[\zeta_{k}(e_{ij})\right]\right)+o_{p}(1),$
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\pi_{k}(\widehat{Z}_{ij})-a_{k}\right)=\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\zeta_{k}(e_{ij})-\mathbb{E}\left[\zeta_{k}(e_{ij})\right]\right)+o_{p}(1),$
 where
@@
 Then, by CLT,
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\hat{Z}_{ij})-\bm{a}_{K}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Xi}_{K}\right),$
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\widehat{Z}_{ij})-\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}_{K}\left(\bm{0},\bm{\Xi}_{K}\right),$
 with $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ given by
-$\xi_{kl}=$ $\mathrm{E}\left[\zeta_{k}(e)\zeta_{l}(e)\right]$
-$=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
-$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
+$\xi_{kl}=$ $\mathbb{E}\left[\zeta_{k}(e)\zeta_{l}(e)\right]$
+$=$ $\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
+$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathbb{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathbb{E}\left[e^{4}\right]-1\right].$
 Along with (A.11), (3.6) in Theorem 2 is validated. This completes the proof of Theorem 2. $\square$
-Proof of Theorem 3.
+A.3 Proof of Theorem 3
 The proof of Theorem 3 is similar to that of Theorems 1 and 2.
@@
 $f(x)=(1-\delta_{N})\phi(x)+\delta_{N}q(x).$
-We first examine the convergence of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}-\sigma^{2}$ as before.
-For $\hat{\mu}-\mu$, (A.1) still holds under $H_{1L}$, and $\sqrt{N}(\hat{\mu}-\mu)\to_{d}\mathcal{N}(0,\sigma^{2})$ as $N\to\infty$.
-For $\hat{\sigma}^{2}-\sigma^{2}$, under $H_{1L}$, (A.2) still holds, and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\mathcal{N}(0,2\sigma^{4})$  [^1] Here under $H_{1L}$, the asymptotic normality of $\sqrt{N}(\hat{\mu}-\mu)$ and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})$ are both derived from CLT for triangular arrays since the distribution of $e_{ij}$ changes over $N$.
-It can be verified that the conditions of Lindeberg–Feller CLT hold for both random sequences, see, for example, Proposition 2.27 of Van der Vaart, (2000) .
+We first examine the convergence of $\widehat{\mu}-\mu$ and $\widehat{\sigma}^{2}-\sigma^{2}$ as before.
+For $\widehat{\mu}-\mu$, (A.1) still holds under $H_{1L}$, and $\sqrt{N}(\widehat{\mu}-\mu)\overset{d}{\to}\mathcal{N}(0,\sigma^{2})$ as $N\to\infty$.
+For $\widehat{\sigma}^{2}-\sigma^{2}$, under $H_{1L}$, (A.2) still holds, and $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})\overset{d}{\to}\mathcal{N}(0,2\sigma^{4})$  [^1] Here under $H_{1L}$, the asymptotic normality of $\sqrt{N}(\widehat{\mu}-\mu)$ and $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})$ are both derived from CLT for triangular arrays since the distribution of $e_{ij}$ changes over $N$.
+It can be verified that the conditions of Lindeberg–Feller CLT hold for both random sequences, see, for example, Proposition 2.27 of van2000asymptotic .
 since
-$\mathrm{E}\left[e^{4}\right]=$ $\int_{-\infty}^{\infty}x^{4}f(x)\mathrm{d}x$
+$\mathbb{E}\left[e^{4}\right]=$ $\int_{-\infty}^{\infty}x^{4}f(x)\mathrm{d}x$
 $=$ $(1-\delta_{N})\int_{-\infty}^{\infty}x^{4}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}x^{4}q(x)\mathrm{d}x$
@@
 Next, we clarify the relationship of the denoted constants.
-Recall the definitions of the constants $d_{1k}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)f(x)\mathrm{d}x$ and $d_{2k}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)e]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)xf(x)\mathrm{d}x$.
+Recall the definitions of the constants $d_{1k}=\mathbb{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)f(x)\mathrm{d}x$ and $d_{2k}=\mathbb{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)e]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)xf(x)\mathrm{d}x$.
 Then under $H_{1L}$, as $N\to\infty$, we have
@@
 and
-$\mathrm{E}[\pi_{k}(Z)]=$ $\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)f(x)\mathrm{d}x$
+$\mathbb{E}[\pi_{k}(Z)]=$ $\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)f(x)\mathrm{d}x$
 $=$ $\left(1-\delta_{N}\right)\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)q(x)\mathrm{d}x$
@@
 Now we start from (A.3) again.
-Under $H_{1L}$, we still have (A.3), (A.4) and (A.5).
-(A.6) also holds as a result of (A.12) and (A.13).
+Under $H_{1L}$, we still have (A.3), (A.4), and (A.5).
+Also, (A.6) holds as a result of (A.12) and (A.13).
 The remainder terms $R_{N1}$ in (A.7) and $R_{N2}$ in (A.8) are still $O_{p}(N^{-1/2})$.
-Combining (A.3), (A.4), (A.5), (A.6), (A.7), (A.8) and (A.14), we obtain (3.8) in Theorem 3.
-It suffices to derive the asymptotic normality of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$. Under $H_{1L}$, straightforward calculation yields that
-$\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]=$ $\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{l}(\Phi(x))f(x)\mathrm{d}x$
+Combining (A.3), (A.4), (A.5), (A.6), (A.7), (A.8), and (A.14), we obtain (3.8) in Theorem 3.
+It suffices to derive the asymptotic normality of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\widehat{Z}_{ij})$. Under $H_{1L}$, straightforward calculation yields that
+$\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]=$ $\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{l}(\Phi(x))f(x)\mathrm{d}x$
 $=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{k}(\Phi(x))\phi(x)\mathrm{d}x$
@@
 $=$ $\delta_{kl}+o(1),$
-$d_{3k}=$ $\mathrm{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xf(x)\mathrm{d}x$
+$d_{3k}=$ $\mathbb{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xf(x)\mathrm{d}x$
 $=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xq(x)\mathrm{d}x$
 $=$ $c_{1k}+o(1),$
-$d_{4k}=$ $\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}f(x)\mathrm{d}x$
+$d_{4k}=$ $\mathbb{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}f(x)\mathrm{d}x$
 $=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}q(x)\mathrm{d}x$
@@
 and
-$\mathrm{E}\left[e^{3}\right]=$ $\int_{-\infty}^{\infty}x^{3}f(x)\mathrm{d}x$
+$\mathbb{E}\left[e^{3}\right]=$ $\int_{-\infty}^{\infty}x^{3}f(x)\mathrm{d}x$
 $=$ $(1-\delta_{N})\int_{-\infty}^{\infty}x^{3}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}x^{3}q(x)\mathrm{d}x$
@@
 Thus, by (3.8) and CLT,
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
-$\to_{d}\mathcal{N}_{K}\left(\bm{\Delta}_{K},\bm{\Sigma}_{K}\right),$
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\widehat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\widehat{Z}_{ij}\right)\right]^{\top}$
+$\overset{d}{\to}\mathcal{N}_{K}\left(\bm{\Delta}_{K},\bm{\Sigma}_{K}\right),$
 since the asymptotic covariance matrix is still $\bm{\Sigma}_{K}$ from the calculation above. Besides,
-$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
+$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
 by definition of the non-central $\chi^{2}$ distribution. This completes the proof of Theorem 3. $\square$
-Proof of Theorem 4, 5 and 6.
+A.4 Proof of Theorem 4, 5 and 6
 The proof is similar to that of Theorems 1, 2, and 3.
 Note that in the case of different means and same variance, we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$.
-To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ without distinction of the null and alternatives.
+To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ without distinction of the null and alternatives.
 Our goal is to prove that
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]e_{ij}-\frac{\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]}{2}\left(e_{ij}^{2}-1\right)\right\}$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]e_{ij}-\frac{\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]}{2}\left(e_{ij}^{2}-1\right)\right\}$
 $+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.15)
-We first focus on the estimation effects of $\hat{\mu}_{j}-\mu_{j}$ and $\hat{\sigma}^{2}-\sigma^{2}$.
-For $\hat{\mu}_{j}-\mu_{j}$, by CLT, we have $\sqrt{N_{j}}(\hat{\mu}_{j}-\mu_{j})\to_{d}\mathcal{N}(0,\sigma^{2})$ for each $1\leq j\leq J$ as $\min\{N_{1},\ldots,N_{J}\}\to\infty$.
+We first focus on the estimation effects of $\widehat{\mu}_{j}-\mu_{j}$ and $\widehat{\sigma}^{2}-\sigma^{2}$.
+For $\widehat{\mu}_{j}-\mu_{j}$, by CLT, we have $\sqrt{N_{j}}(\widehat{\mu}_{j}-\mu_{j})\overset{d}{\to}\mathcal{N}(0,\sigma^{2})$ for each $1\leq j\leq J$ as $\min\{N_{1},\ldots,N_{J}\}\to\infty$.
 Besides, since
-$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\mathrm{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\right]=\frac{J}{N}\sigma^{2},$ (A.16)
-we have $N^{-1}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}=O_{p}(J/N)=o_{p}(N^{-1/2})$ from $J=o(N^{1/2})$. Analogously, by
-$\mathrm{E}\left[\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{1}{N_{j}^{4}}\left[3N_{j}(N_{j}-1)+N_{j}\mathrm{E}\left[e^{4}\right]\right]=\frac{3}{N_{j}^{2}}+\frac{\mathrm{E}\left[e^{4}\right]-3}{N_{j}^{3}},$
+$\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{2}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\mathbb{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\right]=\frac{J}{N}\sigma^{2},$ (A.16)
+we have $N^{-1}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{2}=O_{p}(J/N)=o_{p}(N^{-1/2})$ from $J=o(N^{1/2})$. Analogously, by
+$\mathbb{E}\left[\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{1}{N_{j}^{4}}\left[3N_{j}(N_{j}-1)+N_{j}\mathbb{E}\left[e^{4}\right]\right]=\frac{3}{N_{j}^{2}}+\frac{\mathbb{E}\left[e^{4}\right]-3}{N_{j}^{3}},$
 we obtain
-$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}\right]$
-$=$ $\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\mathrm{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\left(\frac{3}{N_{j}}+\frac{\mathrm{E}\left[e^{4}\right]-3}{N_{j}^{2}}\right),$ (A.17)
-which implies $N^{-1}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}=N^{-1}O_{p}(\sum_{j=1}^{J}N_{j}^{-1})=o_{p}(N^{-1})$ since $\sum_{j=1}^{J}N_{j}^{-1}=o_{p}(1)$.
-For $\hat{\sigma}^{2}-\sigma^{2}$, we have
-$\hat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}-\sigma^{2}\right)$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}$
+$\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{4}\right]$
+$=$ $\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\mathbb{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\left(\frac{3}{N_{j}}+\frac{\mathbb{E}\left[e^{4}\right]-3}{N_{j}^{2}}\right),$ (A.17)
+which implies $N^{-1}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{4}=N^{-1}O_{p}(\sum_{j=1}^{J}N_{j}^{-1})=o_{p}(N^{-1})$ since $\sum_{j=1}^{J}N_{j}^{-1}=o_{p}(1)$.
+For $\widehat{\sigma}^{2}-\sigma^{2}$, we have
+$\widehat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\widehat{\mu}_{j}\right)^{2}-\sigma^{2}\right)$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{2}$
 $=$ $\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
 Next, we will show that
-$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
-$=$ $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.18)
+$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
+$=$ $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.18)
 Indeed, since
-$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
+$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
 $=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)e_{i_{2}j},$
@@
 then we have
-$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{ij}$
-$+\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
+$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{ij}$
+$+\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
 $\equiv$ $A_{N1}+A_{N2}.$
 For $A_{N1}$, we have
-$\mathrm{E}[A_{N1}]=\frac{J}{N}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right],$
+$\mathbb{E}[A_{N1}]=\frac{J}{N}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right],$
 and
-$\operatorname{Var}[A_{N1}]=\operatorname{Var}\left[\left(\dot{\pi}_{k}(Z)\phi\left(e\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e\right]\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}},$
-then $A_{N1}=o_{p}(N^{-1/2})$ since $J=o(N^{1/2})$ as $N\to\infty$. For $A_{N2}$, we have $\mathrm{E}[A_{N2}]=0$, and
+$\operatorname{Var}[A_{N1}]=\operatorname{Var}\left[\left(\dot{\pi}_{k}(Z)\phi\left(e\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e\right]\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}},$
+then $A_{N1}=o_{p}(N^{-1/2})$ since $J=o(N^{1/2})$ as $N\to\infty$. For $A_{N2}$, we have $\mathbb{E}[A_{N2}]=0$, and
 $\operatorname{Var}[A_{N2}]$
-$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\operatorname{Var}\left[\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right]$
-$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\mathrm{E}\left[\left(\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right)^{2}\right]$
-$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{N_{j}(N_{j}-1)}{N_{j}^{2}}\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathrm{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)$
-$=$ $\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathrm{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)\frac{1}{N^{2}}\left(J-\sum_{j=1}^{J}\frac{1}{N_{j}}\right),$
+$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\operatorname{Var}\left[\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right]$
+$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\mathbb{E}\left[\left(\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right)^{2}\right]$
+$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{N_{j}(N_{j}-1)}{N_{j}^{2}}\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathbb{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)$
+$=$ $\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathbb{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)\frac{1}{N^{2}}\left(J-\sum_{j=1}^{J}\frac{1}{N_{j}}\right),$
 then $A_{N2}=o_{p}(N^{-1/2})$. Hence, (A.18) is verified.
-Now we start from (A.3). In the case of different means and same variance, (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-e_{ij}\right)$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}$
-$=$ $-\frac{1}{\hat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
-$-\frac{\hat{\sigma}^{2}-\sigma^{2}}{\hat{\sigma}(\hat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
-$=$ $-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{1}{2\sigma^{2}}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\left(\hat{\sigma}^{2}-\sigma^{2}\right)$
-$+o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}\right)+o_{p}\left(\hat{\sigma}^{2}-\sigma^{2}\right)$
-$=$ $-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
-$-\frac{1}{2N}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.19)
+Now we start from (A.3). For model (3.9), (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-e_{ij}\right)$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}$
+$=$ $-\frac{1}{\widehat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
+$-\frac{\widehat{\sigma}^{2}-\sigma^{2}}{\widehat{\sigma}(\widehat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
+$=$ $-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{1}{2\sigma^{2}}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\left(\widehat{\sigma}^{2}-\sigma^{2}\right)$
+$+o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}\right)+o_{p}\left(\widehat{\sigma}^{2}-\sigma^{2}\right)$
+$=$ $-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
+$-\frac{1}{2N}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.19)
 where we apply (A.2), (A.18), LLN for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$, and CLT for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$ as $N\to\infty$.
 It suffices to show $R_{N1}$ (in (A.7)) and $R_{N2}$ (in (A.8)) are $o_{p}(N^{-1/2})$ in order to obtain (A.15). The technique is similar to that in the proof of Theorem 1. For $R_{N1}$, we have
-$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
-$\lesssim\frac{1}{N\sigma^{2}}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}+\left(\hat{\sigma}^{2}-\sigma^{2}\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
+$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{2}$
+$\lesssim\frac{1}{N\sigma^{2}}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{2}+\left(\widehat{\sigma}^{2}-\sigma^{2}\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
 $\asymp O_{p}\left(\frac{J}{N}\right)+O_{p}\left(\frac{1}{N}\right)\left(\sigma^{2}+o_{p}(1)\right)=O_{p}\left(\frac{J}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 by (A.16) and the assumption $J=o(N^{1/2})$. For $R_{N2}$, we first have
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}$
-$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}\right]$
-$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}+(\hat{\sigma}^{2}-\sigma^{2})^{4}\left(\mathrm{E}\left[e^{4}\right]+o_{p}(1)\right)$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}$
+$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{4}\right]$
+$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{4}+(\widehat{\sigma}^{2}-\sigma^{2})^{4}\left(\mathbb{E}\left[e^{4}\right]+o_{p}(1)\right)$
 $\asymp$ $o_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{N^{2}}\right)=o_{p}\left(\frac{1}{N}\right),$
-by (A.17). Along with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\tilde{\varepsilon}_{ij}^{2}/\tilde{\sigma}^{2}=O_{p}(1)$ from (A.9), we obtain
-$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
+by (A.17). Along with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\widetilde{\varepsilon}_{ij}^{2}/\widetilde{\sigma}^{2}=O_{p}(1)$ from (A.9), we obtain
+$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\widetilde{\varepsilon}_{ij}^{2}}{\widetilde{\sigma}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
 Thus, (A.15) is verified.
 The form of (A.15) can be further refined under $H_{0}$, $H_{1}$ and $H_{1L}$ respectively.
-For example, under $H_{0}$, we can write $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ and $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ in (A.15), and then obtain the results of Theorem 4.
+For example, under $H_{0}$, we can write $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ and $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ in (A.15), and then obtain the results of Theorem 4.
 $\square$
-Proof of Theorem 7.
-The technique is still the same as the previous ones.
-Recall that in the context of same mean and different variances, we denote $p_{j}=N_{j}/N$, $q_{j}=JN_{j}/N$ for $1\leq j\leq J$, and we further assume that as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$.
-We first tackle the estimation effects of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}$.
-For $\hat{\mu}-\mu$, note that $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ here.
+A.5 Proof of Theorem 7
+The technical arguments remain the same as in the previous cases.
+Recall that $p_{j}=N_{j}/N$ and $q_{j}=JN_{j}/N$ for $1\leq j\leq J$.
+And for model (3.10), we further assume that as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$ (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$.
+We first tackle the estimation effects of $\widehat{\mu}-\mu$ and $\widehat{\sigma}^{2}_{j}-\sigma^{2}_{j}$.
+For $\widehat{\mu}-\mu$, note that $\widehat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ here.
 Then,
-$\hat{\mu}-\mu=\frac{1}{J}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{1}{J}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.20)
-From (A.20) we obtain $\mathrm{E}[\hat{\mu}-\mu]=0$, and $\operatorname{Var}[\hat{\mu}-\mu]=J^{-2}\sum_{j=1}^{J}N_{j}^{-1}\sigma_{j}^{2}\asymp J^{-2}\sum_{j=1}^{J}N_{j}^{-1}=N^{-1}J^{-1}\sum_{j=1}^{J}q_{j}^{-1}=O_{p}(N^{-1})$.
-Then $\hat{\mu}-\mu=O_{p}(N^{-1/2})$.
-For $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}$, we have
-$\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}$ $=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}\right)^{2}-\sigma_{j}^{2}\right)=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma_{j}^{2}\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}$ (A.21)
-$=\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}$
+$\widehat{\mu}-\mu=\frac{1}{J}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{1}{J}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.20)
+From (A.20) we obtain $\mathbb{E}[\widehat{\mu}-\mu]=0$, and $\operatorname{Var}[\widehat{\mu}-\mu]=J^{-2}\sum_{j=1}^{J}N_{j}^{-1}\sigma_{j}^{2}\asymp J^{-2}\sum_{j=1}^{J}N_{j}^{-1}=N^{-1}J^{-1}\sum_{j=1}^{J}q_{j}^{-1}=O_{p}(N^{-1})$.
+Then $\widehat{\mu}-\mu=O_{p}(N^{-1/2})$.
+For $\widehat{\sigma}^{2}_{j}-\sigma^{2}_{j}$, we have
+$\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}$ $=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\widehat{\mu}\right)^{2}-\sigma_{j}^{2}\right)=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma_{j}^{2}\right)+\left(\widehat{\mu}-\mu\right)^{2}-2\left(\widehat{\mu}-\mu\right)\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}$ (A.21)
+$=\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\widehat{\mu}-\mu\right)^{2}-2\left(\widehat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}$
 $=O_{p}\left(\frac{1}{\sqrt{N_{j}}}\right)+O_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{\sqrt{N_{j}N}}\right)=O_{p}\left(\frac{1}{\sqrt{N_{j}}}\right),$
-by CLT for $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}e_{ij}\to_{d}\mathcal{N}(0,1)$ and $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)\to_{d}\mathcal{N}(0,\mathrm{E}[e^{4}]-1)$ as $N_{j}\to\infty$ and (A.20).
+by CLT for $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}e_{ij}\overset{d}{\to}\mathcal{N}(0,1)$ and $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)\overset{d}{\to}\mathcal{N}(0,\mathbb{E}[e^{4}]-1)$ as $N_{j}\to\infty$ and (A.20).
 We next prove some preliminary results such that
-$\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}+O_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.22)
+$\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}+O_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.22)
 $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
-$=$ $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.23)
-$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.24)
+$=$ $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.23)
+$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.24)
 $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}=O_{p}(1),$ (A.25)
 as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$. For (A.22), straightforward calculation leads to
-$\mathrm{E}\left[\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right]=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}},$
+$\mathbb{E}\left[\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right]=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}},$
 and
@@
 and so is (A.24). (A.25) is obtained by
-$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\}\right]=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)e]\frac{1}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}=O(1),$
+$\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\}\right]=\mathbb{E}[\dot{\pi}_{k}(Z)\phi(e)e]\frac{1}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}=O(1),$
 and
 $\operatorname{Var}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\right]=O\left(\frac{1}{N}\right).$
-Now we start from (A.3). In the case of the same mean and different variances, (A.3) and (A.4) still hold. The difference lies in the fact that the first term of (A.4) should be further expressed as
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(\frac{\varepsilon_{ij}}{\sigma_{j}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}$
-$-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}.$
+Now we start from (A.3). For model (3.10), (A.3) and (A.4) still hold. The difference lies in the fact that the first term of (A.4) should be further expressed as
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(\frac{\varepsilon_{ij}}{\sigma_{j}}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}_{j}}$
+$-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\widehat{\sigma}_{j}\sigma_{j}(\widehat{\sigma}_{j}+\sigma_{j})}.$
 $\equiv$ $D_{N1}-D_{N2}.$
 For $D_{N1}$, under $H_{0}$, we have
-$D_{N1}$ $=-\frac{\hat{\mu}-\mu}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\left(1+o_{p}(1)\right)$
+$D_{N1}$ $=-\frac{\widehat{\mu}-\mu}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\left(1+o_{p}(1)\right)$
 $=-\frac{1}{N}\left(\sum_{j=1}^{J}\frac{N}{JN_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}\right)\left(\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right)\left(1+o_{p}(1)\right)$
@@
 $=-\frac{c_{1k}}{N}\left[\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\sigma_{j}e_{ij}}{q_{j}}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.26)
-where the second last line is followed by (A.22), and $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ under $H_{0}$.
+where the second last line is followed by (A.22), and $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ under $H_{0}$.
 The last line is followed by $N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=o_{p}(1)$ using LLN.
 For $D_{N2}$, under $H_{0}$, we have
-$D_{N2}=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left\{\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}\right$.
-$\left.-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right\}$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{2}}{N_{j}\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
-$+\left(\hat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
-$-2\left(\hat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
+$D_{N2}=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left\{\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\widehat{\mu}-\mu\right)^{2}\right$.
+$\left.-2\left(\widehat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right\}$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{2}}{N_{j}\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
+$+\left(\widehat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
+$-2\left(\widehat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}\widehat{\sigma}_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
 $=$ $\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{N_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}\left(1+o_{p}(1)\right)$
-$+\left(\hat{\mu}-\mu\right)^{2}\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left(1+o_{p}(1)\right)$
-$-\left(\hat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}\sigma_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
+$+\left(\widehat{\mu}-\mu\right)^{2}\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left(1+o_{p}(1)\right)$
+$-\left(\widehat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}\sigma_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
 $=$ $\frac{c_{2k}}{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.27)
-following (A.20), (A.21), (A.23), (A.24) and (A.25), as well as $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ under $H_{0}$.
+following (A.20), (A.21), (A.23), (A.24) and (A.25), as well as $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ under $H_{0}$.
 It suffices to show $R_{N1}$ (in (A.7)) and $R_{N2}$ (in (A.8)) are $o_{p}(N^{-1/2})$ in this case. For $R_{N1}$, referring to the arguments in the proof of Theorem 4, here we have
-$|R_{N1}|\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}\right)^{2}\equiv R_{N1}^{A}+R_{N1}^{B}.$
+$|R_{N1}|\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}_{j}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\widehat{\sigma}_{j}\sigma_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\right)^{2}\equiv R_{N1}^{A}+R_{N1}^{B}.$
 For $R_{N1}^{A}$, we obtain
-$R_{N1}^{A}=\frac{(\hat{\mu}-\mu)^{2}}{N}\sum_{j=1}^{J}\frac{N_{j}}{\hat{\sigma}_{j}^{2}}\asymp\frac{1}{N}O_{p}\left(\frac{1}{N}\right)O_{p}(N)=O_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
+$R_{N1}^{A}=\frac{(\widehat{\mu}-\mu)^{2}}{N}\sum_{j=1}^{J}\frac{N_{j}}{\widehat{\sigma}_{j}^{2}}\asymp\frac{1}{N}O_{p}\left(\frac{1}{N}\right)O_{p}(N)=O_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
 by (A.20) and
-$\sum_{j=1}^{J}\frac{N_{j}}{\hat{\sigma}_{j}^{2}}=\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}(1+o_{p}(1))=O_{p}(N).$
+$\sum_{j=1}^{J}\frac{N_{j}}{\widehat{\sigma}_{j}^{2}}=\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}(1+o_{p}(1))=O_{p}(N).$
 For $R_{N1}^{B}$, we have
-$R_{N1}^{B}\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}\right\}$
-$=$ $\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\left(\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
-$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)+\left(\hat{\mu}-\mu\right)^{4}\left(\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
-$+4\left(\hat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
-$\equiv$ $R_{N1}^{B,1}+\left(\hat{\mu}-\mu\right)^{4}R_{N1}^{B,2}+4\left(\hat{\mu}-\mu\right)^{2}R_{N1}^{B,3}.$
+$R_{N1}^{B}\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}\right\}$
+$=$ $\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\left(\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\widehat{\mu}-\mu\right)^{2}-2\left(\widehat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
+$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)+\left(\widehat{\mu}-\mu\right)^{4}\left(\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
+$+4\left(\widehat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
+$\equiv$ $R_{N1}^{B,1}+\left(\widehat{\mu}-\mu\right)^{4}R_{N1}^{B,2}+4\left(\widehat{\mu}-\mu\right)^{2}R_{N1}^{B,3}.$
 For $R_{N1}^{B,1}$, we rewrite it as
 $R_{N1}^{B,1}=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\sum_{i_{3}=1}^{N_{j}}\left(e_{i_{1}j}^{2}-1\right)\left(e_{i_{2}j}^{2}-1\right)e_{i_{3}j}^{2}.$
-Note that $\mathrm{E}[(e_{i_{1}j}^{2}-1)(e_{i_{2}j}^{2}-1)e_{i_{3}j}^{2}]=\mathrm{E}[(e^{2}-1)^{2}e^{2}]\mathbbm{1}(i_{1}=i_{2}=i_{3})+\mathrm{E}[(e^{2}-1)^{2}]\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$.
+Note that $\mathbb{E}[(e_{i_{1}j}^{2}-1)(e_{i_{2}j}^{2}-1)e_{i_{3}j}^{2}]=\mathbb{E}[(e^{2}-1)^{2}e^{2}]\mathbbm{1}(i_{1}=i_{2}=i_{3})+\mathbb{E}[(e^{2}-1)^{2}]\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$.
 Then,
-$\mathrm{E}\left[R_{N1}^{B,1}\right]$ $=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(N_{j}\mathrm{E}\left[(e^{2}-1)^{2}e^{2}\right]+N_{j}(N_{j}-1)\mathrm{E}\left[(e^{2}-1)^{2}\right]\right)$
+$\mathbb{E}\left[R_{N1}^{B,1}\right]$ $=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(N_{j}\mathbb{E}\left[(e^{2}-1)^{2}e^{2}\right]+N_{j}(N_{j}-1)\mathbb{E}\left[(e^{2}-1)^{2}\right]\right)$
 $=O\left(\frac{J}{N}\right),$
 which implies $R_{N1}^{B,1}=o_{p}(N^{-1/2})$.
-Similarly, for $R_{N1}^{B,3}$, by $\mathrm{E}[e_{i_{1}j}e_{i_{2}j}e_{i_{3}j}^{2}]=\mathrm{E}[e^{4}]\mathbbm{1}(i_{1}=i_{2})+\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$, we have
-$\mathrm{E}\left[R_{N1}^{B,3}\right]=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(N_{j}\mathrm{E}\left[e^{4}\right]+N_{j}(N_{j}-1)\right)=O\left(\frac{J}{N}\right),$
+Similarly, for $R_{N1}^{B,3}$, by $\mathbb{E}[e_{i_{1}j}e_{i_{2}j}e_{i_{3}j}^{2}]=\mathbb{E}[e^{4}]\mathbbm{1}(i_{1}=i_{2})+\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$, we have
+$\mathbb{E}\left[R_{N1}^{B,3}\right]=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(N_{j}\mathbb{E}\left[e^{4}\right]+N_{j}(N_{j}-1)\right)=O\left(\frac{J}{N}\right),$
 which implies $R_{N1}^{B,3}=o_{p}(N^{-1/2})$. For $R_{N1}^{B,2}$, straightforward calculation leads to
-$\mathrm{E}\left[R_{N1}^{B,2}\right]=\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right]=\frac{1}{N}\sum_{j=1}^{J}N_{j}\sigma_{j}^{2}=O(1),$
+$\mathbb{E}\left[R_{N1}^{B,2}\right]=\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right]=\frac{1}{N}\sum_{j=1}^{J}N_{j}\sigma_{j}^{2}=O(1),$
 then $R_{N1}^{B,2}=O_{p}(1)$. Along with (A.20), we obtain $R_{N1}^{B}=o_{p}(N^{-1/2})$. Hence $R_{N1}=o_{p}(N^{-1/2})$. On the other hand, for $R_{N2}$, we first note that
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}\right)^{4}$
-$\asymp\left(\hat{\mu}-\mu\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}_{j}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)}{\widehat{\sigma}_{j}\sigma_{j}(\widehat{\sigma}_{j}+\sigma_{j})}\right)^{4}$
+$\asymp\left(\widehat{\mu}-\mu\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
 $\asymp O_{p}\left(\frac{1}{N^{2}}\right)+o_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{N}\right),$
 since
-$\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
-$\leq$ $\left\{\max_{1\leq j\leq J}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\right\}\left\{\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right\}$
-$=$ $o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right)$
+$\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
+$\leq$ $\left\{\max_{1\leq j\leq J}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\right\}\left\{\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right\}$
+$=$ $o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right)$
 $=$ $o_{p}\left(\frac{1}{\sqrt{N}}\right),$
-where we use $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}=o_{p}(1)$ from (A.21), and the last step can be obtained through similar arguments above.
-Besides, $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\tilde{\varepsilon}_{ij}^{2}/\tilde{\sigma}_{j}^{2}=O_{p}(1)$ can be derived from the procedure in (A.9) analogously.
+where we use $\widehat{\sigma}^{2}_{j}-\sigma^{2}_{j}=o_{p}(1)$ from (A.21), and the last step can be obtained through similar arguments above.
+Besides, $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\widetilde{\varepsilon}_{ij}^{2}/\widetilde{\sigma}_{j}^{2}=O_{p}(1)$ can be derived from the procedure in (A.9) analogously.
 Then we have
-$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}_{j}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
+$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\widetilde{\varepsilon}_{ij}^{2}}{\widetilde{\sigma}_{j}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
 Therefore, (3.11) is validated. This completes the proof of Theorem 7. $\square$
-Proof of Theorem 8 and 9.
+A.6 Proof of Theorem 8 and 9
 The proof of Theorem 8 is analogous to that of Theorem 2.
@@
 From (3.5) we know that
-$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}(1).$
+$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}(1).$
 Thus,
-$\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)\to_{p}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}.$
+$\widetilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)\overset{p}{\to}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}.$
 Besides,
-$\sqrt{N}\left(\tilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)$
-$=$ $\sqrt{N}\left\{\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]\right$.
+$\sqrt{N}\left(\widetilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)$
+$=$ $\sqrt{N}\left\{\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right]\right$.
 $\left.\qquad-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right\}$
-$=$ $2a^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}$
-$\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.28)
+$=$ $2a^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}$
+$\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.28)
 Under $H_{1}$, by CLT, we have
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right)-a_{1},\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)-a_{K}\right]^{\top}$
-$\to_{d}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{W}_{j},$
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\widehat{Z}_{ij}\right)-a_{1},\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\widehat{Z}_{ij}\right)-a_{K}\right]^{\top}$
+$\overset{d}{\to}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{W}_{j},$
 where $\bm{W}_{1},\ldots,\bm{W}_{J}$ are independently distributed as
@@
 with $\bm{\Lambda}_{K}^{(j)}=\left(\lambda_{kl}^{(j)}\right)_{K\times K}$ given by
-$\lambda_{kl}^{(j)}=$ $\mathrm{E}\left\{\left(\pi_{k}(Z)-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2k}}{2}\left(e^{2}-1\right)\right\}$
-$\quad\left\{\left(\pi_{l}(Z)-\mathrm{E}\left[\pi_{l}(Z)\right]\right)-d_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2l}}{2}\left(e^{2}-1\right)\right\}$
-$=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
-$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathrm{E}\left[e^{3}\right]$
-$+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
+$\lambda_{kl}^{(j)}=$ $\mathbb{E}\left\{\left(\pi_{k}(Z)-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2k}}{2}\left(e^{2}-1\right)\right\}$
+$\quad\left\{\left(\pi_{l}(Z)-\mathbb{E}\left[\pi_{l}(Z)\right]\right)-d_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2l}}{2}\left(e^{2}-1\right)\right\}$
+$=$ $\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
+$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathbb{E}\left[e^{3}\right]$
+$+\frac{d_{2k}d_{2l}}{4}\left[\mathbb{E}\left[e^{4}\right]-1\right].$
 Therefore,
-$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Lambda}_{K}^{(j)}\right),$
-and the asymptotic distribution of $\tilde{\Psi}_{K}^{2}$ in (3.15) can be derived from (A.28) through similar arguments above. This completes the proof of Theorem 8.
+$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Lambda}_{K}^{(j)}\right),$
+and the asymptotic distribution of $\widetilde{\Psi}_{K}^{2}$ in (3.15) can be derived from (A.28) through similar arguments above. This completes the proof of Theorem 8.
 The proof of Theorem 9 can be referred to that of Theorem 3 due to the similar arguments and techniques used.
-To obtain the results in Theorem 9, following the proof of Theorem 8, it suffices to show that under $H_{1L}$, $\mathrm{E}[\pi_{k}(Z)]=\delta_{N}\Delta_{k}$, $\mathrm{E}[\pi_{k}(Z)\pi_{l}(Z)]=\delta_{kl}+o(1)$, $d_{1k}=c_{1k}+o(1)$, $d_{2k}=c_{2k}+o(1)$, $d_{3k}=c_{1k}+o(1)$, $d_{4k}=c_{2k}+o(1)$, $\mathrm{E}[e^{3}]=o(1)$, and $\mathrm{E}[e^{4}]=3+o(1)$, which can be found in the proof of Theorem 3.
+To obtain the results in Theorem 9, following the proof of Theorem 8, it suffices to show that under $H_{1L}$, $\mathbb{E}[\pi_{k}(Z)]=\delta_{N}\Delta_{k}$, $\mathbb{E}[\pi_{k}(Z)\pi_{l}(Z)]=\delta_{kl}+o(1)$, $d_{1k}=c_{1k}+o(1)$, $d_{2k}=c_{2k}+o(1)$, $d_{3k}=c_{1k}+o(1)$, $d_{4k}=c_{2k}+o(1)$, $\mathbb{E}[e^{3}]=o(1)$, and $\mathbb{E}[e^{4}]=3+o(1)$, which can be found in the proof of Theorem 3.
 $\square$
-Proof of Theorem 10.
+A.7 Proof of Theorem 10
 We focus on the first case in Theorem 10, and the latter two cases can be verified analogously.
-We primarily follow well-established steps in the literature, for example, the proof of Theorem 2 in Ducharme and Lafaye de Micheaux, (2004), Lemma 1 in Kraus, (2007), and Propositions 6 and 7 in Duchesne et al., (2016).
-Note that $\{\hat{K}=k\}\subseteq\{N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\}$ for $k^{\prime}\neq k$, then we have
-$\mathrm{P}\left(\hat{K}=k\right)\leq\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\right),\quad k^{\prime}\neq k.$ (A.29)
-Under $H_{0}$, in order to show $\mathrm{P}(\hat{K}=1)\to 1$ as $N\to\infty$, we first show the contrary such that $\sum_{k=2}^{D}\mathrm{P}(\hat{K}=k)\to 0$. In fact, from (A.29) we have
-$\sum_{k=2}^{D}\mathrm{P}\left(\hat{K}=k\right)$
-$\leq$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{1}-\log N\right)$
-$\leq$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq-\log N\right)$
-$=$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N\right).$
-From Corollary 1 we know that under $H_{0}$, $N\hat{\Psi}^{2}_{k}\to_{d}\chi^{2}_{k}$ for each $k=1,\ldots,D$.
-Thus for $k=2,\ldots,D$, $\mathrm{P}(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N)\to 0$ as $N\to\infty$.
+We primarily follow well-established steps in the literature, for example, the proof of Theorem 2 in ducharme2004goodness, Lemma 1 in kraus2007data, and Propositions 6 and 7 in duchesne2016estimating.
+Note that $\{\widehat{K}=k\}\subseteq\{N\widehat{\Psi}^{2}_{k}-k\log N\geq N\widehat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\}$ for $k^{\prime}\neq k$, then we have
+$\mathbb{P}\left(\widehat{K}=k\right)\leq\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}-k\log N\geq N\widehat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\right),\quad k^{\prime}\neq k.$ (A.29)
+Under $H_{0}$, in order to show $\mathbb{P}(\widehat{K}=1)\to 1$ as $N\to\infty$, we first show the contrary such that $\sum_{k=2}^{D}\mathbb{P}(\widehat{K}=k)\to 0$. In fact, from (A.29) we have
+$\sum_{k=2}^{D}\mathbb{P}\left(\widehat{K}=k\right)$
+$\leq$ $\sum_{k=2}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}-k\log N\geq N\widehat{\Psi}^{2}_{1}-\log N\right)$
+$\leq$ $\sum_{k=2}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}-k\log N\geq-\log N\right)$
+$=$ $\sum_{k=2}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}\geq(k-1)\log N\right).$
+From Corollary 1 we know that under $H_{0}$, $N\widehat{\Psi}^{2}_{k}\overset{d}{\to}\chi^{2}_{k}$ for each $k=1,\ldots,D$.
+Thus for $k=2,\ldots,D$, $\mathbb{P}(N\widehat{\Psi}^{2}_{k}\geq(k-1)\log N)\to 0$ as $N\to\infty$.
 Therefore,
-$\sum_{k=2}^{D}\mathrm{P}\left(\hat{K}=k\right)\leq\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N\right)\to 0,$
-as $N\to\infty$, then $\mathrm{P}(\hat{K}=1)=1-\sum_{k=2}^{D}\mathrm{P}(\hat{K}=k)\to 1$, and $N\hat{\Psi}^{2}_{\hat{K}}\to_{d}\chi^{2}_{1}$ follows.
+$\sum_{k=2}^{D}\mathbb{P}\left(\widehat{K}=k\right)\leq\sum_{k=2}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}\geq(k-1)\log N\right)\to 0,$
+as $N\to\infty$, then $\mathbb{P}(\widehat{K}=1)=1-\sum_{k=2}^{D}\mathbb{P}(\widehat{K}=k)\to 1$, and $N\widehat{\Psi}^{2}_{\widehat{K}}\overset{d}{\to}\chi^{2}_{1}$ follows.
 Under $H_{1}^{\prime}$, (A.29) yields that
-$\mathrm{P}\left(\hat{K}=k\right)$ $\leq\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{K_{0}}-K_{0}\log N\right)$
-$=\mathrm{P}\left(\hat{\Psi}^{2}_{k}-k\frac{\log N}{N}\geq\hat{\Psi}^{2}_{K_{0}}-K_{0}\frac{\log N}{N}\right).$
-From Theorem 2 we know that as $N\to\infty$, $\hat{\Psi}^{2}_{k}\to_{p}0$ for $k=1,\ldots,K_{0}-1$, $\hat{\Psi}^{2}_{K_{0}}\to_{p}\bm{a}_{K_{0}}^{\top}\bm{\Sigma}_{K_{0}}^{-1}\bm{a}_{K_{0}}>0$ and $\log N/N\to 0$.
-Thus $\mathrm{P}(\hat{K}=k)\to 0$ for $k=1,\ldots,K_{0}-1$, and
-$\lim_{N\to\infty}\mathrm{P}\left(\hat{K}\geq K_{0}\right)=1-\lim_{N\to\infty}\sum_{k=1}^{K_{0}-1}\mathrm{P}\left(\hat{K}=k\right)=1.$
+$\mathbb{P}\left(\widehat{K}=k\right)$ $\leq\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}-k\log N\geq N\widehat{\Psi}^{2}_{K_{0}}-K_{0}\log N\right)$
+$=\mathbb{P}\left(\widehat{\Psi}^{2}_{k}-k\frac{\log N}{N}\geq\widehat{\Psi}^{2}_{K_{0}}-K_{0}\frac{\log N}{N}\right).$
+From Theorem 2 we know that as $N\to\infty$, $\widehat{\Psi}^{2}_{k}\overset{p}{\to}0$ for $k=1,\ldots,K_{0}-1$, $\widehat{\Psi}^{2}_{K_{0}}\overset{p}{\to}\bm{a}_{K_{0}}^{\top}\bm{\Sigma}_{K_{0}}^{-1}\bm{a}_{K_{0}}>0$ and $\log N/N\to 0$.
+Thus $\mathbb{P}(\widehat{K}=k)\to 0$ for $k=1,\ldots,K_{0}-1$, and
+$\lim_{N\to\infty}\mathbb{P}\left(\widehat{K}\geq K_{0}\right)=1-\lim_{N\to\infty}\sum_{k=1}^{K_{0}-1}\mathbb{P}\left(\widehat{K}=k\right)=1.$
 Then, for any $x\in\mathbb{R}$, as $N\to\infty$,
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{\hat{K}}\leq x\right)$ $=\sum_{k=1}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\leq x,\hat{K}=k\right)$
-$\leq\sum_{k=1}^{K_{0}-1}\mathrm{P}\left(\hat{K}=k\right)+\sum_{k=K_{0}}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\leq x\right)\to 0,$
-since $\hat{\Psi}^{2}_{k}\to_{p}\bm{a}_{k}^{\top}\bm{\Sigma}_{k}^{-1}\bm{a}_{k}>0$ for $k=K_{0},\ldots,D$. Therefore, for any $x\in\mathbb{R}$,
-$\lim_{N\to\infty}\mathrm{P}\left(N\hat{\Psi}^{2}_{\hat{K}}\leq x\right)=0,$
+$\mathbb{P}\left(N\widehat{\Psi}^{2}_{\widehat{K}}\leq x\right)$ $=\sum_{k=1}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}\leq x,\widehat{K}=k\right)$
+$\leq\sum_{k=1}^{K_{0}-1}\mathbb{P}\left(\widehat{K}=k\right)+\sum_{k=K_{0}}^{D}\mathbb{P}\left(N\widehat{\Psi}^{2}_{k}\leq x\right)\to 0,$
+since $\widehat{\Psi}^{2}_{k}\overset{p}{\to}\bm{a}_{k}^{\top}\bm{\Sigma}_{k}^{-1}\bm{a}_{k}>0$ for $k=K_{0},\ldots,D$. Therefore, for any $x\in\mathbb{R}$,
+$\lim_{N\to\infty}\mathbb{P}\left(N\widehat{\Psi}^{2}_{\widehat{K}}\leq x\right)=0,$
 which means the test is consistent against the alternatives given by (4.2). $\square$
-Intuition behind the revised limiting null distribution (4.3).
-We primarily follow the arguments in Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), and Kraus, (2007), and provide a heuristic derivation of the approximation $H(x)$ for the first case, as the latter two cases can be derived analogously.
+A.8 Intuition behind the revised limiting null distribution (4.3)
+We primarily follow the arguments in kallenberg1999data, janic2000data, and kraus2007data, and provide a heuristic derivation of the approximation $H(x)$ for the first case, as the latter two cases can be derived analogously.
 Note that
-$\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x\right)=$ $\mathrm{P}\left(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1\right)+\mathrm{P}\left(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2\right)$
-$+\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3\right).$ (A.30)
-The third term on the right-hand side of (A), $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3)$, can be neglected under $H_{0}$ (Kallenberg and Ledwina, 1995b, ).
-The event $\{\hat{K}=1\}$ is approximated by $\{N\hat{\Psi}^{2}_{1}-\log N\geq N\hat{\Psi}^{2}_{2}-2\log N\}=\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\hat{K}=2\}$ is approximated by $\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})>\log N\}$.
-We first investigate the asymptotic distribution of $(N\hat{\Psi}^{2}_{1},N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1}))^{\top}$.
-The limiting distributions of $N\hat{\Psi}^{2}_{1},N\hat{\Psi}^{2}_{2}$ are functions of a bivariate normal vector $(R_{1},R_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\bm{\Sigma}_{2})$.
+$\mathbb{P}\left(N\widehat{\Psi}_{\widehat{K}}^{2}\leq x\right)=$ $\mathbb{P}\left(N\widehat{\Psi}_{1}^{2}\leq x,\widehat{K}=1\right)+\mathbb{P}\left(N\widehat{\Psi}_{2}^{2}\leq x,\widehat{K}=2\right)$
+$+\mathbb{P}\left(N\widehat{\Psi}_{\widehat{K}}^{2}\leq x,\widehat{K}\geq 3\right).$ (A.30)
+The third term on the right-hand side of (A.8), $\mathbb{P}(N\widehat{\Psi}_{\widehat{K}}^{2}\leq x,\widehat{K}\geq 3)$, can be neglected under $H_{0}$ (kallenberg1995data).
+The event $\{\widehat{K}=1\}$ is approximated by $\{N\widehat{\Psi}^{2}_{1}-\log N\geq N\widehat{\Psi}^{2}_{2}-2\log N\}=\{N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\widehat{K}=2\}$ is approximated by $\{N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})>\log N\}$.
+We first investigate the asymptotic distribution of $(N\widehat{\Psi}^{2}_{1},N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1}))^{\top}$.
+The limiting distributions of $N\widehat{\Psi}^{2}_{1},N\widehat{\Psi}^{2}_{2}$ are functions of a bivariate normal vector $(R_{1},R_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\bm{\Sigma}_{2})$.
 We generally denote elements of $\bm{\Sigma}_{2}$ as $\begin{pmatrix}a&b\\ b&c\end{pmatrix}$ and $\rho=b/\sqrt{ac}$.
-The distribution of $(R_{1},R_{2})^{\top}$ can be constructed through two independent standard normal variables $G_{1},G_{2}$ if we denote $\tilde{R}_{1}=\sqrt{a}(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})$ and $\tilde{R}_{2}=\sqrt{c}G_{2}$, then $(\tilde{R}_{1},\tilde{R}_{2})^{\top}\overset{d}{=}(R_{1},R_{2})^{\top}$.
-Thus, $N\hat{\Psi}^{2}_{1}\to_{d}R_{1}^{2}/a\overset{d}{=}\tilde{R}_{1}^{2}/a=(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})^{2}$.
-For $N\hat{\Psi}_{2}^{2}$, straightforward computations yield that
-$N\hat{\Psi}_{2}^{2}$ $\to_{d}\left(\tilde{R}_{1},\tilde{R}_{2}\right)\bm{\Sigma}_{2}^{-1}\left(\tilde{R}_{1},\tilde{R}_{2}\right)^{\top}$
-$=\frac{1}{ac-b^{2}}\left(\tilde{R}_{1},\tilde{R}_{2}\right)\begin{pmatrix}c&-b\\ -b&a\end{pmatrix}\left(\tilde{R}_{1},\tilde{R}_{2}\right)^{\top}$
-$=\frac{1}{ac-b^{2}}\left(c\tilde{R}_{1}^{2}-2b\tilde{R}_{1}\tilde{R}_{2}+a\tilde{R}_{2}^{2}\right)$
+The distribution of $(R_{1},R_{2})^{\top}$ can be constructed through two independent standard normal variables $G_{1},G_{2}$ if we denote $\widetilde{R}_{1}=\sqrt{a}(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})$ and $\widetilde{R}_{2}=\sqrt{c}G_{2}$, then $(\widetilde{R}_{1},\widetilde{R}_{2})^{\top}\overset{d}{=}(R_{1},R_{2})^{\top}$.
+Thus, $N\widehat{\Psi}^{2}_{1}\overset{d}{\to}R_{1}^{2}/a\overset{d}{=}\widetilde{R}_{1}^{2}/a=(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})^{2}$.
+For $N\widehat{\Psi}_{2}^{2}$, straightforward computations yield that
+$N\widehat{\Psi}_{2}^{2}$ $\overset{d}{\to}\left(\widetilde{R}_{1},\widetilde{R}_{2}\right)\bm{\Sigma}_{2}^{-1}\left(\widetilde{R}_{1},\widetilde{R}_{2}\right)^{\top}$
+$=\frac{1}{ac-b^{2}}\left(\widetilde{R}_{1},\widetilde{R}_{2}\right)\begin{pmatrix}c&-b\\ -b&a\end{pmatrix}\left(\widetilde{R}_{1},\widetilde{R}_{2}\right)^{\top}$
+$=\frac{1}{ac-b^{2}}\left(c\widetilde{R}_{1}^{2}-2b\widetilde{R}_{1}\widetilde{R}_{2}+a\widetilde{R}_{2}^{2}\right)$
 $=G_{1}^{2}+G_{2}^{2}.$
-Then $N(\hat{\Psi}_{2}^{2}-\hat{\Psi}_{1}^{2})\to_{d}(\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{2}$.
-Since $(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2},\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\mathbf{I}_{2})$, we obtain that $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ are both asymptotically $\chi_{1}^{2}$ distributed and are asymptotically independent.
+Then $N(\widehat{\Psi}_{2}^{2}-\widehat{\Psi}_{1}^{2})\overset{d}{\to}(\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{2}$.
+Since $(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2},\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\mathbf{I}_{2})$, we obtain that $N\widehat{\Psi}^{2}_{1}$, $N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})$ are both asymptotically $\chi_{1}^{2}$ distributed and are asymptotically independent.
 Now we calculate the approximation $H(x)$. We will treat $H(x)$ separately for $x\leq\log N$, $\log N<x<2\log N$ and $x\geq 2\log N$. For $x\leq\log N$, note that
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,\hat{K}=2\right)\approx\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}>\log N\right)=0,$
-by $N\hat{\Psi}^{2}_{1}\geq 0$. Thus, we have the approximation
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)$
+$\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}\leq x,\widehat{K}=2\right)\approx\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}\leq x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}>\log N\right)=0,$
+by $N\widehat{\Psi}^{2}_{1}\geq 0$. Thus, we have the approximation
+$\mathbb{P}\left(N\widehat{\Psi}^{2}_{1}\leq x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\leq\log N\right)$
 $\approx$ $\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]\equiv H(x),\quad x\leq\log N,$
-from the asymptotic $\chi_{1}^{2}$ distribution and asymptotic independence of $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$. For $x\geq 2\log N$, we have
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,\hat{K}=2\right)$
-$\approx$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
-$=$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)-\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right).$
-As $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ is approximately $\chi_{1}^{2}$ distributed, we obtain
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
+from the asymptotic $\chi_{1}^{2}$ distribution and asymptotic independence of $N\widehat{\Psi}^{2}_{1}$, $N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})$. For $x\geq 2\log N$, we have
+$\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}\leq x,\widehat{K}=2\right)$
+$\approx$ $\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}\leq x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)$
+$=$ $\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)-\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}>x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right).$
+As $N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})$ is approximately $\chi_{1}^{2}$ distributed, we obtain
+$\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)$
 $\approx$ $2\left(1-\Phi\left(\sqrt{\log N}\right)\right)\approx 2\frac{\phi\left(\sqrt{\log N}\right)}{\sqrt{\log N}}=\sqrt{\frac{2}{\pi N\log N}},$
 by the fact that $t\phi(t)/(t+1)\leq 1-\Phi(t)\leq\phi(t)/t$ for $t\to\infty$. Besides,
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
-$\leq$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x\right)$
-$\leq$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>2\log N\right)$
+$\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}>x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)$
+$\leq$ $\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}>x\right)$
+$\leq$ $\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}>2\log N\right)$
 $\approx$ $\exp\left\{-\frac{1}{2}\times 2\log N\right\}$
 $=$ $\frac{1}{N},$
-where we use the distribution function of the limiting $\chi^{2}_{2}$ for $N\hat{\Psi}^{2}_{2}$.
-Hence $\mathrm{P}(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N)$ converges to 0 much slower than $\mathrm{P}(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N)$, and thus the latter can be neglected.
+where we use the distribution function of the limiting $\chi^{2}_{2}$ for $N\widehat{\Psi}^{2}_{2}$.
+Hence $\mathbb{P}(N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N)$ converges to 0 much slower than $\mathbb{P}(N\widehat{\Psi}^{2}_{2}>x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N)$, and thus the latter can be neglected.
 Therefore, for $x\geq 2\log N$, we have the approximation
-$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)+\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
+$\mathbb{P}\left(N\widehat{\Psi}^{2}_{1}\leq x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\leq\log N\right)+\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)$
 $\approx$ $\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]+2\left[1-\Phi\left(\sqrt{\log N}\right)\right]$
 $\equiv$ $H(x),\quad x\geq 2\log N.$
-For $\log N<x<2\log N$, Kallenberg and Ledwina, 1995b suggested the linearization as follows
+For $\log N<x<2\log N$, kallenberg1995data suggested the linearization as follows
 $H(x)=H(\log N)+\frac{x-\log N}{\log N}[H(2\log N)-H(\log N)],\quad\log N<x<2\log N.$
-This completes the derivation of the approximation (4.3) for the null distribution of the test statistic $N\hat{\Psi}^{2}_{\hat{K}}$. $\square$
-Implement issues in simulation studies.
-In our simulation study, we adopt the orthonormal Legendre polynomials on $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$.
+This completes the derivation of the approximation (4.3) for the null distribution of the test statistic $N\widehat{\Psi}^{2}_{\widehat{K}}$. $\square$
+A.9 Implement issues in simulation studies
+In our simulation study, for the proposed smooth tests, we adopt the orthonormal Legendre polynomials on $[0,1]$ for $\{\pi_{k}\}_{k=1}^{\infty}$.
 The properties of orthonormal Legendre polynomials lead to the following expressions for constants $c_{1k}$ and $c_{2k}$:
@@
 Therefore, in our experiments, we only compute the non-zero elements based on the above two formulae for $k=1,\ldots,K$ with $K\leq 5$.
-Additional simulation results.
-The results of Experiment III are reported in Tables 14, 15, 16 and Figure 5.
-The main findings are broadly consistent with those in Experiment III, with two additional observations.
-First, Table 14 indicates that the limiting $\chi^{2}_{1}$ distribution for the data-driven test yields valid control of the Type I error.
-Second, the empirical distribution of $\hat{K}$ exhibits increased stability under both the null and the alternative hypotheses, with instances where it degenerates to a single value with probability one.
-Both phenomena are due to the increase in total sample size resulting from the larger number of groups, which is consistent with the theoretical results.
-Table 14: Empirical rejection rates under $H_{0}$ in Experiment III, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
-The empirical power under $H_{1}$ equals 1 across all settings.
-$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
-10 0.028 0.046 0.062 0.036 0.054 0.060 0.048
-20 0.046 0.066 0.078 0.052 0.068 0.056 0.052
-30 0.050 0.056 0.044 0.040 0.048 0.044 0.042
-40 0.054 0.054 0.042 0.040 0.044 0.058 0.052
-50 0.052 0.062 0.052 0.056 0.054 0.048 0.042
-60 0.062 0.050 0.060 0.048 0.048 0.052 0.046
-70 0.068 0.052 0.052 0.054 0.044 0.058 0.054
-80 0.052 0.034 0.044 0.062 0.052 0.048 0.046
-90 0.048 0.048 0.034 0.050 0.058 0.042 0.040
-100 0.064 0.048 0.052 0.048 0.052 0.058 0.054
-110 0.044 0.048 0.058 0.058 0.032 0.060 0.056
-120 0.052 0.058 0.066 0.044 0.052 0.080 0.078
-130 0.046 0.050 0.038 0.058 0.062 0.058 0.052
-140 0.058 0.046 0.032 0.054 0.046 0.056 0.050
-150 0.046 0.040 0.060 0.050 0.054 0.038 0.038
-Table 15: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment III
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-10 0.994 0.006 0 0 0
-20 1 0 0 0 0
-30 0.992 0.008 0 0 0
-40 0.994 0.006 0 0 0
-50 0.988 0.012 0 0 0
-60 0.994 0.006 0 0 0
-70 0.998 0.002 0 0 0
-80 0.992 0.008 0 0 0
-90 0.996 0.004 0 0 0
-100 0.994 0.006 0 0 0
-110 0.996 0.004 0 0 0
-120 0.998 0.002 0 0 0
-130 0.990 0.010 0 0 0
-140 0.994 0.006 0 0 0
-150 0.998 0.002 0 0 0
-Table 16: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment III
-$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
-10 0 0 0 0.120 0.880
-20 0 0 0 0.024 0.976
-30 0 0 0 0 1
-40 0 0 0 0 1
-50 0 0 0 0 1
-60 0 0 0 0 1
-70 0 0 0 0 1
-80 0 0 0 0 1
-90 0 0 0 0 1
-100 0 0 0 0 1
-110 0 0 0 0 1
-120 0 0 0 0 1
-130 0 0 0 0 1
-140 0 0 0 0 1
-150 0 0 0 0 1
-Figure 5: The sample means and error bars of $\hat{K}$ in Experiment III
+A.10 Additional simulation results
+The results for Experiment III are reported in Tables 13, 14 and Figure 5.
+They are broadly consistent with those obtained in Experiment III, with several additional findings.
+First, when $J=10$ and $m$ is relatively large, the total sample size $N$ exceeds the applicability range of the Shapiro–Wilk test (which requires $N\leq 3000$), and hence the corresponding results are not reported in Table 13.
+In contrast, our proposed test is not subject to this limitation and remains applicable.
+Second, Table 13 shows that the limiting $\chi^{2}_{1}$ distribution for the data-driven test provides valid control of the Type I error.
+Third, the empirical distribution of $\widehat{K}$ becomes more stable under both the null and alternative hypotheses, in some cases degenerating to a single value with probability one.
+These observations can be attributed to the increase in total sample size induced by a larger number of groups, which is in line with the theoretical results.
+Table 13: Empirical rejection rates under $H_{0}$ in Experiment III. The empirical power under $H_{1}$ equals 1 for all sample sizes and all considered tests.
+Smooth tests JB KS
+$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
+10 0.028 0.046 0.062 0.036 0.054 0.060 0.048 0.050 0
+20 0.046 0.066 0.078 0.052 0.068 0.056 0.052 0.040 0
+30 0.050 0.056 0.044 0.040 0.048 0.044 0.042 0.054 0
+40 0.054 0.054 0.042 0.040 0.044 0.058 0.052 0.046 0
+50 0.052 0.062 0.052 0.056 0.054 0.048 0.042 0.066 0
+60 0.062 0.050 0.060 0.048 0.048 0.052 0.046 0.050 0
+70 0.068 0.052 0.052 0.054 0.044 0.058 0.054 0.040 0
+80 0.052 0.034 0.044 0.062 0.052 0.048 0.046 0.042 0
+90 0.048 0.048 0.034 0.050 0.058 0.042 0.040 0.044 0
+100 0.064 0.048 0.052 0.048 0.052 0.058 0.054 0.050 0
+110 0.044 0.048 0.058 0.058 0.032 0.060 0.056 0.048 0.002
+120 0.052 0.058 0.066 0.044 0.052 0.080 0.078 0.046 0
+130 0.046 0.050 0.038 0.058 0.062 0.058 0.052 0.046 0
+140 0.058 0.046 0.032 0.054 0.046 0.056 0.050 0.056 0
+150 0.046 0.040 0.060 0.050 0.054 0.038 0.038 0.042 0.002
+Table 14: Empirical frequency of $\widehat{K}$ in Experiment III
+Under $H_{0}$ Under $H_{1}$
+$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
+10 0.994 0.006 0 0 0 0 0 0 0.120 0.880
+20 1 0 0 0 0 0 0 0 0.024 0.976
+30 0.992 0.008 0 0 0 0 0 0 0 1
+40 0.994 0.006 0 0 0 0 0 0 0 1
+50 0.988 0.012 0 0 0 0 0 0 0 1
+60 0.994 0.006 0 0 0 0 0 0 0 1
+70 0.998 0.002 0 0 0 0 0 0 0 1
+80 0.992 0.008 0 0 0 0 0 0 0 1
+90 0.996 0.004 0 0 0 0 0 0 0 1
+100 0.994 0.006 0 0 0 0 0 0 0 1
+110 0.996 0.004 0 0 0 0 0 0 0 1
+120 0.998 0.002 0 0 0 0 0 0 0 1
+130 0.990 0.010 0 0 0 0 0 0 0 1
+140 0.994 0.006 0 0 0 0 0 0 0 1
+150 0.998 0.002 0 0 0 0 0 0 0 1
+Figure 5: The sample means and error bars of $\widehat{K}$ in Experiment III